Admin CHW Pump Replacement

12/26/16 – Contract sent to Tony for approval on this project.
01/26/17 – DK says controls will be done my same team as controls project. Gleason has sent all required paperwork to Tony.
02/01/17 – Project kickoff to be schedule soon. CHW pump size will need to be coordinate with new chiller.
02/16/17 – Kicked off project with Tony, Bert and Gilbert. We are collecting chiller information now and rates for LCC equations. We will choose the chiller during the LCC and design around one chiller during design.
03/09/17 – Pump is on hold pending new size from LCC equations.
03/15/17 – Emailed Dave Gleason for update as work was supposed to begin last week and be done by end of March.
03/21/17 – Tony talked to Dave and they are starting next week with pipe fitting work. Steve Yockey is waiting on TEI to provide new pump size for selection work.
04/11/17 – ThermalTech provides new bypass size and flow for DK to install the bypass.
04/26/17 – TTC checked in with Dave Gleason on progress. Given the cooling season is here, they are going to have to work weekends and off time to fix these issues. The new bypass size was sent to DK but I do not know if it was installed yet.
05/02/17 – TTC checked up on project again with DK for progress.
05/25/17 – Debra stated they will replace the pumps and piping at the same time, later in June.
06/28/17 – Met Debra today on project update and they are getting employees finalized for this project understanding it will be weekend work at this point. Pumps are on order, VFD”s are a new change order, TTC to give to Tony, will be from new project on new PO for up charges.
07/18/17 – Met with Design time to redesign the piping for the new larger pumps on teh CT and CHLR job. Spoke with the manufacturer about sizes and head. Debra submitted new change order for entire project. TTC to estimate the entire job again for Bert with Construction budgets.
08/10/17 – We are waiting for the weather to break to start the pump replacement.
08/26/17 – Debra to drain system and install isolation valves Saturday but this was postponed due to material arrival issues. They will try the week after Labor Day.
09/25/17 – Isolation valves are installed and working. DK needs to insulated still. TTC met with DK and went over remainder of project on site. The scope of the original project lines up perfect with the chiller replacement project except this project does not need to install the bypass piping on the fifth floor. The CHW pumps will be installed by DK in this project, the VFD”s and submittals must be ordered ASAP, and the change orders have been sent for approval to DK and the County.
11/2/17 – Pump replacement work is schedule for December 2017. Debra is currently on site and working on this job.
02/02/18 – Punchlist distributed to Debra Kuempel for project closeout.
04/04/18 – Only insulation remains to be completed. DK committed to wrapping this project up including controls by 4/30/18.
04/25/18 – Debra stated that all work on this project is complete including insulation.  Final paperwork is being completed by DK now for all closeout paperwork.  The final invoices are being delivered to Tony with a cover letter or remaining items.
08/02/18 – TTC inspected all final work including insulation and it looks really good.  No additional punchlist items.  Project is ready to archive.


SPHQ – RTU-5 Replacement

11/14/16 – Prebid meeting was held on site.

11/15/16 – TTC investigate using the existing architectural and mechanical plans to reword the ductwork on RTU-5 so that the unit will keep its ductwork up on the second floor and penetrate the side wall on the second floor and intercept the ductwork where it comes up through the floor. TTC spoke to Kevin Maue, Greg Smorey and Tom Speer about the original build today.

11/16/16 – Addendum #1 was sent out this morning for bidding answering prebid questions.

11/18/16 – Addendum #2 was sent out.

11/22/16 – Addendum #3 was sent out.

12/07/16 – Contracts sent to Tony for Debra Kuempel. Waiting on approval and kickoff meeting.

01/13/17 – Discussed equipment vendor selection with Feldkamp and JCI. Carrier was submitted but they are probably going to purchase JCI as an equal vendor for this project. Spoke to Jackson and Tony about this.

02/02/17 – Precon schedule for 2/14.

02/17/17 – Met contractor on site to review possible new placement and routing for ductwork on this job. ThermalTech to draw up new location and routing and get THP approval.

02/27/17 – Revised downflow design completed and sent to contractor for pricing on Thursday, revised again on Friday and emailed, revised again on Monday and sent to contractor.

02/28/17 – Final design and schedules revised again and sent to Feldkamp and Owner. Submittals rejected twice this week but third submittal was sent back Furnish as Corrected. Also told Tony today that the redesign used $2040 of the const budget and leaves about $500 for Const Services.

03/09/17 – Unit is ordered, a revised set of submittals was sent to Feldkamp this week. A controls meeting is scheduled for 3/22 at ThermalTech with Alpha and Habegger.

03/15/17 – TEI sent PO provided by Tony to DK for processing and requested project schedule from Feldkamp for this job.

03/21/17 – THP – THP and TTE had a conference call to discuss scope and fee. TTE to review fees with County.03/21/17 – Controls meeting rescheduled until 3/28 at 9am at TEI.

04/03/17 – TTE requested a formal proposal.04/21/17 – Control submittals have arrived. ThermalTech to review and return ASAP.

04/11/17 – THP provide a proposal for review.04/25/17 – Updated Drawings were given to Jackson and Feldkamp was on site locating the curb on the roof per the new structural drawings. Unit schedule to arrive in May for installation.

04/13/17 – THP proposal approved.

04/25/17 – THP analyzed AHU loading and proposed shifted the unit. Unit location was shifted. THP provided roof top support channel detail.

05/19/17 – TTC stopped by for field inspection. Curb looks good, ductwork looks good, fire dampers look good. DK trying to locate electric for new power to pull new feeders. Pictures from roof support work appeared to be acceptable to ThermalTech and THP.

06/27/17 – Project is complete and we are punchlisting the work.

07/19/17 – Punchlist was distributed today to both Debra Kuempel and Feldkamp Enterprises.

08/03/17 – Met Feldkamp on site to get quote for new boiler options using allowance money leftover on this job.

08/17/17 – Met Jackson onsite today and collected more information for the Alpha quote. Talked to DK and they are coming out to change the restrictor plate on SVH-204 soon, they will also be giving Alpha the K values to program into the VAV boxes on the FE system so we can see the correct CFM’s. I also emailed Alpha a full email explaining that I wanted to simplify and reduce their scope and price on this job and that all of the controls we need are already installed and on the graphics.

08/30/17 -All prices are in for the heating boiler and we can award this work to Feldkamp as part of this job. Still working on a cost from Alpha.

09/08/17 – When the electrical inspector (Ken Spille) came to the field for the final, he had the wrong drawings for the work with him. DK contacted TEI (Mike Barth) and we delivered another copy of the correct drawings to IBI. Ken told DK in the field that once those drawings were dropped off and processed at IBI, that he would get the final for the permit. IBI is now telling DK that they have to call and schedule a final inspection for the project (although they have already had a final), and that IBI cannot close out the project. We told DK that we need to get this closed out, and since the inspector is not recalling his field statements, please work this out and get the final completed. DK working with IBI to resolve.

09/25/17 – Feldkamp and DK were both on site last week installing the new hot water boiler form Lochinvar. We went with the 155,000 btuh boiler for this installation in the fire tube design which is a new offering from Lochinvar. The piping is connected, the exp tank is relocated, the gas piping is connected and the venting is hooked. Feldkamp confirmed the venting is ready and clear of caps and ready to use. This week Feldkamp is to startup the boiler and Alpha controls or Feldkamp will complete the installation of the systems. TTC stopped by last Thursday but not contractors were on site, TTC may stop by again to review the work if he can around his meetings and vacation.

10/24/17 – Boiler is up and running. The controls enables and disables the boiler based on outdoor air temperature. The owner has the boiler set to 140F and can adjust as needed seasonally. TTC sent out the allowance usage and change order for this project and asked Feldkamp to close it all out. The roofer came out and applied another patch to the job around the unit to quell a roof leak and hopefully that works. Everything seems to be up and running well.

10/25/17 – Allowance usage came back from Feldkamp and was forwarded to Tony Matre for signature. Boiler went out on alarm this morning. Feldkamp dispatched to investigate.

11/28/17 – In the past month we have had Feldkamp, Alpha and Corken all on site working on the boiler system. The boiler is running well and Feldkamp has upgraded the gas supply to the boiler, added the combustion tee for tuning up the boiler plant and added the neutralizer to the boiler drain. Alpha has reprogrammed the hot water pump to run through the system at all times and we are awaiting the controls upgrade to supply the setpoint to the boiler. Then Corken can come in and provide a full startup on the boiler.
03/30/18 – After a lot of work to get revised O&M’s for this job from the contractor that TTC felt were good enough for the owner, the final PDF was downloaded and printed by ThermalTech and three O&M’s were created for Tony Matre. I am not sure why TEI is making O&M’s for the contractors nowadays but in an effort to bring the project to completion, they were made and UPS ground to Facilities. The only item we know is open is the CAFM updates for these drawings.


800 Broadway Polygonal Pavilion Roof Repair

Current Comments (10/26/2011)
October 26, 2011 Update THP develops test repair options and met with SSRG to review for pricing. THP to submit separation SOSS for contractor work.

Past Comments

Due to findings, weather sensitive repairs, and schedule; THP proposed completing a safety survey and trial repairs. County requested a letter report summarizing needed work to be possibly incorporated into a larger 2012 project. A SOSS for the letter reported was requested.

September 23, 2011 THP met with County to review conditions of roof.

Early September THP completed field review and design analysis.

August 3, 2011 SOSS approved.

July 18, 2011 THP Submitted SOSS for review.


2011 General Structural Services

Current Comments (10/26/2011)
Hourly budget funds exhausted for F24A. THP directed to submit for additional $5,000 as F24B.

Past Comments

July 26, 2011 County approved SOSS F24A.

July 18, 2011 THP submitted SOSS F24A for additional funds.

Hourly budget funds exhausted; Bert Watts requested a SOSS for the second half of the year.

Bert Watts requested THP provide a SOSS for the study and repair of the 800 Broadway limestone panels.

THP reviewed reported cracking of Hillcrest Oak Cottage and provided a summary letter on May 5, 2011. Bert Watts requested an SOSS to complete a study of the cracking issues.

THP completed initial review of 800 Broadway limestone condition and provided summary email for review.

THP reviewing 800 Broadway LEED EB proposal structure.

May 5, 2011 SOSS approved.

April 28, THP provided SOSS.

April 24, 2011 SOSS requested SOSS.


Parkhaus Garage Repairs

Current Comments (10/26/2011)
October 26, 2011 Update – Contractor noted they are waiting on Joint and several warranty from the manufacturer, which is anticipated later this week.

Past Comments

All pay requests minus retainage approved; waiting on warranty and closeout paperwork.

August 31, 2011 Substantial letter completion letter issued; along with final deduct change order due to liquidated damages.

All work complete as of early August.

July 22, 2011 RAM issued via email that the new hand rail embed should be fabricated and on site at the end of the week of July 25.

July 12, 2011 THP issued new hand rail embed replacement details to RAM. RAM submitted details for embeds to be fabricated.

July 7, 2011 THP approved and forwarded to Hamilton County, RAM request for payment, Pay Application No 4, in the amount of $20,617.20.

July 6, 2011 THP noted, RAM damaged some existing handrail embeds when removing existing handrail. THP to review conditions and issued repair details.

July 5, 2011 RAM re-mobilized on site to begin replacement of the handrails at the NW stair and liquidated damages begin.

June 9, 2011 THP approved and forwarded to Hamilton County, RAM request for payment, Pay Application No 3, in the amount of $29,794.20.

May 26, 2011 THP approved, with exceptions noted, the steel hand rail shop drawings for NW stair tower.

May 23, 2011 RAM demobilized from the project site so Hamilton County has full use of the garage facility. RAM will re-mobilize on site once the steel handrail shop drawings and fabrication are complete. Per liquidated damages letter RAM is scheduled to return for installation of the steel hand rail on July 5, 2011.

May 25, 2011 steel hand rail shop drawings for NW stair tower received.

May 23, 2011 substantial completion date but RAM has not satisfied the substantial completion date requirements. No steel hand rail shop drawings submitted for NW stair tower. Liquidated damages letter pertaining to steel hand rail to be issued by THP per Bert Watts (Hamilton County) direction and approval. Steel hand rail shop drawings and fabrication were never weather dependent work items.

May 23, 2011 THP issued Field Report No 05 for Phase 2 (final) membrane strip installation.

Weekend of May 20, 2011, RAM completed vault slab placement over anodes, and remaining Phase 2 membrane strip installations. Inclement weather postponed this phase of membrane installation till this date.

May 17, 2011 Anode testing report by Bushman and Associated issued.

May 13, 2011 THP issued Field Report No 04 for completion of Phase 1 membrane strip installation. THP also approved and forwarded to Hamilton County, RAM request for payment, Pay Application No 2 in the amount of $54,344.40.

May 10, 2011 Membrane manufacturer (Neogard) letter pertaining to exposed membrane aggregate being compromised was issued.

May 6, 2011 THP issued Field Report No 03 for Phase 1 membrane installation. Inclement weather prevented completion of all of Phase 1 membrane strip installation areas.

May 6, 2011 Final T&M tickets submitted by RAM for anode relocation.

April 29, 2011 Change Order No 01 work items approved by all parties for a deduct to the projects Cash Allowance in the amount of $8,735.00. Total project budget remains unchanged.

April 26, 2011 THP issued Change Order No 01 to reflect Bulletin No 01 pricing so RAM may begin the additional work. This Change Order added two weeks to the substantial completion date. New substantial completion date is May 23, 2011.

April 25, 2011 Hamilton County accepted Bulletin No 01 pricing which was within the projects original Cash Allowance.

April 25, 2011 RAM submitted Bulletin No 01 pricing.

April 21, 2011 RAM began membrane work on Levels 2 through 5.

Week of April 18, 2011 Anode placement began in the vaults.

April 15, 2011 THP issued Bulletin No 01 to RAM for pricing for additional concrete and waterproofing repairs.

April 13, 2011 Progress Meeting.

April 8, 2011, THP issued Field Report No 02.

April 7, 2011 RAM completed construction joint sealant replacement on Levels 2 through 5.

As of March 24, 2011 RAM has not submitted steel shop drawings for hand rail.

March 18, 2011 THP approved RAM’s Pay Application No 1, in the amount of $40,194.80, and forwarded to Hamilton County.

March 3, 2011, THP issued Field Report No 01.

Anode installation scheduled to begin week of April 4.

Weather conditions for this weekend appear favorable, so RAM may perform sealant replacement April 1 weekend.

January 6, 2011 preconstruction meeting completed. Work to scheduled to begin February 7 to allow for weather.

THP has all submittals except cathodic protections.

Contract approved.

County commissioners reviewing contract.

Contract signed by Ram reported to be returned to County.

September 27, 2010 THP sent updated contract for County review.

September 9, 2010 THP sent draft contract for County review.

Bids received and Ram Construction apparent low bidder.

Bids due August 18, 2010.

Pre-bid meeting scheduled for August 4, 2010 at 10:00 a.m.

Project out to bid on July 23, 2010.

Project Documents sent to County for review on July 19, 2010.

THP and Jim Bushman accessed vaults to confirm conditions for implementing corrosion protection system on June 23, 2010.

THP is coordinating work with corrosion consultant.

Parkhaus SOSS approved May 14, 2010.

SOSS submitted 3/3/10.

THP anticipates SOSS delivery by 3/5/2010. (2/24/10

Kickoff meeting complete.

THP needs to schedule a kickoff meeting.

THP working with subcontractor to develop proposal.


1021 Patrol Headquarters Waterproofing Repairs

Current Comments (10/26/2011)
October 19, 2011 THP, SSRG and County met to review the leak to determine if work is warranty or a new item. Due to lack of water/leak; the team decided to wait until rain/leak event and review the condition at that time.

New leak reported in same switch gear room in late September during a heavy rain.

Past Comments

September 15, 2011 County finalizing process of SSRG final pay application.

September 14, 2011 SSRG sent prevailing wage information to County.

September 11, 2011 THP forwarded final pay application and close-out manual to the County.

SSRG completed punch list items early August and THP verified items were complete.

July 13, 2011 THP published punch list to SSRG. Steven noted the work will be completed and THP is holding onto the final pay application (retainage) until the work is complete.

THP is waiting on warranties to process final pay application.

THP plans to punch on June 30, 2011.

Change Order No. 1 work (Window repairs) complete as of Friday, 5/20.

Change Order No. 2 sign by HGC, awaiting County approval.

THP forwarded Change Order No. 2 for the urethane injection of the electrical and sprinkler rooms to HGC for signature on 5/16.

During heavy rains of late April, leaking at the switch gear room and sprinkler room (new leak) occurred. THP recommended injection of crack from interior of both room. April 25, 2011 THP and HGC met to discuss the condition and request pricing.

During restoration repair, additional window sill and steel structure repairs associated concrete sill were identified. THP requested cost from HGC to complete the repairs. Change Order #1 approved by the County.

Subsurface waterproof and exterior injection work complete by HGC.

January 27, 2011 Work on hold until weather is within allowable specifications (early spring). The County requested all submittals and background checks be completed this February.

March 30, 2011 THP sent HGC the change Order form for the repair work associated with the window deterioration.

March 28, 2011 THP received approval from the County to proceed with a change order for HGC to repair the deteriorated window.

March 25, THP received final pricing from HGC regarding the work to repair the deteriorated window.

March 25, 2011 HGC finished the drain tile and landscaping work which completed all base bid work.

March 23, 2011 THP requested pricing from HGC for associated window deterioration.

March 23, 2011, HGC completed injection repairs and installation of the drain tile / back fill.

January 27, 2011 Work on hold until weather is within allowable specifications (early spring). The County requested all submittals and background checks be completed this February.

THP is waiting on submittals and contractor background check.

County conducted preconstruction meeting on 12/29.

Project needs to schedule a preconstruction meeting.

The requested project schedule and submittals from Structural System.

Contract approved.

October 29, 2010 the County provided updates to the proposed Contract. THP updated and returned the contract.

October 8, 2010 THP sent draft contract to County for review.

September 30, 2010. Bids were within budget and THP recommended Structural Systems (HGC).

Bids are due September 30, 2010.

September 21, 2010 Addendum #1 was issued, which was the meeting minutes from the prebid meeting.

September 17, 2010 bid documents were sent to HGC, ZSR, Lithko Restoration, and Langenheim & Thomson Company.

September 16, 2010 THP provided bid documents to the County.

September 9, 2010 County provide comments to review documents.

September 3, 2010 THP provided review documents to the County for review. Project estimate is below County threshold for publically bid project, so project bid is by invitation.

County requested construction budget and THP reviewing.

THP provided owner review documents on August 23, 2010.

County approved drainage system.

County reviewing THP proposed approaches of a new slab or drainage system (THP on hold).

THP and County had kick of meeting on June 21, 2010 to complete repairs recommended in report

THP issued letter report on June 18, 2010.

THP arranging with contract excavation of wall area.

Excavation needed for investigation is planned for mid-March, due to adjacent utility lines and required handwork.(2/24/10)

SOSS approved 2/3/2010.

THP submitted SOSS 2/1/2010.

THP needs to submit SOSS and complete survey.

Kick off meeting completed with Bert Watts and John Millar.


Memorial Hall 2010 Facade Assessment

Current Comments (4/27/2011)
Project complete.

Past Comments

Can project be closed out?

March 8, 2011 report entered into Archibus.

March 4, 2011 report sent to County.

THP working on report.

January 3, 2011 THP & Lithko completed a water test of the south wall parapet.

January 2, 2011 Lithko restoration a safety survey to remove all loose stone material. THP completed façade lift inspection during safety survey.

THP working on project.

January 3, 2011 THP & Lithko completed a water test of the south wall parapet.

January 2, 2011 Lithko restoration a safety survey to remove all loose stone material. THP completed facade lift inspection during safety survey.

December 10, 2010 THP completed initial visual survey. Survey revealed loose sandstone and THP recommended completing a knock down survey by a contractor. Due to weather conditions repairs should occur in the spring. THP getting pricing from the contractor.

November 15, 2010 SOSS approved.

October 26, 2010 THP submitted SOSS for review.

October 26, 2010 THP and County completed kick off meeting.

October 19, 2010 County requested SOSS from THP.

August County and THP discussed preliminary Memorial Hall assessment numbers.


CAB 10th Floor Community Development Expansion

2/9/10-Meeting with Bert Watts at Community Development

2/18/10- Issued SOSS to Bert Watts.

2/19/10-Bert questioned GBBN verses TEI scope and permitting costs.

2/22/10-Terry Cannon of TEI issued email with scope, SOSS and permitting resolutions.

3/6/10- Received signed SOSS.

3/8/10- Met with Mike Florimonte to field measure adjacent storage space. Receive directive to eliminate one workstation for a total of 3. The fourth area to become storage.

3/15/10-Received Purchase Order (538771) from Bert Watts.

3/17/10-Susan Walsh approved rotated third workstation.

3/31/10-Final Design Development drawings submitted to Bert Watts for Susan Walsh’s review.

4/6/10- Building Permit Application, Drawings, Electronic Submittal and Check delivered to Cincinnati Buildings and Inspections.

4/21/10- Approved Building Permit paid for and picked-up from Cincinnati Buildings and Inspections.

4/22/10- Approved Permit and Drawings Delivered to Bert Watts.

5/18/10- Inspector Larry Keifer has requested an additional sprinkler head be added in the expansion space.

6/24/10- Sprinkler permit has been filed and still pending.

7/29/10- Bert Watts declared this project closed and ready for Archives.


Sheriff HQ – Electrical Device Identification and Arc Flash Study

05/30/08 – Ginter sent the panelboard schedule in excel format. TEI forwarded to County and sent allowance usage certificate to Ginter for signing. Ginter has to provide an approved IBI permit at conclusion of this project.

04/09/08 – All Work is complete. TEI will finish and stamp the permit drawing for Ginter. Ginter will get IBI for making the conference room on emergency panel for colonel. This will use the last of the allowance money and closeout the project.

01/07/08 – Ginter has completed punchlist items. MAH will conduct back check when meeting on extra work takes place. The Colonel would like to have some receptacles on emergency power in the conference room so they can use as a command center. The work will be discussed with final approval from Tony Matre.

12/14/07 – Change order was requested for Ginter to access the $2000 allowance. TEI has requested additional itemized backup and an explanation of the additional devices from Ginter. Jim Ginter sent the additional items to TEI and MH is reviewing for approval today. MH created the punchlist for the project and is distributing the punchlist to the contractor today as well.

12/11/07 – Construction meeting was conducted and punchlist was created moving toward a final closeout. Punchlist items will take a few weeks because of the lead time in ordering engraved labels. Aside from some misunderstanding in the labeling requirements the work appears very thorough. Ginter turned over mark-ups. Ginter to submit pay apps and TEI to write a deduct change order for unused allowance.

11/30/07 – As promised Ginter reported today that this project is complete. TTC responded back to their email that we would use the const mtg for our punchlist and to advise about allowance usage on this project so TEI could write a deduct change order for closeout. TTC also sent all the closeout documents to Ginter for use on this job.

11/28/07 – Ginter provided an update today stating the the ID project was going well. TEI will setup a const mtg the week of 12/10.

11/09/07 – TTC distributed Pre Const Mtg minutes from last week. Ginter is on the job and ready to complete by the end of this year.

08/31/07 – TTC received word from HCFD that the insurance issue has been resolved. TEI awaiting word from HCFD so that this project can move into the Construction Phase.

08/15/07 – NR wrote: “Any idea about when this project is going to be ready for construction?” Bert( HCFD) responded: “I am still working out Insurance issues with Michael Ginter, he is working on getting me the correct Certificate of Insurance. Jim and I are working on getting the LEADS background checks completed.” TEI awaiting word from Hamilton County Facilities Department to move into construction phase.

08/10/07 – TEI awaiting word that the contract has been signed by Ginter Electric so that this project can move into the construction phase.

07/30/07 – NR received word from Bert Watts at the Hamilton County Facilities Department that the contract was sent to Ginter Electric last week for their signature. TE is awaiting word that the contract has been signed so that this project can move forward with the construction phase.

06/29/07 – TE awaiting word that the contract has been signed to that this project can move into the construction phase.

06/22/07 – Contract has been awarded to Ginter Electric. The contract has been sent to Mark Donnelly; TE is awaiting signed contract to proceed with pre-construction meeting.

06/15/07 – Hamilton County Facilities is reviewing the bid tabulation. TE is awaiting word about which contractor has been awarded the contract so that we can move forward with this project.

06/08/07 – This project opened for bid yesterday. TE is awaiting the bid tabulation from the Hamilton County Purchasing Department.

06/01/07 – This project is set for bid opening on June 7, 2007.

05/25/07 – The pre-bid meeting was held on site today. Those in attendance included Kevin and Steve Kallmeyer of Kallmeyer Electric, Jim Ginter of Ginter Electric, Nick Richmond of ThermalTech Engineering, and Walt Hendrick/ Mike Jackson/ Ken Caruso of Hamilton County. A site tour was given as well as a specification and drawing review. This project is expected to go to bid on June 7th of 2007.

05/21/07 – Project went out for bid on May 18, 2007. Pre-bid meeting has been scheduled for Friday, May 25 at 10:00am. Meeting place is the lobby of the Patrol Headquarters. NR contacted Captain Mark Hendrick to discuss working hours. It was decided that normal working hours would best suit this project because of limited access to certain areas of the facility. NR sent an email to Mark Donnelly informing him of this.

04/27/04 – Drawings for the Device Identification portion of this project are complete and will be present for Hamilton County review in the meeting today, April 27 of 2007. The Arc Flash Labels are at the printer and should be completed within the next 2 weeks.

04/20/07 – Arc Flash Labels are being dropped off to the printer next week. The Device Identification portion of this project is complete and the drawings will be complete next week.

04/13/07 – Single Line Diagrams have been established and Arc Flash Study has begun. TE found discrepancies with the As-Built drawings and what was found in the field so they are making another field visit today, April 13, 2007. This visit is meant to trace the source of the panels in the barns and out-buildings. There is an automatic transfer switch in the Property Room that needs to be traced also. Dan Myer of TE will also be on site to meet with a county employee to locate above ceiling equipment.

04/06/07 – Kick-off meeting was held. Project is moving along as expected.

03/30/07 – Kick-off meeting is scheduled for next Thursday, April 5th of 2007.

03/23/07 – Project progressing as expected. Kick-off meeting tentatively scheduled for the first week in April. TE to set up meeting with Hamilton County Facilities.

03/20/07 – Fieldwork is progressing and is about 80% complete.

03/14-07 – NR created project journal and filled in back data from previous meetings. Project awaiting signed SOSS to be returned to TE. 90% of field work has been completed for the Device Identification portion of this project. Field notes not inputted electronically per awaiting signed SOSS.



09/29/10 – Project closed and archived.

08/16/10 – We talked to Debra Kuempel today. O&M’s were submitted and I believe sent to the County last November. We have been holding for the final change order. Kevin and Donna at DK agreed the Change Order Credit was going to be $5761 for this job. Laurie is typing, we will delivery for signature to DK today and then they can get their closeout docs in order and we can close this job.

08/13/10 – Payapps updated for both contractors. As Built dwgs completed by TEI. QCM needs to sign the change order and send to Bert.

07/28/10 – Check on DK change order work and approve as-builts for project completion.

05/26/10 – DK is installing the fiberglass pan in June. We are awaiting QCM closeout docs.

05/15/10 – TTC is working on closeout documents for QCM. We have the final payapp but I don’t know if we have the O&M manuals so I contacted QCM again to check the status.

04/26/10 – Project is complete save the fiberglass pan issue.

03/23/10 – DK got an okay to provide pricing for a new fiberglass drip pan for the panel as we wait for THP to work on better site drainage.

02/22/10 – TTC signed additional payapps for QCM and DK for payment on this job. We are ready to close it out. There is an issue about a ground fault alarm and wiring on the fire pit. DK said they may just replace the wire and be done with it.

12/31/09 – Debra Kuempel did the overnight work on the conduit sealant upgrade and I believe we are down to final closeout. We have a warranty issued with the pit having a grounding issue at the fire alarm panel. DK has sent a technician to repair.

10/20/09 – TTC sent out punchlist today to contractors with instructions to closeout project. We are accepting CO#09 for blacktop from QCM. We have to work out the $10,000 allowance with Debra Kuempel and the $7500 with QCM for this project to write the final Change Orders. The email included instructions to closeout and finalize change orders and allowances.

09/22/09 – The project is nearly complete. TEI will be performing the punchlist in the next few weeks. DK is going to be on site on Thursday to finish up their work. Queen City Mechanical submitted change orders for processing today. We received a new COR #8 for piping. TEi asked for more info on this change order. We should be able to close out in a month.

08/26/09 – Final tie-ins for both fire protection and fire alarm are being made. By the end of this week, everything should be 99% complete and ready for punchlist.

07/27/09 – Bert and Terry talked last week about having TEI modifying our permit to list the Fire Pit as existing equipment. TTC called but the Inspectors office was closed. I will try back again later. The County has asked for the CWW to submit their drawings and apply for a permit for the work separately.

06/23/09 – All the permit issues have been resolved. TTC contact QCM mechanical by email and phone to get an update on status of fire main trench. No update but TTC will check on 06/24 when he visits the site for the FA project. DK has not been back to the site since the new power was brought back online. We believe they will return when they work on the trench with QCM.

05/22/09 – DK has finished the transformer relocation. The building is back up on Duke power from the new transformer. The cables are being collected for pickup. Prus provided the new fire main and fire pit for us. QCM will be out 05/26/09 to begin trenching the parking lot for the remaining work on fire alarm and fire sprinkler work.

04/03/09 – TTC sent an email update to the team for the pit, given it is located in the parking lot. Tony is going to look at the final location. Bert and Tony to advise. Luckily, we talked about this happening in the pre construction meeting.

02/19/09 – GCWW has brokered a deal to have Ray Prus provide the water tap to the R/W for QCM from the new water main. Ford Development will build the pit from that location. We have asked for a credit from QCM for the smaller scope.

02/0509 – QCM informed us that a new water main is being installed in Hamilton Ave. I advised QCM to investigate using that main instead of the one we originally projected for use.

02/02/09 – TTC modified the project schedule and sent out for approval to all the vendors. Jackson questioned the actual shutdown and switchover date but TTC told him that was not set in stone yet. The City has approved the water permit. We are waiting on ODOT approval of the plans before digging.

01/05/09 – Payapp #3 came in for DK and ThermalTech approved and sent to Bert. This payapp went to Bert first which may delay this payment.

10/16/08 – ThermalTech sent profile drawing to QCM for submission to state. TEI sent new permit drawings to Debra Kuempel for permit application. DK is getting price from Duke for work on this project per the scope of work on the specs as TEI wrote.

08/29/08 – Contracts approved. Pre Const Set for next Thursday.

08/22/08 – No update from County. TEI inquired about proceeding. Tony answered that contracts will be in front of the Board next Wed. 08/27.

08/08/08 – Tony reported the Prosecutor has contracts for approval. They are pushing hard.

07/31/08 – County reported that all contracts are at Prosecutor’s Office awaiting approval.

07/30/08 – TEI has delivered all contracts to Contractors for signing. TEI sent stamped permit drawings to Queen City Mechanicals for CWW permit application. QCM stated that CWW will want permit tap fee as soon as permit is awarded. This process may be the longest lead item of the whole job.

07/18/08 – Pre Award Meeting complete. All contractors recommended for project. Debra and Queen City Mechanical both stated they understood the project, the schedule and agreed their price was valid for this scope.



01/24/11 – TTC asked for a written letter from KFI to proceed with closeout as the punchlist is not fully complete as the scratches on the sign have to be touched up.

10/28/10 – We have visited the site several times as the foundation has been poured, electrical has been run and masonry work is ready to begin.

09/29/10 – Holding for State to complete road work.

08/13/10 – Payment 1 has been sent.

07/28/10 – TEI to approve payapp and send to Bert. It looks good, have to visit the site to see the sign.

04/26/10 – We are on hold until June.

03/23/10 – It has been reported that the sign is complete.

02/24/10 – John Brockman called with a question about backlighting the can. TEI will return the call.

02/05/10 – TTC visited Ernst signs and worked out final details on the sign building scheduled for April. TTC updated all parties.

12/31/09 – TTC submitted a permit answer to HamCo building inspections. The next step is theirs. We may have to resubmit to start the project again.

09/22/09 – The new street is very close to the sign. The County is working with ODOT to finalize the sign work. Our contractors are preparing a new shop drawing to the County permit office with the galvanized steel frame.

07/29/09 – No updates because Street Widening is still on hold.

06/23/09 – KF has not returned our call or emails about sign permit approval. At this time it is being held for review by Marty in HamCo Bldg Dept awaiting our clarification on how the sign post will be embedded into the sign base and what material it will be. We may need a meeting or conference call with KF and Brockman signs.

05/22/09 – Zoning has approved our sign permit. The bldg department is questioning the supports for the sign. Tom Speer is checking his design for us.

03/02/09 – Pre-Construction meeting held. KFI will order materials and store sign at HQ to allow invoicing. On hold until street widening is complete.


SHERIFF TARGET RANGE Match House Investigation

01/29/15 – County requested THP review the Match House walls.

02/23/15 – THP sent email with general recommendations and an opinion of probable cost.

02/24/15 – County requested a full letter report and authorized additional field work.

03/26/15 – THP anticipates publishing the final letter report.

03/26/15 – THP published final letter report.


PBS – North Park Project

03/02/05 – Submitted Change Order #1 deduct for $3000 for non used allowance.

10/15/04 – Final walkthrough performed. Waiting for final project closeout paperwork from Owensville. Punchlist contained zero items.

Construction is underway with Owensville is the Electrical Vendor for this project.


PBS – Sound System Repair Project

03/21/06 – TEI asked for an engineering change order which covers extra time spent by ThermalTech during the construction phase and occupancy permit time on this project. Joe said he signed it and went on to his boss Eric Stuckey. This project is officially complete.

02/13/06 – TTC finally approved the closeout documents, final pay applications # 6 & 7 and as-built drawings for this project. The only thing left is TE request for additional payment.

12/02/05 – TE requested a revised final payapp reflecting $6200 change order that was approved for the north end zone speakers. All final paperwork and as-builts have been received from Durrell in hard copy and CD. TTC made a copy of this information and is sending all originals to Joe Feldkamp for his use included 3 binders of O&M’s required by spec. TE will have have a hard copy of the 1/2 sized dwgs, just the CD.

11/30/05 – Chris Schaeffer called to inform Terry that the County reached settlement with Ed Simon and his lawyer’s and that a court appearance would not be needed.

11/28/05 – Upon final check the sound system is working, the new speakers are installed and working and the final paperwork has all arrived including the as-builts for this building. TTC has to write an engineering change order for ThermalTech’s Extra Time and WJHW Extra Time on this project and submit it for payment for the year 2005.

11/14/05 – Durrell and TJ’s total change order totaled $6200 ($5600 + $600). Joe Feldkamp approved. TTC confirmed with both contractors. Chris Horrell and Scott have scheduled Durrell’s visit for this Wed 1/16/05. Work to be completed by Thursday. TE shall send out the hard copy paperwork to all vendors.

11/10/05 – TTC was contacted by Chris Schaeffer about a possible court appearance in December over the bidding of this project. TTC discussed County bidding procedures and current ORC requirements for bidding and how the Sound Engineer, Audio Contractor and Electrical Contractor was selected for this project.

09/12/05 – Pay Application #4 is being resent to TE. TTC will approve and send to County for payment. TJ delivered payapp to TE today.

09/08/05 – TTC emailed contractors about closeout for project and punchlist. Sound was excellent last Friday when TTC went to the Colts-Bengals preseason game. TTC sat in section 315 and was impressed with quality of sound and clearness of PA announcer.

08/09/05 – Durrell reported that, Tom and I went over these items and will have most knocked out this week prior to this weekends game. What will be left will have to do with getting QSC down to make the cosmetic changes to the amplifier control system to make it that much more user friendly and correctly control the amplifiers that are being displayed.

08/08/05 – TTC coordinated the final change order amounts with TJ Williams and will modify the change orders for final signing today. Both change orders are complete. TTC will deliver them by hand to PBS tomorrow and try to get Bill Meyer to stop down and sign them. Punchlist was compiled and emailed out to contractors (hyperlink below).

08/01/05 – WJHW and Durrell completed the punchlisting work. Ron Baker is going to send the punchlist to TTC for distribution. TTC talked to Scott Simpson about items that he needed added from his end and left a message for Joe about his items. The consensus appears to be that everything is very nearly complete and everyone is very happy. Hopefully this will lead to a quick closure and a working sound system ready for the season in a very short time period.

07/19/05 – We have reached resolution on the PBS Sound Change Orders. Amy will approve change order #1 8855.00 for the countertops and change order #2: 2,060.74 for restocking fees from TJ. TE will write the new change orders for these amounts and send them to TJ for signatures and then onto JF for signatures by the County. WJHW and Durrell have scheduled the Engineering Commissioning for this project on 07/27 thru 07/29. TTC will be out of town on Wednesday but will assist on Thurs/Fri as needed. Punchlist should be fine as Durrell and Scott Simpson have been working hand in hand during the project. TE and WJHW will supervise and TE will try to set up and Const Mtg on Friday for a final update on progress before we start the as-built drawings for this job. Durrell reported by email on 07/18/05:

Last week Durrell got all the speakers working (with only two exceptions) and wired to the correct amplifiers. The West amp room is complete, all racks laced and cleaned, all amplifiers working. The East we ran out of time Friday, but we have all amplifiers working with the correct speakers, we need to spend one day lacing up and minor cleaning.

We found one 2″ driver not working on the East side that will have to be accessed and one crossover on a full range speaker on top of the north end zone – each will require one day’s work.

The west club area is finished and working and the east is about two hours away (got held up due to a wedding party).

The control room is up and working. We had a MediaMatrix card fail (used backup for tuning) and Scott Simpson has to provide us with info where to mount the two computer screens and the Volume meter, both minor installs. The console and all the other equipment is in a functioning correctly including the software upgrades. We want to rebuild the control block that exist for the control screens for QSControl – some coordination will need to exist between Scott Simpson, Tom Moores and we would like Ron Baker’s input – when the commissioning is done.

Durrell will schedule two days prior to WJHW’s coming to PBS for their commissioning to be sure all of these minor items are finished. We are waiting on Ron Baker for that date (date has been set see above)

07/08/05 – Eric Stuckey, Joe Feldkamp, Diana Craynon and Bill Meyer called today to inform me that the change order has be stopped and no work will be done on it and it is most likely canceled until later. We informed TJ Williams and TTC is supposed to cancel Monday’s field trip to work on it.

07/05/05 – TTC emailed Eric Stuckey the final Change Order #01 cost including the cabling and the countertop for approval. PDF. TTC met Chris Schaeffer, Scott Simpson, Steve Durr, Sam Kinman and Ed Simon and his associates downtown today to supervise a field visit and walk through with the County for the County lawsuit.

06/27/05 – TTC sent out CD’s to HamCo for law suit.

06/23/05 – TJ Williams sent the revised ACO out today. PDF. Terry and Joe asked for TJ to redo the change order to reelect the credit that the County will be receiving in place of the speaker parts that won’t be installed. TTC and JF will then talked to Amy about writing a change order to finish the project off with the cabling upgrade. The Const Mtg verified that many, many of the issues with the speakers were resolved by fixing wiring and termination points. ThermalTech believes that by knowing this it is a fair inference that the 3000 foot cable splices are probably not infallible.

06/22/05 – Const Mtg #02 was held. Durrell reported that all base work is complete except for software upgrade on MM. Durrell offered to do all the CO work except for the cabling for the $50,000 allowance. TTC agreed that this work be a fair barter. We decided to accept the countertop change order from jb schmitt. We will pursue a change order to get the cabling issue taken care of this year and Joe and TTC will talk to Amy about it after TJ provides an update T&M price NTE for the work.

06/16/05 – TTC visited the site earlier in the week with JB Schmitt and they turned in a quote today for the new countertop work at PBS. TTC forwarded the quote to the team for a decision. Countertop = $5695.00 and New Window = $3160.00.

06/06/05 – Durrell sent an email request Topper be added to all correspondences on this project. TE forward to Joe Feldkamp for approval. PDF.

05/31/05 – TE tried to reach WJHW for a conference call to discuss moving forward with additional bidding on Change Orders for this project. CO# 2 and CO# 4 are significant and may have to be bid out separately. TJ Williams called today to see about CO approvals. TTC sent the Allowance email to Bill Meyer. TTC should have sent last week when it was approved but somehow that didn’t happen. TE and WJHW still need to talk to Joe about Club Level work and scope.

05/24/05 – TE issued a recommendation letter to PBS about Change Orders for this project. TE spoke to Durrell and told them they must proceed with scheduling on the assumption that the Change Orders # 2 and # 4 are not approved. The County will issue a formal approval as soon as it can. This may lead to scheduling conflicts later near substantial completion but bidding and change order policies must be adhered to for this project. PDF

05/17/05 – Consultants and JF are in consultation about approving change order request. TE has requested direction from their sub WJHW on what is essential and what is not. First TJ payapp is approved and sent to JF.

05/11/05 – TJ submitted 5 change orders today to the project. PDF TE setup a construction meeting for Thursday at 1:30 p.m. to discuss project progress. TJ contracts are signed and in security for contractor to pick up.

04/27/05 – Contract has been approved by the Commissioners. TTC advised the contractors that the contract and PO would soon be in their hands. TE and WJHW will have subs complete by the end of the week.

04/26/05 – WJHW has submittals for AV work on the job. TE has reviewed the submittals fro EC compliance and found no errors so we will approve them with “No Exceptions Noted” to that end, we have told TJ the same. TJ inquired about contract approval and TE told them Wed at Board meeting, we should know. Please advise us when you hear information about Contract.

04/22/05 – TE received submittals for TJ yesterday. TE is checking for EC items. TE shipped one to WJHW for AV items. TE has delivered all dwgs and specs to TJ for permit application. JF reported that contract is in front of board for approval on 04/27/05. Everyone is still working at their own risk to date. Durrell has sound system working in bowl for weekend draft picks. County Prosecutor asked for some video footage of splice problems in the amp rooms. PBS reported that some have been fixed but other still remain to be documented and that Durrell may know best which ones to document. JF and TTC communicated about this pre-work in email PDF.

04/19/05 – TE sent out an email stating the Durrell may be doing some additional work in the Club levels that is above the scope of work outlined in the contract. Monies for this work shall be determined later.

04/18/05 – TJ sent their employee list into the County for this project. PDF. JF reported that contracts have been submitted and the project is up for Board approval on 04/27/05.

04/12/05 – TE had TJ resign new contracts forms and delivered to Kerry Dowling for approval. Kerry said that she would sign these on 04/11/05 and TE would have City Dash pick up the newly signed contracts and have them delivered back to Joe Feldkamp for submission to the County Commissioners for approval. TE sent 3 sets of dwgs to Dallas for WJHW to sign for TJ to turn into IBI for permit sets. TE and JF told TJ at the pre construct mtg on 04/11/05 that all work done by Durrell and TJ before receiving the contract and the PO is “at your own risk” type work.

02/23/05 – TE delivered contact to Joe Feldkamp. TTC and JF discussed contractor over phone and coordinated paperwork due to prosecutor. JF will submit contract for prosecutor review. TE will hold on notification to contractors until all paperwork is finalized. TTC would like to see 2 week turnaround in contract to keep scheduled intact.

02/22/05 – TE has received all required bid paperwork and finalized the Contract. Contract, Drawings, Spec Books and Drawings will be sent to Joe Feldkamp today.

02/16/05 – County and PBS indicated to proceed with Base Bid and Alternate 2 for this project.

02/16/05 – TTC and WJHW discussed the Alternates from the bids and are basically recommending that the County accept Alternate # 2 at this time and wait on Alternate # 1 because of the unknown variables included with updating the system. TE recommends fixing all the existing problems first before adding a new layer of master control that my contribute new and undiscovered problem of us to troubleshoot in the field plus upgrading to the BIAMP system will add additional dollars to the existing SOSS as this is a scope change for engineering (WJHW is working on an adder price). We will write the contract with Alternate #2 for the time being.

02/08/05 – Bids opened. TJ Williams appears to be the low bidder. TE has contacted contractor for additional paperwork and will begin contract work immediately. County needs to provide a signed bid evaluation form with Alternate recommendations for contract.

01/25/05 – Addendum 3 distributed. Durrell asked a new question for Addendum 4. The bid was extended to February 3, 2005.

01/21/05 – TE coordinated with WJHW and County to extend bid date one week and to send out Addendum 03 on Monday with updated information for bidding and alternate. Purchasing Department instructed TE to contact ICB Audio via email about non-inclusion into the accepted bidder’s list for this project.

01/20/05 – TE distributed Addendum 01 on 01/11/05 and Addendum 02 on 01/20/05. Addendum 03 should come out on 01/21/05. TE has written a rejection letter and given it to Purchasing to give to ICB as they do not meet the Professional Stadium Experience Requirement on this project. The SoundCom qualifications have been sent to WJHW for review and we are currently waiting on SoundCom to confirm a couple of questions about which office will serve this project but they most likely will be approved for bidding as they do meet the quals. TTC asked if the bid should be extended since SoundCom is being added. TE is awaiting the County’s decision. The second prebid was attended by about 8 contractors. Muzak declined to bid and left the prebid after we stated that OEM equipment must be used for repair.

01/19/05 – Prebid is finished. Addendum # 01 wads distributed. Addendum # 02 is going out tonight. Addendum # 03 is scheduled out tomorrow. ICB will not be an approved vendor. TE has to provide a letter to Purchasing for the use, modify and send.

01/14/05 – Joe Feldkamp called to inform TE that because the prosecutor wants a firm number for cost to provide as-built drawings that we can change the bid for to add an alternate to have the AV contractor provide a complete set of as-builts encompassing the entire scope of work at the stadium. TE shall modify the bid form accordingly.

12/24/24 – TE invited Audio Specialist Contractors. Project Manual is on the street, on schedule. Prebid set for January 5th. WJHW will be in town for the prebid. Everything appears to be set for bidding att his time.

11/04/04 – TE found and sent the AV drawings to WJHW today. Final design is still scheduled for final delivery on 12/08/04.

10/25/04 – WJHW sent an informal report from the field trip last week. PDF

10/19/04 – TE, WJHW, PBS and HamCo had a meeting and set the following schedule for this project (see below)

10/19/04 – TE and WJHW met onsite with the owners and ironed out the scope of the project, schedule of the project and the expectations of the project. WJHW will begin design and target first week of December for final engineering documents. TE will release minutes in a follow up email. Contract will be bid to a local EC with a Audio Specialist assigned as a sub.

10/15/04 – TE will meet WJHW onsite on Tuesday for kickoff meeting. WJHW will be onsite late Monday afternoon. They have coordinated their time with Scott Simpson.

ThermalTech solicited consultant bids and wants to team with WHJW from Texas for this project.

Conference call on August 17 between Eric Brown, Bob Bedinghaus, Joe Feldkamp and Terry Cannon with WHJW was held to discuss itemized consultant cost and scope of consultant work on this project.

TE is awaiting a signed SOSS from the County.

The County is awaiting a one page report from WHJW outlining the added benefits of using WHJW on top of the Durrell report.

Project scheduled to be signed in August 04, Designed in Sept – Oct 04, Bid Nov 04, Awarded in Dec 04, Construction Jan 05.


PBS – Occupancy Redistribution Project

03/20/06 – Joe approved TEI’s cost on the ECO for Sound project. This project is officially complete and ready to be archived.

11/30/05 – After speaking with City Permit office they confirmed that TE did pay for these permits. TTC called Randy this morning and told him that these permits were paid for on 10/18/05. TTC is awaiting a call back to clear this up.

11/25/05 – Randy Campion called to inform TE that there is a remaining balance on these permits. TTC called the city to see what that cost was because TE has already paid for the new permits with a check. Once this is cleared up Randy will send new Certificates of Occupancy.

10/18/05 – All revisions have been accepted and approved by the City. New permits costs = $354.00. New occupancies have been set to 1021 (East Club) and 986 (West Club). According to Randy Campion once the permit is paid for, which TTC did today, the permit will close and information will be sent to Fire Dept. They will issue a new occupancy for private club events. to PBS.

09/28/05 – The City received all the information from Cole & Russell but need it properly inserted into the permit dwg sets. TTC will stop at the permit office and manually do this on Friday or earlier if possible.

09/20/05 – After the City sent the letter emailed by TTC to PBS yesterday requesting additional egress information, TTC talked to Cole & Russell about this information. C&R believes that they already have all that information collected from the sliding door project and will call Carl Graeber and discuss directly with him. They can get all that information to the City on Friday or early next week. TTC will check back with city and C&R about progress at that time. We believe the city will accept this and process the permit app immediately thereafter for final inspection.

09/08/05 – TTC contacted city. Permit should be reviewed by 09/13/05. TTC will check back at that time.

08/25/05 – The permit is still being reviewed by the City. TTC will continue to check on status.

0815/05 – TTC met Randy Campion at the PBS site on Monday, Randy approved the 821 and 851 increased occupant load and will sign the CO paperwork and send it to PBS. After the meeting, PBS said that the 30% OA was not a problem so that they want even higher occupant loads. TTC turned in another set of prints with the higher numbers of 986 and 1021. The city made us turn in for all new permits: 2005P06669 and 2005P06670.

07/18/05 – City called with Approved Permit numbers today. Permit was split up into 2 permits 2005P05544 AND 2005PO5589 ($284). TTC will pick-up permit on Tuesday morning. Inspector, Randy Campion has to come to the site to approve and give us a new CO for the space. This will be coordinate with HamCo & PBS in the next couple of weeks.

06/30/05 – TTC got GD to sign occupancy permit drawings in regards to HVAC OA requirements. TTC submitted for permit. Permit # 2005P05544. TTC informed all parties that permit can be checked at www.cincinnati-oh-gov at anytime. Amit was also promoted to Project Manager, so we are unclear if he will review this permit.

06/16/05 – Eric called today and TTC reported that Amit had returned one of his phone calls stating that he was seriously behind in permit reviews. With that knowledge, TTC stated that we should just go ahead and subject a revised CO application and talk to him them as opposed to conferencing in with him early.

06/06/05 – TTC sent an email outlining TE’s field findings to Joe, Sam and Eric. TE need direction from PBS to proceed on next level or work.

06/03/05 – Sam Boehringer of TE collected field data to update actual occupancies of the space.

05/25/05 – Eric Brown setup a meeting with the CFD department about occupancy requirements of the club level space.


PBS – Deep Well Pump Replacement

03/23/06 – TTC spoke to John Wagner of Moody’s today and here is where the problem is: the actual repairs cost $15,573 not $6573, the 6573 +the 9000 already paid = the total cost of repairs on this job. This means that they are still giving you back $4427 on the job. John outlined this in a letter in September which I am having faxed to me. With this information, I will redo the paperwork and send it up to John for use on completing this project. I have to write a new change order, allowance usage and final pay application to make all the numbers work out right.

03/20/06 – TEI received final paperwork on this project today from Moody’s. This will be forwarded to Joe for payment on this project. TEI to begin phase II service contract immediately.

02/16/06 – Moody contacted TEI to request AIA forms. TTC sent them out in US mail for their use. They should complete these, notarize them and get them back for closeout.

02/06/06 – After an inquiry by Sam Kinman, TTC found the payapps sent to ThermalTech for approval. They were approved by Sam on 11/21/05. The problem is that they are not on AIA invoices which means there is no where for TTC approve. Assuming they are approved by TTC on these forms, the Contractor still has not turned in standard county documents for project closeout. TTC will forward a copy to Moody’s and ask for return of these documents. TTC completed a full packet of information required and mailed to Moody’s today via US Mail. TTC will follow up with a phone call today as well.

11/30/05 – Sam Kinman reported today in an email back to TTC that the deep well pump replacement job is complete. Since TE wasn’t used very much during construction, TTC assumes the punchlist is fine and no additional work needs to be done. TTC will archive the project from the database and closeout the project number. TE has never received a payment to approve, nor the final County paperwork to approve for that payment, nor manage any of the allowance monies. All this will need to be done before the project is closed out in the County’s books.

08/08/05 – TTC contacted the County about status of the Deep Well Pump Contracts. The previous entry does not state that the contracts were sent to PBS.

07/18/05 – Bid were opened. Moody was apparent low bidder at about $29,400. $20,000 in Alternates can be taken (ALT #1 and ALT #3). TE has contract written for $49,400 and will give 4 original copies to PBS for approval with Board. Bid PDF

07/05/05 – TTC worked on July 4 to get Addendum 2 complete and sent it out for distribution. PDF

06/27/05 – TE worked on Add01 and it will go out to the Contractors on 06/28/05. PDF

06/13/05 – TE sent out Project Manual for bidding. PDF

06/06/05 – Comments received from Sam and John Wagner of Moody in Dayton. TTC will update Project Manual and prepare for bidding ASAP. PBS talked about including a 3-year service contract on top of the bid. TE has done these in the past and can add this scope to the project if so directed by the County for bidding. This would be an engineering scope change to this project but it makes very good sense to include it now (12 hours x $80 = $960) and that includes the contract modifications for the Alternate bid to take optional years.

06/01/05 – Comment received from Joe and Eric. Nothing from Sam yet.

05/30/05 – TTC sent a project manual out for review over the weekend to PBS and County. Once comments are received, TE will distribute to public for bidding. All dates must still be coordinated with HamCo Purchasing.

05/26/05 – TTC contacted Moody’s for some specific questions about parts replacement. The Project Manual is built for front end. TTC needs to add dwgs and mechanical section. Bid Form needs to be modified for this job.

05/17/05 – The team had serious discussions about project schedule today. PBS re-iterated their concern about flooding. TE provided a timetable for construction that appeared not to be acceptable to PBS. TE had discussions with JF about project and will try to speed up project manual creation significantly to lessen risk. It is TE’s belief that County Purchasing has already decreed that this project does not meet Emergency Status.


PBS – Pro Shop Metering Project

03/29/06 – TTC reviewed the billed KW and KWH to the PBS Pro Shop and the numbers looked good. The pro shop is using less than 1% of the total power consumed at the building.


50.91 21,863

34.91 28,950

79.91 25,448

59.28 16,077

72.30 14,661

63.67 16,528

53.30 4,166
02/01/06 – TTC check on the meters and to review monthly billings at the site. The meters appear to be reading in the Megawatts which seems really high at first glance. TTC to reconcile with actual building data.

11/16/05 – TA visited the site today. Reviewed punchlist and determined that Bond Electric has not completed the issues described on the punchlist. TE called Bond to tell that them that can’t get paid until these items and the final paperwork is complete.

11/15/05 – As-Built drawings for this job were turned over to Wayne Trent of HamCo Facilities today.

06/17/05 – Const Mtg was held at PBS. Sam did not attend. Equipment is not on site yet but Bond is ready to get to work as soon as it arrives.

02/28/05 – Purchasing called to inform TE that Bond does NOT have a performance bond or check submitted for this project. We should not let them begin work without this paperwork in place. TA to confer with Jack Boland today and emphasize the need for this check immediately.

02/10/05 – Contracts and Bid Results are in County’s hands. TE awaiting further direction on this project.

11/17/04 – Bid Results

10/12/04 – Project is out for bidding.

Addendum#1 distributed.

Addendum#2 distributed.

Addendum#3 distributed


PBS – Deep Well Pump Service Contract

02/02/07 – After many contract revisions, phone calls and problems, the Prosecutor finally allowed preventative only maintenance contract that Kerry wrote and gave back to everyone to use. Other repairs and emergency repairs must be bid our as prevailing wages. TEI does not and will not every agree with this approach. We took prices for many other services to save the County money but those were not accepted as part of this job by Hamco Prosecutors office.

01/10/07 – TTC got Kerry’s new email address and forwarded the revised contract to her for any final changes that she would like to make. Kerry confirmed that the packet was received at her office.

01/08/07 – TTC hand delivered the final contract with Kerry’s marks to security for Jenny and Joe to pick up and resubmit to the Prosecutor’s Office.

12/26/06 – TTC will finalize the deep well pump contract this week.

03/29/06 – TEI has not heard from Daktronics with a registered service repair list yet so I called and got Greg’s email and sent him the request again.

03/21/06 – TEI finished the SOSS and will drop it off at the HamCo Facility Department on Wednesday for this project.


PBS – Energy Study

03/20/07 – Project ready to archive.

11/30/05 – TTC and GD have to meet internally to review the 10/11/05 findings and energy usage. GD believes there can be savings found in schedules and space control but we will need “buy in” from facility department to implement. TTC has to setup meeting with GD to review other items but County end of year crunch on projects has dampened progress. All work will be done by end of year.

10/11/05 – TTC and GD met on site and reviewed the whole facility with Sam, Joe, Eric and the financial manager.

05/10/05 – Kickoff meeting for project with JF, TTC.


PBS – Video Wall Replacement Project

12/19/07 – Nor Comm has provided the AIA closeout paperwork, warranty letter and payapps. TTC signed and City Dashed it to Joe Feldkamp. This project is complete.

09/03/07 – TTC has received and approved the O&M’s. TTC is hand delivering the box of O&M’s and final payapp to PBS tomorrow at 9am at the Deep Well Pump prebid.

08/28/07 – TTC contacted Nor-Com again about final payment via email with a Read Receipt.

08/20/07 – Nor-Com contacted Joe today about final payment. TTC emailed Joe back station what happened to Payapp #2 as listed below. To date, we have not received any additional paperwork.

07/02/07 – TTC has contacted Howard Milkis and Nor-Com several times for closeout on this project. No response to TEI. TTC will try to contact the office directly today and find out how we are going to close out this project.

05/14/07 – PBS did not use TEI service much at all on site during this project. TEI has been in contact with Joe and Scott numerous times during construction as things were going well. TEI contacted owner today for punchlist and closeout items for Nor Comm. As far as TEI knows, this project is complete.

03/30/07 – TE received submittals for review and promised to work on them this week. Nor-Com has a schedule forthcoming.

03/16/07 – Scott and Joe held a kickoff meeting. TTC to follow up and publish some const mtg notes this week. A second Const Mtg is to be scheduled. PBS would like TV in before draft day.

02/13/07 – All contracts and information is in the Prosecutors office for review. Comments have been added and the contract should be approved very soon. TEI provided a copy of the bid results to Kerry Dowling per her request last week. Joe Feldkamp is following up with the Prosecutor Office.

01/30/07 – TEI finished the revised contract and returned it to PBS for use on this project. Now TEI is waiting for approval and signatures to move into a pre construction meeting.

01/10/07 – TTC hand delivered the contract based on the Base Bid cost to Joe for submission to the Prosecutor’s office. TTC included a $2500 allowance in the project after the bids were opened.

12/26/06 – TTC will finalize the deep well pump contract this week.


PBS Video Switcher

10/26/09 – Final paperwork is received. TTC approved and mailed to Joe.

09/22/09 – Project is complete. We have approved payapp #2. Once all the closeout paperwork has been received and training is complete. We can close out this project.

08/25/09 – One payapp has been approved. Punchlist has started. Training continues.

07/29/09 – Project has been bid and awarded to Rosco. Work is proceeding and Chris Williams will be able to check in on progress on August 4th when he visits Cincinnati for the Brewers game.

06/23/09 – This project is in contract approval stage at HamCo.


PBS Z-Ramp Replacement

09/14/09 – Additional documentation has been sent to Kerry for contract approval.

08/25/09 – Joe needs additional bid information for the project. United needs to provide final bid paperwork to Joe.

0729/09 – Project has been bid. Prices were tabulated and sent to bidders. Final contract was $16,300 and included $500 in allowance monies. TEI has written the contract and given to the County for signing.

06/23/09 – This project is out for small project bidding.


Parkhaus Garage Assessment

September 7, 2011 THP sent assessment report to County

THP anticipates completing the report by August 5.

July 25, 2011 THP completed all field survey.

Assessment placed on hold until Ram completes stair tower work.

THP anticipates field work in next two weeks and report to be complete by June 30.

April 26, 2011 SOSS approved.

County reviewing SOSS.

April 18, 2011 THP provided requested SOSS to complete garage assessment for the Parkha


Northside MR/DD Masonry Review

THP closed out project.

October 29, 2010 THP provided a summary letter of masonry work.

THP anticipates publishing a letter to GBBN documenting field condition on October 29, 2010.

October 14, 2010 THP notified GBBN that lintel repairs were not recommend and THP will follow-up with a letter documenting the field visit.

September 29, 2010 the County, GBBN and THP met to review the lintel condition of the west wall of the Northside MR/DD facility.


DDS-Beckman Adult Center Change of Use Remediation Project.

This is second SOSS for overall project. This second SOSS is to await results of complete code investigation and response from Colerain Fire Department to determine complete scope of work for project.

This project was determined to have ThermalTech as the lead design professional. GBBN is a consultant only.

5/17/10- GBBN received final Structural Analysis for Beckman Change of Use Code Review.

5/25/10- Arcodect delivered final Code Review for implementation into Permit Set of Construction Documents.

5/26/10- Final Permit Set of Construction Documents printed, signed and stamped. Set to be hand delivered to TEI at the Partnering Meeting for issuance to the Hamilton County Planning Department.

6/4/10- Issuance of ITB#047-10 by Hamilton County Purchasing Department.

6/14/10- Attend Pre-bid Meeting at Beckman Adult Center.

6/15/10- Issuance of Addenda #1 with Pre-bid meeting minutes.

6/23/10- Bid question sent directly to GBBN. Bidder instructed to send question to Purchasing per the bidding documents.

7/1/10- Building Department Response Letter received from TEI with two items concerning occupany separations. Building Department requesting drawing showing Use Group separations, net/gross areas and occupancy locations on floor plan.

7/02/10- Responses for Building Department Response Letter received from Wayne Meyer at Arcodect.

7/19/10- Newly created Occupancy Summary Drawings for Building Department Response Letter delivered to TEI.

9/15/10- Construction meeting was held.

9/27/10- TEI announced the Plans Examination was completed and approved for Building Permit by Hamilton County.

1/27/11- TEI and Kevin Maue confirmed the permit is now closed on this project. The Certificate of Occupancy and the Colerain Township Fire Department Letter of Release have been obtained.

1/27/11- Kevin Maue indicated at the Partnering meeting this project can be archived.


DDS-Beckman Adult Center Change of Use Code Investigation

3/19/10-Met with Jim Bowman, Colerain Fire Department, and Kevin Maue at Beckman Adult Center for walk-through on current renovation project.

3/20/10-Kevin Maue requested SOSS be submitted. This is first of two SOSS’s for this overall project.

3/21/10-Preliminary Code Investigation performed and submitted to Kevin Maue by email.

3/23/01-Fee proposal from Arcodect (code consultant) received. GBBN spoke to Cindy Rutherford, Hamilton County Building Department to discuss Change of Use requirements and submissions.

3/31/10-Schedule of submissions to Hamilton County Building Department submitted by email to Kevin Maue.

4/1/10-SOSS to be revised and re-submitted showing THP as a structural consultant to GBBN for any seismic implications per directions of Kevin Maue at Partnering Meeting.

4/7/10- GBBN met with Kevin maue, Terry Cannon (TEI) and Tom Frimming of Communale Fire Protection to review proposed AFSS implementation.

4/14/10- Signed SOSS without Ralph Linne’s signature and approval delivered by Kevin Maue to GBBN.

4/27/10- Signed SOSS with Ralph Linne’s signature and approval delivered by Bert Watts via email to GBBN.

4/27/10- Bryan Korb of GBBN assigned to project. Bryan to meet with Wayne Meyer code consultant at Beckman on Thursday and Friday for site investigation

4/29/10- GBBN to meet with owner, Colerain Fire Department, and Hamilton County Building Department to review building and change of use scope.

5/12/10- Kevin Maue announced ThermalTech Engineering would be the lead design professional on this project and GBBN would be the consultant. GBBN established estimated hours for completion to TEI for SOSS.


DDS Signage Refacing

10/01/09 – Kevin Maue of MR/DD contacted GBBN about the pending name change as mandated by the State of Ohio. Sign letters on 1520 Madison to be removed as part of ceremony and will be handled as a purchase order. The signage refacing of all other sites will be on a SOSS through the Partnership contract.

10/05/09- ASI Modulex was approved to handle the purchase order work for 1520 Madison based on a submitted quote to MR/DD.

10/07/09-Existing electronic drawings sent to ASI for signage permit with City of Cincinnati.

10/23/09- Shop drawings for 1520 Madison approved for layout and lettering style.

10/30/09- Signage Refacing RFP sent to four selected bidders.

11/24/09- Adams Signs was selected for Signage Refacing and P.O. 536683 released.

12/08/09-1520 Madison sign permit approved. ASI
to order building sign letters.

1/04/10-Adams Signs re-submittal shop drawings approved.

2/19/10-Adams Signs has completed all installations. Kemper Meadow Road sign has bleed through of old signage.

2/22/10-Email sent to Rick Cain of Adams Signs about corrective measures at Kemper Meadow Road location.

2/25/10-Received signed SOSS from Bert Watts at monthly partnering meeting.

3/15/10-Resent email to Rick Cain with photos of Kemper Meadow. Rick claimed the corrections would be completed soon.

4/3/10- Email sent to Rick Cain of Adams Signs about corrective measures at Kemper Meadow Road location.

4/10/10- Rick Cain called to indicate Adams Signs very busy but will try and expedite completion of remediation measures to Meadow Road location.

5/19/10- Kevin Maue sent an email to Adams Signs indicating Application for Payment non-payment until corrective work is accomplished at Meadow Road location.

6/18/10- Notified Adams Signs to submit warranty information to comply with close-out documentation.

6/23/10- Sent Invoice from Adams Signs to Kevin Maue.

6/24/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

6/29/10- Rick Cain called to indicated work is commencing to author standard warranty details.

7/28/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

7/30/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

8/24/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

9/27/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

10/25/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

11/25/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

12/14/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment. Rick Cains email was returned as undeliverable.


MRDD Franks Center Renovations

04/15/09 – CR Architecture + Design awaiting final punchlist items to be completed and closeout documents.

01/30/09 – Substantial completion issued.

01/30/09 – Final punchlist completed.

10/29/08 – Phase I punchlist.

09/22/08 – Additional service issued for Kitchen Counters.



02/14/08 – Punch list submitted.

02/13/08 – CRA sent additional service signed 2/13/08 to Bert

02/13/08 – TEI attended the const mtg today for MRDD. EDI is almost complete. Additional concrete will be installed around the new air handlers.

12/07/07 – CRA job meeting held on 12/05/07

11/14/07 – CRA received and approved shop drawings and substitution request for automatic door operators.

11/10/07 – The construction has commenced.

11/01/07 – Pre-construction meeting was held. See Meeting Minutes below.

10/26/07 – GC/EC submitted revised insurance certificate. GC/EC contracts have been approved approved. HVAC contract is still pending.

09/28/07 – Awaiting GC/EC to submit revised insurance certificate for final executed Contract.

08/16/07 – Contracts submitted 07/31, CRA received comments, re-issued 08/16.

08/02/07 – C+RA and TE submitted the Building Permit Application to the Village of Evendale Building Department.

06/25/07 – Addendum issued. Bids are due 07/02/07.

06/15/07 – A pre-bid meeting was held on location yesterday, June 14, 2007. This project is set to open for bid on June 26, 2007.

06/08/07 – This project has been sent to Cole & Russell for review. TE is awaiting comments back.

04/27/04 – Final review meeting is scheduled for next week.

04/20/07 – Updated drawings are being sent out today, April 20 2007.

04/13/07 – Project is progressing as expected.

04/06/07 – Preliminary set of drawings were sent to Architect today, April 6, 2007. Awaiting review comments.

03/30/07 – Final field work has been assessed. TE working on design.

03/23/07 – Project is progressing as expected.

03/20/07 – Kickoff meeting complete. Check CRA for updates.



02/15/08 – Add service CR-077a for $6,755.00 signed by MRDD on 2/12/08, forwarded to Bert Watts.

01/16/08 – Mtg held re additional service request my MRDD.

01/17/08 – Meeting held at site for scope definition. TJ and MAH attended. A few deletions and additions from original scope. ZZ ran meeting and will produce meeting notes. TJ and MAH will produce notes and forward to ZZ.

12/20/07 – Add service requested by MT

12/07/07 – CRA Told by MT that project will be from 2008 budget.

10/10/07 – Received signed SOSS.

09/06/07 – SOSS submitted.

05/18/07 – Kickoff meeting was held.


MRDD – Franks Center Renovations

03/31/08 – Bids due 4/11/08

03/31/08 – Pre-bid meeting

03/21/08 – Addendum 1 issued

03/08/08 – Issued for bid

02/15/08 – CRA Add Service CR-084a $6,999.00 signed by MRDD on 2/12/08, forwarded to Bert Watts.

01/11/08 – SOCC CR-084a modified per WM email 1/11/08

12/20/07 – CRA add service requested by MT 12/20/07

11/09/07 – SD review meeting was held at Franks Adult Center. CRA presented Schematic Design drawings and color and finishes samples.

11/02/07 – Next review meeting is scheduled for 11/08/07 at Franks Adult Center at 3pm.

10/24/07 – Meeting was held at Franks Adult Center to discuss the scope of the project and verify existing conditions.

09/24/07 – CRA received signed and approved SOSS.

08/30/07 – CRA submitted the SOSS.


MRDD – Storage Buildings

01/02/08 – Final warranty info received 1/2/08, forwarded to MT

11/26/07 – Received As-Built drawings, forwarded to MT on 11/26/07.

10/01/07 – Awaiting closeout documents. Need As-Built drawings.

08/15/07 – CO #1 and Pay Application #3 sent to Michael Trabue.

07/27/07 – Construction is complete. Pay Application #2 sent to Michael Trabue.



08/29/08 – Triton is still working on warranty issues. Boiler control boards. Humidity issues. RCP problems.

02/25/08 – Triton and JCI have been working to remedy the remaining warranty issues on this project which include boiler flame outs, VAV box control, min airflows through VAV boxes, winter heating. TEI has to look at cold restrooms in the 1956 building for Kevin Maue.

12/17/07 – JCI and Triton had to go back to the site today to continue to work on DDC issues since the RTU-2 is still not running properly. Triton put the unit in had on Friday at 40%. TEI does not have an update on Monday from the troubleshooting field trip.

12/14/07 – TTC was on site today with JCI and Carrier going over warranty issues on the job including the LON talk interface between RTU-2 and DDC. Scott fixed the misnomers on the RTU-1 unit in the points list. Scott fixed the economizer program so RTU-1 would actually run in economizer mode when conditions were right (there was a programming glitch in the system) and he addressed the setback temperatures which are correct as programmed. Training will be schedule the second week of January 2008 with MRDD. There was a cold call in the lobby today from Wayne Mckee and Triton and JCI was dispatched to investigate the problems with the space. The hot water is 156F so there should be plenty of hot water available for the system.

12/04/07 – CDFP is scheduled the FM200 training directly with the MRDD folks and they will keep TTC in the loop. CDFP sent a change order in today for additional services on this project that TTC forwarded to ZZ for approval on this project for the extra sprinkler work, additional fan test.

11/28/07 – Jim Lape emailed to ask about closeout on this job. TEI stated that TJ should be 100% complete on this project. Triton is also done but we don’t have final closeout documents from them on this job. TTC is trying to get Grady to give us all closeout docs for this job.

10/17/07 – TTC visited the site and produced some more items for JCI and Triton to take care of on this job. He emailed pics and instructions to the contractors for completion. Everyone is working on completing all the work this week. Triton says the final balance reports are complete. TTC said today that he would adjust the manual high limit controllers to 210F and let the factory 190F shut the boilers down on auto reset. TTC also said he would caulk up the hole in the concrete hatch patch while he was on site. TTC will also check the installation of the electric heaters in the stairwell today while he is on site.

10/08/07 – TTC answered Jim Lape’s email about door control hardware. On drawing E400 #12 & #13 state that TJ will wire up the devices furnished & installed by the GC.

10/01/07 – TTC was on site last week performing commissioning. The control system was not ready to be back checked by TEI. JCI had not performed the functional test and TTC did not have the CX sheets. Nevertheless the system is running and with another week of actual work, it will be fine. TTC requested trend reports for verification of working systems.

09/19/07 – TTC approved JMW change order for roof drain and sent email to Jim, Zach and Kevin to that effect. TTC reviewed Triton invoice and asked for change order modifications before actual billing to the County. Change Orders were not reflected on actual payapp but they were on the schedule of values.

09/17/07 – Discussed Change Orders # 09,10,11 with Tim and Kevin Maue. I approved them as submitted. The steel issue is still are large issue that doesn’t appear close to resolution. TTC responded by email about controlling the pressure of the system by setting the static residual to 5 psig and letting the building set the pressure from there. If we have to change the relief valves to 65 psig from there, we can. The static of the building on the basement pumps should be 26.64 psig (50 ft bldg ht / .4328 ft h20/ psi + 5 psig to keep air out).

09/12/07 – Triton is on site checking the existing VAV boxes for repairs. They found one bad damper and believe 6 of the 7 boxes that are not operating properly. We are concerned that we have paid for some of this work already at the beginning of the job, but we will pay for the repairs to the dampers now and fix the electric heaters.

09/04/07 – JCI has given training but the controls are not complete and they still need some work. Boiler are not operational at this time and Denny Lammers is still need to program the internals of the boilers so JCI can communicate and give them their firings rates and monitor alarms. TTC dropped off TJ Williams hard copy as-builts drawings for the project that need to be included into the O&M’s that Kevin already has in his office.

08/23/07 – TEI printed out 2 more originals for Change Order #1, signed them, and sent them to T.J. Williams with instruction to sign them and forward them to HCFD.

08/17/07 – TEI received as-built drawings from TJ. Williams today. TEI will update the electronic drawings with the information and forward them back to CRA early next week. TTC delivered the O&M’s for EC, PC and FPC to Jim Lape at his office today.

08/16/07 – TTC and Cintas performed and passed the final server room fire alarm test on Wednesday. TTC has the paperwork in the folder and will include in the O&M. Triton said that the handrail was ordered and will be put back by Pat Murphy.

08/06/07 – TJ will do the security camera and emergency lighting change order work at T&M not to exceed but they will not come down in price. TTC sent this info to Jim Lape and approved on his end. TEI needs direction from the owner. TTC contacted MRDD about installing gaskets behind wall plates in server room.

08/02/07 – TTC talked to Kevin today about TJ outstanding issues at the MRDD. Security Camera change order and Emergency Lighting change order. JMW has O&M’s into TEI and they all closeout paperwork into Kevin Maue. TEI received TJ Williams closeout paperwork on this project as well. NV had TTC give to it to NR for approval.

07/23/07 – TTC spoke to Tim Myers today about project closeout, change orders, punchlist and controls. TTC discussed the allowance usage with Tim which totaled $13459 to date not including the refrigerant changeout which is a different cost. TTC emailed CRA about distributing the correct paperwork for change order pricing from Triton on this job and listed 6 items that need to be priced for this project.

1. Extra Diffuser for 1956 building. 17 in Qty. Linear Slots and 2×2’s to meet job requirements.
2. Refrigerant Check, Removal, Replacement on RTU-1. This work wasn’t included in the allowance cost but we authorized Triton to do the work on a separate ticket.
3. Overtime for IT closets and occupation over one weekend (i believe KM authorized triton to do this work)
4. Replacement cost for handrail? Triton removed it but someone else threw it in the garbage. Hobert is supposedly looking for it today but Triton will have to replace it if it can’t be found.
5. Demolition of all housekeeping pads in boiler room. This was Empire work but Triton had to do it all because Empire refused to do it. If Triton is going to pay the steel cost as a change order to Empire which I now believe they will after speaking with Tim Myers this morning, Empire either has to reimburse them this cost or deduct from the steel change order cost. Tim brought up a good point. A $2029 redesign fee is steep. That is 20 hours of a ThermalTech Principal time or 40 hours of a contractors time. It did not take that long. Redrawing that steel took about 6 to 8 hours. That fee should from the sub-contractor $500-$700. Please try to get them to give us a “real” cost and the housekeeping pad back charge could go away. As a side note, the new designed decreased the amount of steel and grating required, why are they paying an upcharge of $2522.97?
6. Replacement of 2 electric heaters. Provide 2 recessed ceiling heaters in the 1956 stairwell in lieu of the scheduled heaters. Return the surface mounted fin heater and the wall heater.
07/17/07 – TTC has scheduled a backcheck of the plumbing punchlist on Thursday. He will meet Jason on site for the walk through. Triton has committed to Approved Products being on site to finish the balance portion of the work starting on 07/23. TTC sent 2 sketches to CRA for plumbing ASI’s.

07/03/07 – The job is mostly complete aside from outstanding issues. The MRDD people are in and using the building. Items that we are still working on include: balancing HVAC & water, relocating a fp line out of the elevator shaft, final test of FM200, hot water boiler loop, controls, roof drain near unit, and punchlist by all contractors.

06/26/07 – All TEI punchlist were generated last Friday. TEI gave Jim Lape a complete set of extra materials required on this project to included in substantial completion for the project. TTC was onsite with JMW plumbing early Monday morning to discuss punchlist items which they are addressing. TTC spoke with Kevin about roof drain issue on 1956 building and we agreed that changing the scupper might be the best option. We can drain it down to the canopy roof.

06/15/07 – TE will be generating a punch list next week.

06/07/07 – RTU is still surging. we are trying to find out the status of the jci & carrier contractors. JMW plumbing will be on site on Friday to demo the 4″ plumbing vent from the elevator room. we believe fire protection is complete and ready for testing.

06/04/07 – Inspections are scheduled for Tuesday 06/05. MRDD personnel are moving equipment into the building today. ttc did not approve payapp #6 for triton and asked them to back down the % on the ground floor of the 1956 bldg to 90% and controls down a bit lower.

05/30/07 – TEI was on site on Tuesday for our weekly walk through. TEI contacted the vendors with questions, TCO milestones and coordination issues at the MRDD to help get it ready for the MRDD move on June 4. Items include ceiling layout, diffuser locations, ductwork insulation, EF flashing, fire protection work, elevator details, sump work, controls, etc.

05/21/07 – Triton reported that all fire dampers were completed over the weekend and that mason would be finishing up wall penetrations today and tomorrow. Kevin Maue seemed unimpressed with the quality of the work being done over the weekend and hopes they pass inspection. Tim, Terry and Kevin will walk the site on Tuesday AM to check progress and get commitments for completion.

05/17/07 – Triton will not be working this weekend but will have another crew on the MRDD job starting on Monday and they will have masonry sub helping with the patching of the duct shaft. The new RTU will not be ready to be ran this weekend.

05/14/07 – A lot has happened since my last entry. We have attended weekly const mtgs. We have attended several additional field trips for construction issues. The 1992 building has reached occupancy. The 2nd floor move was today. The 1992 RTU-1 unit is up and running but not under full control. The FM200 system is not fully tested yet even though the room is built. The 1956 building has ceiling grid in the entire building and now diffusers and fire protection piping is being installed.

04/17/07 – TTC and TA walked the project with all the contractors in lieu of a const mtg. The RTU is on the roof. The Server room is nearly complete. Cintas is working to complete the testing procedure and set it up for us. Shaft wall is offset form fire dampers. TE approved shifting the wall over to catch the sleeve of the fire dampers. The wall holes in the return shaft are open and Triton asked about filling them in (even the ones that weren’t done by them). TE directed Triton that they could sleeve the chimney with the boiler vents. The elevator room exhaust fan will reuse one of the openings that goes up to the chimney. A gooseneck may have to be added to the top of the chimney to keep the exhaust 10ft away from our unit intake.

04/13/07 – Project progressing as expected.

04/06/07 – Project progressing as expected.

03/30/07 – Project progressing as expected. Triton reported that the Liebert unit can’t be started today because of a breaker size issue (req 40A we have a 20A spec’d). TA resolved with TJ and Triton will have startup technician for unit after power is provided. TE received submittals today from JCI and GD has been looking them over all day. KM & TTC emailed Triton with acceptance on the retro-commissioning items on RTU-1.

Replace the Supply Fan Motor $3566.00 – APPROVED

Replace two bad compressor contactors $408.00 – APPROVED

Repair static pressure control for exhaust $2720.00 – APPROVED

Replace one bad condenser fan motor $1149.00 – APPROVED OR REPLACE WITH ALL 6 DEPENDING ON PRICE TBD

Perform leak check of Ref system, repair and replace refrigerant as needed $2339.00 – HOLD

Cleaning of coils and inside if unit including combing of coils as needed $1633.00 – APPROVED

Repair of Inlet Guide Vane and controls $3931.00 or replace same with VFD control $5108.00. – APPROVED BASED ON AVAILABILITY OF PARTS

03/23/07 – Lobby lights are on order and all issues with this item have been resolved. 480V electrical service has been changed over by Duke Energy, has been inspected and is up and running. 240V service to be removed the last week of March 07, tentatively.

03/20/07 – 02/19/07 – Last week the structural steel modifications began for the Carrier unit changes. Triton provided a dwg to SSA and CRA showing the new steel dims. SSA is sending back a proposal. The return duct shaft is open and must be sealed to provide proper airflow back into the unit. CRA is checking on some mold issues in that shaft from a roof leak. The FPC dwgs have been submitted to the city. The new FM is located and routed through an existing coredrills in the connector. TA release approval on change orders for TJ Williams for emergency lighting battery packs and dishwasher power in the 1992 building. Triton is still behind on the VAV boxes for the 1992 building and will have to work with Empire for ceiling pad out.

02/09/07 – TTC walked the site last Tuesday to check on progress. MEP is moving forward okay. Duct chase holes have not been cut yet. Triton is installing ductwork on the 1992 building

01/17/07 – Triton has coordinated with Empire on existing work. Kevin is worried that we don’t have enough written documentation that we told Triton to get this work coordinated and scheduled with the owner

01/15/07 – Somehow this job has not been tracked. Here’s the catch up data. The MC is Triton, the EC is TJ Williams, the PC JMW plumbing. We have had several job meetings. Triton has been poorly represented at these meetings. Last week after missing a rescheduled job meeting, Triton came on board with some initiative in catching up to the rest of the contractors. Triton has added more people to the job and is beginning new HVAC in the 92 building. Triton has demolished two floors of work in the 56 building. Triton has 2 men finishing the demo of the multi-zone unit and boilers in the basement of the 56 building. Triton has submitted an RFI questions who is cutting the holes in the shaft floor for his ductwork. TTC responded the GC and this was confirmed by CRA. The RFI was replied to but Empire responded they did not agree. The work is shown on the Demo and Arch dwgs for the GC and the summary of work says superstructure work is by the GC. The CWW has denied the underground permit and tap because the CGE pit is in the Right of Way of the street. TEI has negotiated with them and Cin-Day Fire Protection to provide the new line inside the building through the connector and throughout the building. This should save money and remove trenching and such in the yard. Kevin Maue has agreed that he likes this approach.

09/14/06 – Weekly design meetings continue. We reviewed cost estimates with the client and architect. Final estimates and drawings are due on September 25th for review and bidding is scheduled shortly thereafter.


MRDD Thomas Kidd Renovation

08/15/08 – KM reported the variance was accepted.

07/24/08 – KM and Tim Jones rehashed unit placement and easement issues at Thomas Kidd. This issue is still not resolved. The unit has never been moved and has always been on all review prints in this location. It does sit on an existing easement the MRDD already has.


MRDD Jackson Center Renovation

07/28/09 – Peck dropped off O&M’s today for this project. TEi will review them and give them to kevin for Closeout.

06/23/09 – This project is complete. I am not sure if Kevin has received all his closeout documents, please check and confirm at monthly meeting.

05/22/09 – We are closing out the project with TJ Williams next week. All work is complete and we are down to closeout documents. the Substantial completion for this project is Feb 20, 2009.

04/15/09 – CR Architecture + Design awaiting final punchlist items to be completed and closeout documents.

04/10/09 – Back punchlist issued.

03/09/09 – TTC sent out punchlist for Mechanical, Electrical and Plumbing and finalized Allowance & Change Order Payments with the punchlist so the contractors had everything they needed to close out the project.

02/19/09 – Project was punchlisted yesterday. Punchlist due at the end of the week.

02/02/09 – TTC continues to attend const mtgs for TEI. TTC drafted a letter to TJ Williams outlining the problems with the licensed door operator. The letter states that TJ Williams is responsible for hiring the licensed sub-contractor per the drawings and specs. We are however going to try to minimize the cost to TJ of this extra work that we believe they did not have in their original bid.

12/23/08 – Const Mtg 9 shows we are ahead of scheduled. TEI needs to address the sink in room 47, the trench drain, the curve sink is on site, the junction boxes need to be t&m by tj.

11/24/08 – Phase II is complete. TEI must do a Phase II punchlist. Peck requested a sheave change on HV-1 to get new airflow. TEI approved and sent to Kevin. TEI completed new allowance change orders for TJ Williams and sent them to John Parfitt for approval. TJ Williams revised payapp was received.

10/16/08 – Phase I is complete. Phase II has begun. Trenching for new sink was this week. No reports after TTC clarification was given. We are hoping things go well.

09/10/08 – TTC contacted Carrigan & Grimm about WC-4 for this project. MRDD wants light lens on warehouse lighting. Phase 2 is first. Phase 1 is second. All permits are approved.

08/25/08 – HVAC and Plumbing submittals are complete. TEI sent back to contractors. Building & HVAC permits are ready. $2,769.87 for the building permit $1,045.45 for the HVAC permit.

08/22/08 – TEI sent permit drawings to the project team. TEI sent out all HVAC submittals back to PHB.

08/08/08 – No update on approved IBI permit for this project.

07/29/08 – TTC hand delivered third revision contracts to KM at MRDD. We are awaiting project kickoff.

Bids came in acceptable. Pre Award meetings are complete. Contractors are approved.


MRDD Due Diligence Report

09/21/09 – Kevin called with final comments on the final draft. TTC incorporated these and other comments into the final printing. TEI made 3 copies and mailed 2 of them to Kevin. Per Kevin’s request we will assemble the report into pdf format and email as well.

08/25/09 – Reports are due Sept 4 to Kevin Maue.

07/29/09 – All buildings have been visited. The reports are about 50% updated. Cost estimates, write-ups and grade cards remained to be inputted.

06/18/09 – TTC turned over the final Beckman report for Kevin to review. TTC spent a few days proving that the EH rate is the best rate for the County in these buildings. During this investigation, we found that several of the MRDD buildings are already on the EH rate but are also buying gas on the open market from a broker. TEI has visited Beckman, Northside and Franks so far and completed the field work. The Beckman report is complete. The Northside report is going to be sent to Kevin next. TTC will have Folck & Counts write up and estimate the Franks Adult Center report next week. Kevin said that we will be visiting the schools next. Try to get a date and time for Danny and Amy to do the next building.


MRDD 1520 Madison Road Renovation

12/31/09 – There are minor control issues with the economizer in question but the building and controls are working as designed. TTC archived on 12/31/09.

10/20/09 – Kevin has JCI coming out to change return air humidity to zone average humidity on RTU-1 on Thursday. TTC may attend. Three items remain on the service punchlist. TTC contacted Brett to get answers about economizer control,

09/22/09 – I believe the final controls punchlist items to be complete is to acquire the software CD’s and backups of the current settings.

08/25/09 – Building is overcool in mornings. JCI is supposed to come out and fix timing and response loops. Many emails and phone calls have been made trying to get this done.

07/29/09 – Today we received updated As-Builts on the controls for the building from Chuck Vedder. I forwarded them to Kevin Maue for his keeping. I am also printing out a full set of 11×17’s to keep in the basement of the MRDD building again for our use. Everything seems to be running really really well.

06/23/09 – Triton, JCI and Carrier have finally been to the building enough to help us change the programming on the unit to get it to coordinate with the JCI system. This project may finally be finished. Things are working well in both buildings.

02/02/09 – Triton is reporting that Chuck Vetter will finalize this project for 1520 Madison. TTC contacted TJ Williams for the smoke detector work on the 1956 unit. TTC sent a summary email for controls to Brett at Triton for final completion items. The unit is finally running well but the building pressurization & the duct heaters are not working quite right yet.

10/16/08 – Grady says Triton will put together a final sequence that incorporate everything that is really working on the project and the way it works.

09/26/08 – Const Mtg 2 is over. TEI responded to TJ Williams RFi about another light in the hallway. TEI is approving the PHB Honeywell controller for the HVAC unit.

08/25/08 – Triton emailed today to discuss remaining closeout items. TTC explained about OA sensor.

08/22/08 – Triton is ordering a new board for Boiler 2. Triton relieved the pressure on the hot water system to prevent relief valve issues. Triton may have to get TJ to troubleshoot the overloading problem on the recirculation pumps. Kevin emailed that the OA sensor was flaky again.

07/29/08 – Brett Harpel of Triton discussed final control sequences with JCI. Modifications to system were made a month ago and again last week. Building is running fine. KM wants TEI to bless the control system as it meets design intent of the specs. JCI replied stating TEI sequences were in fact in place on RTU-2 but local overrides were keeping the system from operating as designed on that day.

Triton asked to switch the disconnects serving the RCP to troubleshoot low voltages problems at Boiler-2.

Triton to test expansion tank pressure to see if it is causing over pressurization.


MRDD – Beckman Fire Protection and Fire Alarm Upgrades

10/28/10 – Phase 1 & 2 are nearly complete. Punchlist is scheduled for 10/29/10 at 8am. All lights are installed, we want to change bulb color to cool white. We need price from Quality on isolation valve. QFP has been advised to replace the ceiling tiles in the main Administration area from the allowance monies.

09/27/10 – Next Const Mtg was rescheduled to Oct 6, 2010.

07/28/10 – Contracts have bee approved and signed. Awaiting first const mtg.

07/23/10 – Bids have been opened and approved. Pre Award meetings are complete. QCE is the electrical. Quality is the Fire Prot Contractor. Contracts have been written and approved by the Prosecutor dept. We are submitting a package back to the bldg dept to answer their concerns on permit issuance.

05/26/10 – Drawing stamped and signed. I will pick up the GC dwgs on Thursday and take to the permit office after Partner meeting. Budgets look good. Emergency lighting is an issue to design and may add more detail to for bidding. ITB was assigned 047-10. Bid set due June 4.

05/16/10 – Client review meeting brought up several issues that we have to fix before permitting. We will permit on 05/26.


MRDD Beckman Plumbing Renovations

10/12/10 – Driekast has replaced all the fixtures. The new ones are acceptable and exactly what we wanted.

09/29/10 – Driekast has provided prices for fixture replacement. Need a direction from Kevin on replacement work.

08/13/10 – BBN has offered the fixtures. Driekast is working on an install price.

07/23/10 – We offered to pay Driekast to install new fixtures if they can get the mfgr to furnish better hands free operators on Beckman. No updates on that progress.

05/26/10 – Fixture replacements still an issue. QCE closeout documents arrived this week.

05/10/10 – Terry contacted Dave about replacement fixtures at Beckman.

04/26/10 – Driekast has several items on the punchlist to complete, TTC and KM have to decide about their change order request. Hands free faucets wiggle, wiring has to be strapped down, balancing has to be addressed, they need to add balance dampers.

03/23/10 – Final punchlist has been sent to contractors. Final inspection is scheduled for Thursday. We have to resolve the air flow issues with the GC on this job in the new restrooms. We have to back check the work on the final punchlist, check with the owner and wrap up any final conditions for closeout and acceptance.

02/22/10 – Const Mtg on site. DC attended and back checked punchlist and took additional air flow readings to be distributed. Change order for heavy duty toilet submitted by Driekast and given to Kevin. The final phase is almost complete.

01/28/10 – Phase I is complete. Phase II is almost complete and will be ready to punchlist next week. Some change orders are in the works and we have discussed handicap water closets being WC-2 in both the men’s and women’s restrooms.

12/31/09 – Phase I is complete. Contractors are working over the break to cut the hallways for the new sink. We have approved the stainless warp around by Driekast for the new sink.

09/17/09 – All contracts have been typed and given to Kevin for prosecutor review.

08/25/09 – Bids have been received. Driekast was low plumber. QCE was low electrical. Butler County Electric was low General and Mech.

07/29/09 – This project has been sent out for bidding. Bid opening scheduled for 08/25. The prebid is scheduled for 08/11 at 2:30pm.

06/18/09 – TEI and CRA completed the design development review with the client. This project is scheduled be out for review back to the owner in mid July and bidding in late July.


JC – Kitchen Trench Drain Repair Project

04/05/05 – Queen City Final Paperwork arrived at TE for review. TE to pass along to Mark Hardin soon.

01/14/05 – TTC sent out allowance confirmation with final instructions to Brian for project final payapp.

01/11/05 – Change Order Approved. TTC signed today and sent to QCM for signing. $8028.68. QCM removed 2 hours from the change order for miscommunication on the handle removal.

11/18/04 – Tiles are installed, Steam kettles are being installed by Peck, Piping is being finished by Peck, QCM will be down this evening to anchor the remaining leg holes. TE would like County to request second shift of Peck workers to complete steam piping so kettles are ready to go by 5 a.m. on Friday morning. Bernie is to call TTC back when he gets off the phone.

11/09/04 – The trench pour is not smooth enough to meet TE expectations. QCM said they will grind down to make smooth and backfill with sand based concrete to meet specs. Siemering is onsite prepping the floor for new tiles. Peck will be in next week to finish the steam kettles. The project schedule is very tight and Mark Hardin and Terry Cannon are trying to keep on top of items as they arise to meet the November 18th schedule. The kitchen should be ready for use on Friday morning November 19th.

11/08/04 – Epoxy on new concrete issue has been resolved. We have directed QCM to follow a new schedule. TE and County met Siemering tile onsite this morning to review the scope of the work. Floor tiles won’t be onsite until Monday 11-15. To accommodate this change of plans, the following new schedule is being incorporated into this project: QCM will demo forms 11-08, County will process Siemering background checks, QCM will provide new studs for kettle installation to replace the ones damaged during removal on 11-09,11-10. Siemering will prep the floor on 11-11 and 11-12, QCM will apply first coat of epoxy on 11-11, no QCM work on Friday, second coat on 11-13, attach stainless steel angles on 11-13, Siemering will begin tile installation on 11-15, finish on 11-16, QCM will install steam kettles on 11-17, Peck will finalizing steam kettle piping on 11-18. Project will be punchlisted by TE. This is our drop dead timetable.

11/04/04 – TTC visited the site. Things are going well. New schedule looks like this: 11-4 pour new trench let cure, 11-5 acid etch new conc, 11-8 apply first coat of epoxy, 11-9 apply second coat of epoxy, 11-10 start tile work, 11-11 complete tile work, 11-12 reinstall steam kettles. PHB will have to finish steam piping by 11-11 and we confirmed today with Mark Hardin that only condensate would be stainless piping. PHB can continue to work during the day while QCM works at night.

11/03/04 – PHB is on site doing steam work. QCM has found and removed the valve handle. The new concrete will be poured on 11/04/04. TE has received and is approving the Pitt-Guard Epoxy Coating (USDA and FSIS approved) for use on the trench surface. QCM is to follow the recommended installation procedures for this epoxy for proper installation including concrete preparation (acid etched or brush blasted) and pour the concrete so that it is self-priming as listed on the cut sheet. QCM is also to provide 3 vandal resistant floor drain covers on the existing trench floor drains and two stainless steel angles to support the stainless grating being put back in place.

11/01/04 – Peck reported that they are all set to begin the steam piping work and have coordinated with Tony about this project. They should be on site and ready to work in the next couple of days.

11/01/04 – Karen Draper from the City Health Department returned my call and told me that our plan to install smooth concrete with an epoxy coating in the Justice Center trench with vandal resistant flood drain grates. After my latest phone call we actually discussed the existing condition and the Health Department is very much inclined to approve our new design with the concrete and epoxy coating where we remove all instances of standing water under and around the SS trench in lieu of the concrete/epoxy hybrid. Karen said that any FDA or USDA approved epoxy adhesive and sealant will suffice so QCM can submit what they plan to do to TE and we will look it over and give it our blessing as soon as we receive it. QCM please also provide the stainless steel angle irons along each side to support the grating, attach with adhesive and i would recommend that anchors be installed in the vertical sides about every 6 ft o.c. to hold the angles in place while the grate will hold down the top. PDF

10/27/04 – QCM is on site and beginning work. TTC showed up at 8:00 to coordinate final details as all the important people took care of business from 7:30 – 8:00 p.m. QCM must submit security releases for the coring company TODAY!

10/25/04 – QCM is ready to begin work on 10/25, they need approved security checks but are planning on being on site at 7:30 p.m. as scheduled.

10/11/04 – QCM delivered the project schedule and it was forwarded to MWH and HS. TE approved the timetable for mid November completion.

10/11/04 – TTC & HS are asking contractors to give a price for replacing the insulation on the steam piping behind the steam kettles while the system is down in October/November. TE is exploring a insulation base paint that can be applied. Hy-Teach Thermal Solutions is one site for this paint that TE found. We have calls into their engineering department for information.

09/24/04 – Notice to Proceed was given by MWH via a phone call. TE to setup up pre construction mtg.

Bidding is complete.


JC – Battery Room Fan Coil Upgrade

07/12/05 – Final paperwork arrived today for this project. TTC copied the documents, signed the final pay application and dropped it into the mail for the County.

06/30/05 – Filter rack is fixed. Final paperwork has not been received. O&M’s have been delivered to the County via a hand delivery by TTC. TE contacted PHB today about final paperwork status. PHB will get out soon.

06/15/05 – O&M’s Received today. They were approved and will be held for final payapp paperwork. PHB should be down in the next day or two to fix the filter rack and then they will send in all final paperwork.

06/09/05 – PHB called to questions the $2500 allowance on this job because it was not in the contract. TTC checked the bid documents and the contract and PHB is right. The original bid of $24,400 did not leave enough room to add the allowance alternate without going over the $25k limit by ORC. We will now have to write a change order to increase the budget slightly to accommodate the newly added automatic air vents which are a necessity in a chilled water system like the one we are using at the JC. Final paperwork should be down to you in about a week.

06/07/05 – PHB reported that they would be down to finalize the punchlist about June 20 – 27 or so.

06/01/05 – TTC sent an email reminder to PHB to finalized punchlist and finish the filter box.

05/05/05 – Punchlist distributed to PHB. PDF

04/05/05 – TE stopped at site today to punchlist the FCU job. SBO did the site work. Punchlist will be distributed very soon.

03/15/05 – PHB requested a change order for the new automatic air vents. TE asked for the cost and will suggest the County accept this allowance usage as part of the project that wasn’t included in the original scope. We are awaiting a cost associated with this additional work. I can see now that the last payapp was approved for too much money because we have a $2500 allowance that hasn’t been used at all, but that is okay now because we need to pay for this additional work and that will come out of the allowance. I guess PHB didn’t break out the allowance as a separate line item on the schedule of values.

03/15/05 – TTC coordinated with Howard and Bernie since payapp only left $2300, TE wanted to make sure that controls and electric and vents were all installed and working by the end of the week so that TE can do a punchlist soon. TTC left a messages on Bernie’s cell phone.

02/28/05 – IBI Permit set in PHB hands. Fill up and startup of new FCU should happen within a week of electric hookup and IBI approval. Tentatively we could be looking at startup at the end of next week.

02/25/05 – TTC notified PHB today that the return transition is not necessary on this project as it would prohibit access to the return filters. The supply transition is being fabricated at this time and PHB should be out to install it in the next week or so.

02/16/05 – TTC checked on the status of this project today. Piping looks good. Insulation needs to be installed. Thermostat has to be installed. Unit has to be started. Pressure test has to be removed. Final piping must be filled and checked.

02/07/05 – Doug Turner was onsite making the tie-in connections over the last weekend. They drained the system and refilled it and made the tie-ins

02/02/05 – Permit has been applied for by Peck for this project.

01/25/04 – Peck was onsite to review project. TE is to provide stamped signed dwgs for permit and to discuss supply diffuser and return grille on fan coil unit.

09/23/04 – Bids Received and Posted. TE awaiting County’s approval to proceed with contract. See bid results below

09/13/04 – Bidding Underway


JC – Task Force Room Upgrade

09/18/05 – The city has approved the whole permit. This job is complete.

07/22/05 – Steve Ossege called and said the City does not know of any approved independently support handrail system that meets code. He will however approve the platform arrangement discussed on site.

07/18/05 – Ed Enderle of DK called today and quoted $2450 to build a platform and railing for the Task Force Unit. TTC asked them to look into temporary handrails that were eluded to by Steve Ossege. More information to follow.

06/27/05 – TTC and DK have been working on getting the permit issued and closed for a week. Amit Gosh called and verbally told me that he was signing off on the permit today. TTC relayed this information to DK for finalization.

03/24/05 – JCI reported that wiring and status condition has been wired from Task Force unit to DDC. Programming is complete and task is complete.

03/22/05 – DK delivered the final payapp today along with all the final paperwork. TTC will sign and send to Diedra in the Sheriff’s Dept for project completion.

02/24/05 – Debra Kuempel (DK) deliver payapp #2 today. It was approved by TE. All final paperwork should be complete next week.

02/22/05 – DK stated that all final paperwork was sent over last week to TE. TE is going to finalize and send to Diedra Burns this week.

01/25/05 – TE contacted Ed Enderle as all work is complete and signed. TE request final paperwork and signed punchlist!

01/24/05 – TEI called DK and spoke to Michelle Combs in regards to when we may expect Pay Application #2 and final documents. Michelle was not aware that the project was finished and said that she would check with DK personnel and will call back, I told her that Ed Enderle may be of assistance.

12/15/04 – Debra-Kuempel Pay Application #1 not found at Hamilton County. Transmittal and ThermalTech records indicates that the application was mailed September 24, 2004. Doris called and left a voice message for Michelle Combs informing her that the application was lost and asking her to resubmit an original application.

10/12/04 – DK has finished project. The allowance has been used to get a new power source for the computer room unit that doesn’t shutdown the UPS system. TE has instructed DK to finalize paperwork and punchlist to complete this project. Howard reports that things are working okay to date (4 days runtime at this point)

09/29/04 – TE emailed DK for an update and schedule.

09/23/04 – TE directed Debra-Kuempel to redo their Change Order/Allowance price to re-use the existing MCC starter that use to be for a now abandoned pump to the get the cost under $5000 so we can use the allowance money without writing a change order for this project. We have $5000 in allowance money to use towards re-doing the power to the computer room unit. Once this is complete all work on this project and the punchlist items should be taken care of at the same time. Look for completion by September 30, 2004. The will coordinate all field visits with Howard Seal before starting on site.

TE has to finalize power change work with Debra Kuempel.

TE has distributed punchlist.

Compressor short cycling has been eliminated by utilizing the hot gas bypass feature to change the performance of the system.

TE to work on a new power source must be found and the unit hooked to emergency power panel instead of the current UPS connection.


JC – Exhaust Fan Upgrade

06/15/05 – TE received final paperwork and change order signed for this project. TTC will be mailing it down to Diana with final balance report for HC records. This project is complete. The County is getting $3256 back on this job! O&M’s received and forwarded to County.

05/25/05 – PHB Jerry Govert called today to organize final paperwork. TE sent out a change order to recouple the $3256 in remaining allowance money. All final paperwork should be received in a couple of weeks. PHB should include the CO Credit with final paperwork.

04/22/05 – TE does not or has not seen final paperwork on this job. The balance paperwork was marked as received but no pay applications have been received. Please submit with all final paperwork for payment. Please provide all final paperwork and allowance usage details because we did us the allowance on this job.

02/25/05 – TTC ask Ken Shad to give me final paperwork and allowance change orders to wrap this project up today. He said that he would get right on it.

02/16/05 -TTC has requested all final paperwork on this project including balance reports. PHB sent to TE office on Tuesday. TTC to review.

01/18/05 – PHB stated that all work is complete including the 2 new 5 HP motors. The exhaust % increased by 18-20%. TE is awaiting all final paperwork on this job for closeout. Rick Ahrens will stop by to take a look at a motor squeal on Thursday.

12/27/04 – Howard reported today that EF-C1 has high pitches squeal. TE called Peck, but they appear to be out of office today. TTC will try first thing Tuesday morning again.

11/26/04 – PHB has acquired all existing cfm data from the fans. TE is reviewing to provide a direction to the PHB construction team. It appears as though the allowance money should be used to replace the motor on EF-D2 from 3 to 5 HP. The wire size is sufficient. The existing 3 hp bearings will be taken out and put on EF-D1 as that fan needs a bearing replacement. PHB reported that even with the duct cleaning there are still some security grilles that are clogged with dirt and the secure design makes them virtually impossible to clean out from inside the cell area. TE to advise how to handle this situation. Sam Boehringer will be reviewing submitted data on Monday November 29 at TE and then we will setup and onsite meeting to go over the project with the owner in the coming week. We have a $5000 allowance to use for these repairs.

10/15/04 – TE has contacted PHB about and updated schedule for this project.

10/05/04 – Pre construction meeting was held and PHB plans on starting in 2 weeks. Schedule is basically one EF per month.

09/20/04 – SOSS revised for additional Champion monies and will deliver on 09/21/04.

Champion provided duct cleaning on existing systems. Hired through ThermalTech.

Peck Hannaford & Briggs won the bid ($26,050). Base Bid, Alt 1, Alt 2 and $5000 Allowance. EF-C1, EF-C2, EF-D1

Contract written and being reviewed.


JC – DDC Upgrade

03/03/06 – JCI finished the upgrades on Wednesday. The original CD’s were delivered. Some work need to be done to fix what the upgrade busted and then a new backup needs to be done. Then Tony will pay for project. TEI told the contractors this by email today for JCI.

02/22/06 – TTC received final paperwork and approved for this job. TTC is going to approve and give to County even though final CD’s and backups has not been received yet. This is the only JCI project that remains to be backed up and CD delivered to the Owner. The County may at their prerogative continue to hold final payment on this invoice as they do not have the licensed software in hand yet. TTC has emailed Al Bell and has not heard back about getting final CD’s to Howard Seal for this job so I doubt Tony is going to release this money until the material is in his hands. Please let ThermalTech, Howard and Tony know when it arrives on site and please do this ASAP.

01/18/06 – Wayne Jackson reported that a printer with a USB connection would have to be used for the JC ADS server since no printer ports were available. TTC informed JCI that the County has already taking care of this item. Wayne said that he would have Mike Asbrock call Howard Seal directly and prescribe a safe and easy method for deleting the 93,000 alarms in the alarm container. The licensed CD and Backup still have to be provided to the County on this job. No remaining punchlist items remain.

01/10/06 – TTC emailed JCI about installing a report printer on this project and about closeout CD’s for this project.

12/14/05 – TTC emailed Tammy back at JCI to inform her that the project is not complete and cannot be paid until everything is done.

10/24/05 – TTC contacted JCI again about punchlist by email.

10/14/05 – TTC and JC have not seen any final punchlist from JCI yet. TTC contacted main office to get closure on this project!

10/03/05 – JCI reported that they will finalize the punchlist in the next two to three weeks. TTC told Howard about this schedule.

09/16/05 – TTC contacted JCI about received punchlist and trying to get a schedule for completion today.

09/08/05 – TTC contacted JCI and approved their final two pay applications but requested all the final paperwork and TTC has to send out punchlist immediately on this job. Punchlist has been sent to JCI and County.

08/08/05 – JCI Reported that VFD’s have been reprogrammed and are 100% operational. TTC to provide final punchlist on Tuesday, 08/09. When complete this will finalize this project.

07/21/05 – Tony called to say that escrow on the JC DDC project has voided App #14. JCI should resubmit App#14 for $20,921 and App#15 for $23096.00. TTC instructed JCI to send in all final paperwork as well as this project is complete except for the punchlist.

07/05/05 – JCI is re-sending payapp # 14 which appears to be correct. TTC reviewed and asked for 1 original and 3 copies. JCI and its electrical contractor are installing the bypass around the VFD’s. They should be done this week. At that point the project should be complete and TE will do one final punchlist on it.

06/14/05 – JCI prematurely sent a final bill into TE. B&J and JCI aren’t scheduled back onsite until 06/27/05 for the final phase or work on this job. B&J called today to say that all the sheetmetal is prepped for installation. B&J also said that two of the boxes were ordered as 6″ inlet and 10″ inlet but actual field measurements showed the boxes they are replacing are 8″ and 8″. TE approved the new sizes as they can be balanced to what we need in the space. JCI is on vacation

05/02/05 – B&J confirmed that all VAV boxes should be shipped the first week of June. TE advised the JC FM staff.

03/24/05 – JCI has automatic reports working for Howard. The new cooling tower sequence is the only project punchlist item remaining. All commissioning reports (hardware verification) sheets are in the as-builts. JCI will provide new as-builts to insert into set based on the change order work scope currently going on.

03/16/05 – New quote from A&M came in for DDC work at $12,091.53 for B&J Jacobs to do the actual installation. Since they were the original contractor, this would be great. TE can supervise the installation and just turn it over to Ed Braemer when we are complete or still have Ed work with B&J during the installation. I need approval from someone before we let them go. JCI reported that wiring and status condition has been wired from Task Force unit to DDC. Programming is complete and task is complete.

02/28/05 – TTC met with Howard, Peck, Debra and Langdon to solicit additional prices for the Dual Duct Box replacement at the JC. Quotes are due this Friday at Noon. I will coordinate with Ed Braemer on moving forward after we receive prices and we talk to Tony and Howard about authorization. In the meantime, I have asked Ed to call Howard to discuss reporting problems with the new software.

02/16/05 – TTC is collecting a packet to get Peck and Debra to provide a follow up price on the Dual Duct Boxes in the Sheriff’s Area. Howard reports that JCI is almost to the point where they are out of work.

02/15/05 – JCI gave us a price of $3557.40 to hardwire the VFD’s to the Fire Alarm Panels. Should or can we approve this work?

02/03/05 – Construction meeting complete. JCI to install final version of software next monday. TTC to gather additional mixing box estimates for AKM. Fire alarm issue should be hard wired through VFD’s so they can’t be overridden from the front end software (JCI to work on).

01/25/05 – JCI came onsite last night to work on 2 dual duct boxes in Sheriff’s area and to work out some issues with EDB. TE to follow up with HS today on status. Existing boxes don’t have accessible actuator shafts for DDC upgrade. We may have to improvise to modify them for retrofit. EDB, TTC and HS to meet and discuss if needed.

01/24/05 – Ed Braemer or Johnson Controls, calls TEI on required app doc and will follow-up with Tammy of JC to see if she has everything needed to submit the apps.

01/21/05 – Received JC Invoice for Application #13. The request is a 4 page Original Invoice form. Tammy of JC contacted and asked to submit Application on the standard AIA forms and that a completed Affidavit form should accompany the AIA forms.

01/19/05 – TE stopped at the JC to investigate the “howling” on the Stairwell Pressurization Fan.

11/09/04- Submittals on the VFD fans returned. Pay Application #11 returned to Ed at JCI. Both returned to JCI via UPS.

11/08/04 – Voicemail left for Ed Braemer regarding incorrect Pay Application #11.

11/08/04 – TE is reviewing the submittals on the VFD fans this week. They should be complete and returned this week.

09/23/04 – JCI needs County approval to begin. What is the status of the project. They are waiting on a go ahead from you.

New change order approved. Waiting to begin new work.

JCI has completed all phases of the project except for the cooling tower programming.

A new round of change orders has been accepted by the County and the paperwork is en route to JCI for signing.

Steve Steinkuhl to issued a final punchlist for the project by September 3, 2004.


JC – Aramark Office VAV Upgrade

05/23/06 – TTC moved this project to the archive as all work on this job will be done under the fingerprinting ID project that is still open.

11/08/05 – TE heard back by email form Jill Williams and we will set up a quotewire job for this.

11/04/05 – TTC and Tony discussed including the purchase of the Dual Duct boxes today and decided to buy this box with the other 8 for Fingerprinting and ID. Trades will install these boxes.

10/28/05 – Prebid meetings were help on other JC projects but this job will be held for Trades to do work.

10/14/05 -Design is complete. TE will have trades do the work on this job to replace the existing VAV box with a larger box and redo the controls. No permit is necessary on this job.

10/04/05 – TTC & SRB have visited the site and discussed VAV options with Howard. Since this space is not part of the Dual Duct air handler, there will most likely be overheating and over cooling of the office during low load periods as the minimum setting on the VAV box is 145 CFM and the air handler supply temperature is reset based on OA and averaged room temperatures. TTC emailed Tony about just having the work done in house and having JCI do the controls directly.


JC – Boiler Room Exhaust Fan Upgrade

05/23/06 – TTC archived this project because there is no way to make this work without a lot more make up air into the room.

10/28/05 – TE still has not found a viable solution for the heating problems in the space and will look at other usages.

10/24/05 – TTC has not found a usable scheme for boiler exhaust to date. There is not enough cold air to make a dent in boiler room temperature without additional cooling.

10/14/05 – TE cannot find an exhaust fan design that works without additional cooling.

10/04/05 – TTC & SBO have visited the site and are designed the new exhaust fan design to help remove heat from the boiler room.


Justice Center – Backup Chiller Project

09/01/06 – TTC worked out final billing details with Bridgette today by phone. DK is sending over final invoices for payment.

08/30/06 – Steve and Terry finalized the number required on the final payapps. Laurie is typing up the final change order for the project. TTC has to coordinate the change order payment with Tony Matre.

08/29/06 – TTC reviewed the final pay application by Debra Kuempel. Change orders amount were added to the contract amounts and this is not correct so TTC has to contact DK to figure out the correct billing method.

04/03/06 – JC asked DK to insulated the chilled water lines to prevent pooling under the chillers.

03/24/06 – DK has not been on site doing any work. They may wait until it gets warmer outside to make the boiler changeover work. The chiller scope is nearly 100% complete.

02/23/06 – TTC was on site today to look over chiller project. New 12″ mains look great and are installed and cored drilled and anchored.

02/22/06 – We approved some small allowance work today described as: *Cut (2) 11/2″ nipples off existing thred-o-lets on Chilled Water Supply and Return piping. Weld on B/W caps on to existing thred-o-lets to cap lines off. *Install 2″ thred-o-lets with ball valves and caps at the bottom of 12″ Chilled Water Riser for drains on both supply and return piping. Total cost for the work outlined above is $642.00

02/16/06 – Howard asked about replacing valves if they don’t hold and about removing a 1-1/2″ nipple in piping when drained down. These are two agenda items for Friday Const Mtg.

02/10/06 – TTC release this email after talking to Howard this morning:

Steve, When you go to the drain the Chiller plant at the JC, it was our belief in the bid documents that the butterfly valves near the chiller will hold the water in the system and prevent a complete drain down of the mechanical plant. If this turns out not to be the case and we have to drain down the entire building, there is a concern about who is replacing the nitrate in the CHW. It is my belief that this is a condition that will be covered by the allowance we have on this project if it proves necessary. I don’t foresee this being more than $500 or so if in fact we need this. Please let me know how the valves hold next week when you begin to tie into the existing system. If you have any questions, please let me know. thx, ttc

02/02/06 – TTC contacted DK about working on the chiller project and they said they would get started.

01/17/06 – TTC held the pre construction meeting today and distributed the meeting notes. The boiler project will start 03/15/06 and last for 60 days. The chiller project will begin immediately and last for 60 days.

01/10/06 – Pre Const Mtg scheduled for 01/17. Everyone has confirmed for meeting.

12/13/05 – Contracts were returned for corrections. TE corrected, printed and returned to the County for submission to the commissioners.

11/08/05 – Tony gave TE the okay to type the contract for Debra Kuempel on this job. We are taking the base bid and $2000 Allowance.

11/07/05 – TTC opened and tabulated bids for this project. See link below.

10/28/05 – Prebid meeting was held with all contractors. Bids due Friday to TE.

10/24/05 – Drawings are complete up to TTC’s review. Prebid is set for this Friday.

10/14/05 – TTC promised final design complete on Wed 10/19.

10/03/05 – Tony approved our recommendations and we are going with the plan to put in the chilled water tie

09/29/05 – TE recommended to the County that no electrical xfer switches be included in the temporary job because the expense of putting in electric switches just isn’t cost effective to providing emergency generators on site at the time it is needed. Feeding the chillers is problematic. TE will bid the mechanical work out as a small project.

09/27/05 – TE pulled all the old designs from the archive and setup the new project. Electrical is a big issue as this will put the project over $25,000 unless we use generators.


JUSTICE CENTER – Backup Boiler Project

09/01/06 – TTC worked out final billing details with Bridgette today by phone. DK is sending over final invoices for payment. TTC sent out the change order for the job to Steve V.

09/01/06 – Tony approved the $1597 change order by phone today. TTC to advise DK to wrap up the project.

08/30/06 – Steve and Terry finalized the number required on the final payapps. Laurie is typing up the final change order for the project. TTC has to coordinate the change order payment with Tony Matre.

08/29/06 – TTC reviewed the final pay application by Debra Kuempel. Change orders amount were added to the contract amounts and this is not correct so TTC has to contact DK to figure out the correct billing method.

08/24/06 – Al Bell reported that the high pressure sensor was re-ordered and will be installed very soon.

06/26/06 – Howard reported that DK will be onsite tonight and begin the work on the final changes to the job. JCI and DK have not been in contact with each other yet on the new sensor.

06/06/06 – TTC contacted both JCI and DK again about progress. JCI came in last week but needed a different part to make it work. DK is supposed to be down this week but TTC will find out more details from Steve Vater and let everyone know. As of Monday, Howard had not heard from DK about coming back.

05/23/06 – TTC emailed both Debra Kuempel and Johnson Controls to try to get them scheduled to finish this project up. TEI will pay the JCI cost to the job and DK will bill towards the allowance to wrap this up.

04/03/0 – DK is going to be on site the week of 04/10 to finish up the boiler work. JC asked for insulation on chilled water system to prevent condensation. TTC relayed this information to Steve Vater and asked DK to coordinate their time and visits directly with Howard Seal.

03/24/06 – DK has not been on site doing any work. They may wait until it gets warmer outside to make the boiler changeover work. The chiller scope is to b nearly 100% complete.

02/23/06 – TTC visited the site today with the contractor and Howard. TTC wants to remove a check valve from the main steam supply from the courthouse because the pressure drop will be to great to keep this device in line. An existing chain operated gate valve on boiler #3 will suffice in place of the one noted on note #3 on the dwg. The non-return valve can come out as planned. At this temperature outside the steam plant from the courthouse can’t be shutdown until 7pm and it must be back online at 3am. DK will have to coordinate this with Howard and the CH staff.

02/22/06 – We approved this allowance usage today:*Install (2) 3″x21/2″ weld reducer on new 3″ pump condensate piping at tie-in point to existing piping. Existing pump condensate piping is 21/2″ in size. Debra-Kuempel quoted total cost for the work outlined above is $261.00

01/17/06 – TTC held the pre construction meeting today and distributed the meeting notes. The boiler project will start 03/15/06 and last for 60 days. The chiller project will begin immediately and last for 60 days.

01/10/06 – Pre Const Mtg scheduled for 01/17. Everyone has confirmed for meeting.

11/08/05 – Tony gave TE the okay to type the contract for Debra Kuempel on this job. We are taking the base bid and $2000 Allowance.

11/07/05 – TTC opened and tabulated bids for this project. See link below.

10/28/05 – Prebid meeting was held with all contractors. Bids due Friday to TE.

10/24/05 – TTC & SRB visited the site to pickup some last minute details about returning condensate the DA system when the Justice Center is on its own steam. Drawings are complete up to TTC’s review. Prebid is set for this Friday.

10/14/05 – TTC promised final design complete on Wed 10/19.

10/03/05 – Final design is being complete. Job is ready to bid out for small projects.

08/23/05 – TTC visited the site and found that bypass valve is already in place so that piping work does not need to be done. At this point there are a couple of viable options: 1) put in a new valve near the jc side of the bridge or automate an existing pneumatic valve in the courthouse. TTC sent an email to Tony, Howard and Mike looking to explore our options.


Justice Center North Electrical Device Identification

01/10/07 – Project completed in 2006.

08/24/06 – Drawings were delivered to Hamilton County Facilities. There are no open issues on this project.

05/17/06 – Close-out meeting; no open issues. All work by Queen City has been completed. I will be working on getting the drawings completed within the next several weeks and send (3) full sized sets to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Howard Seal). Queen City presented (2) binders, each with copies of the old and new panel schedules. On the new panel schedules the items listed in bold have been verified; the items listed in normal type are transferred over from the old schedules. The binder also includes diagrams of the MCC’s and Switchgear. A CD was also provided which contains all of the new panel schedules in Excel format.

11/15/05 – Contracts have been submitted.


JC – Replace Sallyport Sliding Gates

02/19/07 – TTC is approving and passing along all the closeout documents for this project. TTC is archiving this project today.

01/02/07 – TJ Williams submitted a letter expressing their extreme disappointment with the County not approving the Change Order for the extra work incurred during the Sallyport project. TTC forward to Mark Donnelly for the County to file with the records are address separately if they deem it necessary. PDF

12/26/06 – TTC archived the project today. Legacy has the smallest bit of paperwork to turn in on this project. Once Donnelly has that, this project is complete.

12/14/06 – Mark called today about closeout and Legacy reported: The guides at the justice center are painted. I am not sure the status on TJ Williams. Kenny is in a meeting until about 5. As soon as he gets in I will get an update and get you the information. I have been working on the close out documents and we are about 80% done. We still need to finalize the certified payroll from two of our sub-contractors.

12/04/06 – Close out meeting was held today. 3 items on punchlist were discussed. Pay App 01 was discussed. Legacy to provide a revised Payapp 01 with allowance shown. TTC signed payapp 01 for Mark D to use as soon as a revised schedule was provided by Legacy. Legacy provided one by email but TTC edited it and sent it back out to the contractor with 8% retainage per the project manual. TTC also sent Allowance Clarification 01 to the contractor for signing.

12/01/06 – TTC sent out formalized punchlist items today and reminded everyone about Monday’s meeting. PDF

11/20/06 – The project is complete. There are two punchlist items which remain: touch up paint on door frame and the control panel box that TJ damaged during construction. The doors are working, the fencing and hardware are all installed properly and TEI visited the site for a punchlist walk through. The TJ change order was denied but it was agreed that the allowance money could be used to pay for the extra expenses. TEI has asked for closeout documents on the whole job to wrap it up. We are having a closeout meeting on December 4, 2006 as a final meeting.

10/10/06 – The structures are installed the concrete is finished. We need to let it cure for about 2 days before we install the overhead doors. TJ Williams has some temporary power and control patches installed in the building that have to be replaced. Overhead door is supposed to bring down the doors on Friday of this week. TJ Williams has not started the change order work on the loop control. We are hoping they will do that this week when they come down to do the rest of the work ready on the project. Legacy has a new PM on the job. Ken has taken ill. Howard was able to get the new guy into the building for the concrete today. In a week or so we should have operation of the new doors. The Sheriff may re-stripe the parking lot and move the dumpster outside the scope of our job.

09/15/06 – Legacy contacted us today to report the steel has been ordered. 09/25 looks like a delivery date on that. The doors are in town and waiting to be installed. Legacy may be down next week to do some anchor work on the new structure and Howard confirms that all security cameras are still in operation after TJ Williams completed their partial demo. Saw cutting and more general work scheduled for mid week of next week.

09/04/06 – Change has left our office to the contractor. Project should move forward immediately. Contractor planning on being on site this week.

09/01/06 – Change Order for $3775 was approved today by Mark at meeting in his office. Sheriff will do demo of existing fencing. Legacy will sink new polls for fencing company. TTC approved stainless steel anchors for this project since galvanized were hard to order on time. TEI to write change order asap.

08/30/06 – Legacy reported an error in the change order pricing and a misunderstanding on ThermalTech’s email. New pricing was sent to TTC by fax today. TTC will forward the information to Mark Donnelly on this project for a decision.

08/29/06 – Change Order prices came back today as $1800 for fences with Sheriff demolition and $7195.00 if Legacy does everything on the job. TTC passed along to County. HS was checking with Sheriff’s for their willingness to still do their part.

08/24/06 – Today we approved the electrical submittals and asked for an immediate change order cost for the new door location on this project from Legacy. We have been waiting at least one week for this amount.

08/16/06 – We have had two kickoff meetings for construction. The first issue was the location of the new doors. The Sheriff has now put forth an edict that the busses that enter the site should be closed between the doors. This was not the case during the original design. TEI has issued a change order to collect costing information for this change order. It shouldn’t be too much money as the a lot of the scope is already included. The Sheriff is willing to provide additional demolition of the existing fencing to make this work and Howard stated that the existing dumpster fencing must remain as is, TEI said that is fine. TTC called Ken Jones today and found out that the CO cost is not ready yet and the security checks have not been sent in yet. TTC advised Legacy to move forward immediately with providing this information and the job is ready to begin. Spencer approved the submittals on his and TEI will distribute them as approved to the vendors.

07/25/06 – We received today that the contract was approved. Pre Construction meeting scheduled for 08/02 at 2pm. TEI to contact Legacy Construction for meeting.

06/19/06 – TEI revised the contract and mailed 3 originals to Mark Donnelly for submission for approval. After approval we will convene a kickoff meeting for the owner and contractor.

06/01/06 – Contract was submitted to prosecutor for review.

05/25/06 – Contract was written for Legacy and mailed to Mark for approval.

05/23/06 – Bidding is complete. Addendums were issued. Bid were opening on schedule. TEI drafted our first contract and will give it to Mark for submissions to the prosecutor department. We luckily increased the project budget because the job came in over the estimate but within the 10% margin but still under the actual Archibus estimate.

04/11/06 – Final Client Review sent was sent to Howard and Mark including the work from CRA and THP. There are still some issues to resolve before bid issuance. Mark and Terry increased the budget to $80,000 to cover the additional work added to the job by a more extensive structural drawing set and permits being added back the project. Security Cameras and lights had to be moved to allow for new overhead door placement on project. TTC coordinate with Howard Seal on this point. TEI located the South camera and light on the dwgs and in talking with the Howard the North camera and light were also located on the plans accordingly. County comments are due back on Friday. Final set is due to Purchasing on Monday morning.

04/05/06 – A second design meeting lead to some changes in the door placement. TEI developed a new drawing showing the perpendicular arrangements of the new overhead doors and modified the drawings notes accordingly. Power will be brought from existing connections. The razor wire will be removed and replaced by the contractor. The north door shall be the first to be replaced. The old sliding gates will be completely removed from the job. TEI added special notes to allow other manufacturer’s to be submitted on but they must be approved before bid day by the engineer. The specs were updated and the ITB number and dates have been acquired. A review packages was sent to the team for initial review.

03/21/06 – TTC and Mike Haney spoke today about the project’s structural requirements and THP had some great ideas. Since these are overhead doors they can be turned perpendicular to the building and only be 15 ft wide thus making the doors smaller and the support steel smaller. This sounds like a great idea. TEI will approach the owner about this condition.

03/16/06 – Design meeting schedule on site. Sheriff has approved type of door that we will use for this project. Now we just have to finish the structural design and bid the project out.

02/10/06 – TEI release a review set for comments to the County for this project. TTC still has to discuss product with overhead door and finalize ITB. We have not requested an official ITB from purchasing yet.

02/08/06 – TEI and CR met on site with Howard and Tony and reviewed scope of work.


JC – ID & Fingerprinted Dual Duct Box Upgrade

03/15/07 – All Closeout is complete. Howard is satisfied but JCI has a bit of tweaking to do on the sensors. This project is ready to archive.

02/28/07 – JCI dropped off as-builts and O&M and final payapps for this job. TTC to approve and send to Tony. TTC approved and gave to Laurie to send to Tony.

08/30/06 – Triton has all schedules for the proejct and are putting the submittal data together right now. They are not sure about having 8 or 9 vav boxes on their PO.

08/23/06 – Construction Meeting Held. Job should be completed in 60 days. Triton to provide equipment, JCI to provide installation power and controls. PDF

05/09/06 – TTC compiled the complete state term contract and will deliver it to Tony’s office in the morning of 05/10/06. TEI is sending 4 copies of the contracts, spec book, and 1 copy of the JCI bid documents.

02/16/06 – Tony asked for final contract documents to be sent over so he can move the project forward. TEI will have these to Tony by 02/17/06.

01/10/06 – JCI has provided all contract documents for a state term contract. Sam compiled everything and TTC needs to review it. TTC will then deliver to Tony for use on this project.

12/13/05 – TTC has contacted JCI numerous times about submitted an actual state term contract. Al Bell will work with Joe Schutte and get one prepared for this job.

11/08/05 – TE heard back by email form Jill Williams and we will set up a quotewire job for this.

11/07/05 – TTC opened and tabulated bids for this project. See link below. The project bids came in over $25,000. Tony and TTC decided that the best way to get the job done at this

10/28/05 – Prebid meeting was held. Bids are due 11/04 to TE.

10/24/05 – Project is ready for TTC review. Prebid is this Friday at the JC.

10/14/05 – Project is ready for review and bidding.

10/04/05 – TTC & SRB have visited the site and are redesigning the dual duct boxes. We have contacted JCI about scope of Sheriff’s upgrade to use on this job. Drawings are done. Specs need to be finished and scope needs to be ironed out.


JC – HTX Upgrade

02/26/08 – Final paperwork, O&M’s and As-Builts have been sent to the County for closeout. This job is complete.

02/21/08 – TTC received the final pay application on this project and signed it immediately and sent it to Tony for closeout on this project. ThermalTech agreed to do the as-built drawings for Tony on this project.

12/05/07 – Today we worked on allowance cost and change order submissions by PHB. TTC asked for finalization and closeout paperwork from Bernie on the job. 3 allowance have to be finalized and TTC asked the County if there were any other tasks on this project they needed done with the remaining allowance money.

11/15/07 – PHB reported today that the equipment went in exactly as drawing except for some minor change in piping right above the heat exchanger. TTC said he would accept a markup on this change and incorporate it into the electronic dwgs for the job and give it to the County for final As-Builts. PHB is working on final closeout documents this week and should have them to TEI for final approval very soon, they only items they are waiting on is JCI sequences.

10/17/07 – JCI does not have the right sensors in the HTX system for their controls. We are investigating replacing them or using an old one. this needs to be resolved quickly.

10/01/07 – TTC & GD visited the site last week and reviewed the project installation. The job was very well done by PHB. GD requested that the strainers be installed in the horizontal position to alleviate water hammer. This is an arduous task for PHB but they are going it today. TTC approved payapp #2. JCI said they were almost finished with controls but had to add flex tight to the controls at the valves. Howard was going to run the system in manual to see how it works. TEI still must provide a punchlist for this project.

09/12/07 – Howard called with a questions about isolating the PRV for maintenance. TTC said it was isolated on the PID. TTC will stop down tomorrow to review with contractors and will setup a progress meeting with GD to go over design.

09/03/07 – In the past week we have approved a pad extension onto the front of the existing HTX pad layout to allow the piping to be connected to the existing piping easier. We have approved new butterfly isolation valves from the allowance monies and TTC approved payapp #01 for the project. In the field TTC approved a slight modification to the layout of the steam traps for easier maintenance and TTC and Walt discussed the atmospheric steam vent that is supposed to be connected to a tube inside the trap and came shipped separately.

08/23/07 – PHB asked about new steam control valves vs old. TEI directed them to the Add01 where hot water control valves are reused but not steam valves. Steam are new CV-27 types and 3″ in size. Howard asked about HTX loads and copper tube sheet strength. TTC asked GD and TZ for info. TTC scheduled a field trip on Friday to the building to review the progress of the project.

08/20/07 – PHB did get the new steam isolation valve installed. The Spence valve and 2″ gate valve was also replaced. Howard said that 2 of the hot water isolation valves are corroded and need to be replaced. TTC asked PHB to give us a price to replace this valves. These particular valves are not more than 5 or 6 years old as they were put in during the

08/17/07 – PHB is installing the steam valve, Spence valve and 2″ gate tonight after hours. TTC will check in with JC tomorrow on progress.

08/03/07 – TTC gave PHB requested information on the Spence valve, isolation valve and existing steam pressures. The Spence valve has unions on both sides for easy replacement. The isolation valve shows 6 inch on the dwgs and with the insulation on the piping, there is no easy way to confirm it isn’t 6″. The steam pressure is 65 psig max for normal operating pressure so PHB’s request to buy a 150 psig isolation valve will be fine. TTC & HS picked the location so that the HTX could be isolated from the system for this work.

05/30/07 – TTC contacted PHB about the equipment orders. The HTX vendor asked for more money for sizing the HTX at 10 psig but TEI emailed PHB and showed that 10 psig was called for on the schedule. This may hold up equipment ordering.

05/21/07 – TEI returned submittals via email to Bernie Tobergte today. Now comes the equipment order and scheduling. TEI will deliver hard copies to Tony and Howard.

05/14/07 – TEI has all submittal data and will review on 05/16/07 and return to PHB by week’s end.

03/14/07 – TM & TTC have contract and are ready for kickoff meeting.

02/01/07 – Contracts have been sent to BOCC for confirmation and then we will kick off the project.

01/02/07 – TEI has submitted all contract and bid documents to Facilities for Prosecutor review but as of this date, we have not had any information back on this project.


JC – Revolving Door Replacement

12/07/07 – CRA left message with Steve Jones regarding project status.

09/28/07 – CRA approved door shop drawings and sent to Southern Ohio Door. Next meeting to be held prior to start of construction.

09/18/07 – Pre-construction meeting was held. Notice to Proceed issued at pre-construction meeting.

09/04/07 – Signed contract received.

07/26/07 – CRA received minor comments back from Prosecutor. Contracts revised and sent.

06/21/07 – Bids received. CRA directed to accept and issue contracts.

05/29/07 – CRA reissued documents.

05/23/07 – Bids came in higher than estimated. Directed to raise the estimate to $85,000 and re-bid.

02/20/07 – Kick-off meeting was held at 9am in the Courthouse Room B-103. The possibility of installing remote lock down security revolving doors in the south building was discussed.

01/30/07 – CRA received the approved SOSS.

01/16/07 – CRA SOSS was authored and delivered.



05/22/09 – TEI has not been paid on this project yet.

03/02/09 – All work completed several months ago.

08/25/08 – Progress is being made on project. No Conf Call today.

08/22/08 – Permit has been issued and give to County (almost 2 weeks ago). Construction is proceeding. MAH is handling the conference calls.

08/07/08 – TEI submitted drawings for permit and copies of permit set to County and Andre for review and their records.

07/30/08 – TEI has been to the site several times getting additional field work for this project. Original load locations do not work for new kitchen equipment loads. 30 days panel load data is required. We will try to get permit on 3 day info.

09/05/08 – Permit obtained without metering. Work is in progress.


JUSTICE CENTER Domestic Pump Upgrade

08/24/10 – As-Builts converted and sent to GBBN and Bert.

08/13/10 – Waiting for final payapp and closeout. As-Built dwg is complete.

07/28/10 – Project is complete. TTC to delivery as-built dwgs to Bert.

07/23/10 – TTC is approved O&M’s on these jobs for Bert. Dwgs are being bound. Paperwork looks fine.

05/26/10 – Project is complete. We are awaiting closeout paperwork.

05/13/10 – I approved payapp #2 today and asked Gary for closeout documents.

04/26/10 – Project complete. TEI will release a no item punchlist and begin closeout.

03/19/10 – Corporate came down to investigate a blown fuse situation when opening the cover panel on the new pumps. JCI is looking to hook up the remaining sensors back to the new pump package for Howard. TEI will issue a punchlist on this project.

02/24/10 – QCM pumps have been verified in the shop for payapp. They will start on Monday at 6am and this is confirmed with Howard and Gary.

02/04/10 – QCM scheduled to begin work on 02/22/10. TEI asked for an updated schedule.

01/26/10 – Equipment is ordered and extra pump has been approved.

12/31/09 – Submittals have been approved and sent back to QCM. We still don’t know if Tony has approved the change order for an extra pump.

09/04/09 – Contracts submitted to the County.

08/27/09 – Bid open today. One Addendum was issued.

06/23/09 – Tony said the SOSS has been signed. We are awaiting approval and the PO. The existing Systecon domestic package is variable speed and will most likely be replaced with a new variable speed system that has readily available parts.

06/17/09 – Terry and Mike visited the building to create the SOSS.


JUSTICE CENTER – Kitchen Oven Replacement

10/28/10 – Project is complete and ready for closeout. New paperwork has been received.

10/5/10 – CBS will receive $2,150.00 allowance monies instead of the $5,000.00 due to using $2,850.00 for upgrades to both ovens.

09/29/10 – Project complete. Need closeout documents.

08/13/10 – Waiting for closeout docs. Payapps sent in as paid.

07/28/10 – Oven look good. Punchlist items look to be complete. Training should be complete. Ready for closeout.

07/23/10 – Both Ovens are installed. TTC sent an email punchlist to CBS who said they would address these issues on their next visit downtown. Several of the items on the list require adjustment when the ovens are off. They said it would be no problem. Oven #1 had an issue with a door opening switch which CBS fixed and now all seems to be well.

05/26/10 – Kickoff meeting schedule for June.

05/13/10 – We are trying to setup our next const mtg on this job. Equipment should be here soon.

05/05/10 – Ovens are ready for shipping. Project is ready to startup in two weeks.

04/26/10 – Waiting on Oven delivery to begin the project. Contract checked in and is ready when equipment arrives.

02/24/10 – Tony confirmed the submitted oven is approved by Aramark for this project. Other submittal data is acceptable.

01/26/10 – TEI owes CBS submittals on this project.

12/31/09 – Oven submittals have been received. TEI to review and approve, especially the prison option to be finalized.

09/04/09 – Contracts submitted to the County.

08/24/09 – CBS was the low bidder. TEI interviewed CBS and they seem to have everything included in the bid. The bid very aggressive and want to make sure they can sell the ovens when complete.

07/29/09 – This project went out for bid last Monday. Prebid 08/05. Bid Opening set for 08/24.

06/23/09 – Tony said the SOSS has been signed. We are awaiting approval and the PO.

06/17/09 – Terry and Mike made a visit to the building to look over the equipment and the installation. There is some electrical work and mechanical work to perform but the bulk of the cost will definitely be the equipment itself.


JUSTICE CENTER Kitchen Hood Suppression Upgrade

10/28/10 – Cintas has not sent in AIA pay applications yet for payment.

09/03/10 – The kitchen passed its final inspection. All contractors were instruction to complete closeout documents and send final payapps in.

08/13/10 – All contractors reported for work. Cintas is compete. DK is complete. QCE did not finish yet. Final fire alarm test will be schedule after QCE is finished.

07/28/10 – Work is ongoing. QCE was on site and hooking up fire alarm contact last night. Cintas was not on site but Ansul system appears to be 90% complete.

07/23/10 – Cintas & QCE electric will begin work next Monday on the systems both Fire & Elec. We are scheduling a fire test pm Friday July 30. We are trying to get DK to fix the Aqua Vent system by then.

05/26/10 – Kickoff meeting schedule for Thursday.

04/26/10 – County reports contracts are ready to be signed. Then we will begin.

03/23/10 – No updates

02/24/10 – Under contract review. QCE called to ask about status. TTC updated them.

02/03/10 – TTC sent BW the prebid mtgs notes for contract approval on this job. We are waiting for construction to begin.

01/23/10 – We continue to wait for approval to proceed with this work.

12/31/09 – All bids have been received. TEI has delivered contracts to Tony for approval. We are waiting on approval to move forward.

09/08/09 – Steve Ossege discussed the hood coverage with Terry Cannon. All hoods that cover grease producing equipment. This means that one hood in the JC kitchen must be covered. The other hoods can have the sprinkler heads removed. All hoods must keep the fusible links in place with the hand pull stations. The Aqua Vent bid is still under review at the prosecutors.

08/24/09/ – Daryl Meadows and Steve Ossege will provide city direction for TEI work. Our design is complete. We just need to bid out the Ansul system replacement.

07/29/09 – Kitchen Hood Fire Suppression Upgrade is almost complete. This project is set for small bidding next week. The cost estimates show the project can be under $25,00 but not by much. The drawings are complete. The specs are complete. We will be inviting fire suppression contact

06/23/09 – Aqua Vent project in contract approval stage. Kitchen Hood project in code review and design stage. Cintas has not emailed or called TTC back about code requirements vs final design of this small project. They have given this information to the County before and I suspect they don’t want to design it for us. TEI has to update the hood suppression system, the interface to the fire alarm panel, and new relays for the remaining equipment for FA signals. This might be after the cost estimates are finished for 2009.

05/22/09 – The Aqua Vent repair has been bid out and we are awarding to Debra Kuempel for about $12k. Contracts are being written. TEI sent out the bid tab and told DK to get hard copy paper work to Tony by 05/29/09.

05/19/09 – Cintas has provided the estimate and a scope for the fire protection work on this project. TTC to use for bidding.

02/02/09 – Terry and Mike are finalizing a small project bid manual for this project this week. We are planning on bidding the 3rd week of February.

10/18/08 – Approved but new hoods not required. What upgrade is left to do? TTC to inquire with Howard.

08/07/08 – SOSS submitted for approval to County.


JUSTICE CENTER – Aqua Vent Upgrade

10/28/10 – DK sent final paperwork in for approval yesterday.

09/29/10 – Last 2 change orders were sent to Debra for closeout per Tony’s direction.

08/13/10 – All contractors reported for work. Cintas is compete. DK is complete. QCE did not finish yet. Final fire alarm test will be schedule after QCE is finished.

07/28/10 – DK finished change order work last evening. TTC check it over and was on site. All but one nozzle is working properly. We need a final approval from Howard before testing with City Friday.

07/23/10 – TTC told DK they needed to be done next week with their work. Rick returns Monday for verification. Nozzles have been replaced but 3 still need fixed. TTC will try to have DK fix these when they fix the rest of the equipment on the rinse cycle.

05/20/10 – Sent Tony and Howard final change orders with prices from Debra for wrap up. Awaiting approval.

05/12/10 – Met Debra today on final change order prices. We don’t have everything we need. I will contact Tony with questions on scope of work and change order cost. Terry sent a proposal to Tony and Howard for comment.

04/26/10 – DK sent changes orders for final repair cost to extra items such as the rinse cycle piping and solenoid. TTC and TM to decide next course of action.

03/23/10 – DK reported a list of problems with the Aqua Vent operation. TTC answered several of these answers and asked the County for direction on several others. DK also asked for a sequence of the original Aqua Vents but we don’t have that.

02/24/10 – DK coordinated final work with Howard to be this week and next.

02/03/10 – DK scheduled control panel installation on 02/22/10 when Howard returns. The owner accepted one change order and rejected the extra electrical work change order. DK has not responded to the TEI email about the change orders.

01/26/10 – Aqua vents assemblies have been rebuilt at Debra Kuempel. Control panel parts have been upgraded at DK. Change orders have been submitted for new valves and shock arrestors. We have approved this work. The change order for extra electrical time has not been approved by Tony.

12/31/09 – Rick Kuethe reports the solenoid valves are in and DK will start next week provided their men are approved and Howard Seal gives him the okay to start working.


JC – 20 ton Chiller Replacement

07/29/11 – Glen reported there is no turn down on the screw compressors.

06/29/11 – Tonnage looks to be about 10 to 15 tons. Final came in 11.7 tons. TTC sent review to Tony and Glen Wilhoit for their use.

05/05/11 – Waiting on people and load count from County.

04/25/11 – TEI asked Tony and Glen for more information on the space and to mark up the existing dwgs for space and layout, then we can quickly and efficiently run the model for HamCo.

03/30/11 – No further progress yet. Tony said we might write state term contracts for him.

02/22/10 – Tony asked us to do a load calc in the Sheriff area to confirm a new chiller size for the space serving the Sheriff space for additional cooling.


Justice Center Kitchen Flooring Study

07/06/10- Attended Kick-off meeting with Bert Watts and Tony Matre.

07/15/10- Submitted SOSS to Bert Watts.

9/27/10- Meeting with GBBN, Fred Orringer (Kitchen Consultant) Jerry Vogel and Andre Tudor to discuss the logistics and operational issues for implementation of the possible flooring replacement.

9/27/10- Transcript of meeting sent to Bert Watts for clarification of scope and summarization of issues discussed in the meeting from the same day.

12/3/10- Bert Watts called to clarify questions on scope. It was determined the report would be issued to the Owner 1/18/11.

1/18/11- Final document was sent to Bert Watts electronically for review.

1/31/11- Bert Watts instructed GBBN to request color and slip resistance texture samples from the vendor for presentation to the Sheriff.

2/3/11- GBBN requested UreCon color samples from Key Resin. Key honored request to send color samples for the epoxy topcoat as well.

2/15/11- Color sample chips arrived along with slip resistance samples.

2/17/11- Color charts arrived.

2/18/11- Bert Watts sent a list of proposed kitchen equipment to be replaced in the flooring project. TEI and GBBN are to investigate possible connection and logistics costs for this equipment for inclusion into the final Study.

2/21/11- Bert Watts called to indicated the final report with modifications is due 2/28/11.

3/3/11- Delivered final copies of Flooring Study to Bert Watts.

6/28/11- Bert Watts requests SOSS from GBBN to include ThermalTech as consultant.

6/29/11- GBBN sent final scope document to ThermalTech for coordination of SOSS scoping.

6/30/11- SOSS completed and delivered to Bert Watts.


HC General Structural Services

THP completed billing of $5,000 allow for general services. Funds lasted 10 months from July 2010 to April 2011. THP directed to submit new SOSS, refer to F24.

April 15, 2010 THP requested to review possible leak and stone condition of 800 Broadway Bld. THP and County scheduling site visit.

Past Comments

THP completed site review of 800 Broadway wood deck. THP completing analysis and letter report.

THP submitted four names for badges. County review approval policy.

THP reviewed Hillcrest cracked walls. Walls are cracked and further investigation is necessary. Cracks are not a life/safety concern and cracks do not appear to be a warranty issue from settlement.

February 17, 2011 County requested THP review floor condition at Alms Doepke. Work may be part GBBN floor project, if not THP will notify the County on how it thinks we should be proceed.

February 14, 2011 THP reviewed cabin with cracked walls as requested by the County. As a follow-up THP does not think the cracks are a life-safety concern or warranty related to previous work. THP will follow up with a summary letter.

January 27, 2011 County requested partners submit employees names for background check and badges. Future access to County facilities will be limited to people with badges or escorted by badge person.

January 27, 2011 County distributed HC Safety specification section to added to all facility projects. Partners to determine location with front end documents.

January 27, 2011 THP requested to evaluate 800 Broadway interior wood deck structure for loading capacity. THP to evaluate and if work is minimal work will occur within General Service number, if work is more involved THP to discuss separate SOSS.

THP has $3,200 remaining in the hourly allowance.

Septermber 29, 2010 THP to review MDD North side structure for possible work with GBBN.

County approved updated proposal.

County reviewing initial proposal.

THP provided preliminary pricing of the building envelope assessment of Memorial Hall on 8/17/10.

County requested preliminary pricing for building envelope assessment of Memorial Hall building.

County reviewing SOSS.THP Issued SOSS F9A for an additional $5,000 on July 2, 2010.

Roughly $1500 remaining in general fund.

THP used roughly $2,800 of the $5,000 fee.

THP tracking 800 Broadway storage area floor capacity survey to general fund. (2/24/10)

THP tracking additional Masonry work from Hillcrest project to general fund. (2/24/10)

OSS approved 2/3/2010.

THP submitted SOSS on 2/1/2010.

THP needs to submit SOSS.


Hillcrest Training School Dining Hall Loading Dock Restoration

July 27, ,2011 THP issued AIA Substantial Completion Letter.

July 19, 2011 A&M confirmed (via email) punchlist items completed.

July 12, 2011 A&M submitted final Pay Application for retainage. THP notified A&M that retainage monies will not be approved until all warranty paperwork has been submitted and approved.

July 8, 2011 THP issued (via email) punchlist items to be completed by A&M.

July 8, 2011 loading dock sealant installed.

July 7 & July 27, 2011 THP requested A&M to submit warranty paperwork.

July 7, 2011 new galvanized handrail system installed and THP approved and forwarded to Hamilton County, A&M request for payment, Pay Application No 2 in the amount of $10,070.00.

June 30, 2011 THP issued Change Order Deduct to the contract, in the amount of $3,778.00. The new contract sum is $51,132.00.

June 30, 2011 sealant scheduled to be installed and THP will punch-out the project.

New loading dock handrail system fabrication completed and A&M installed new handrail system week of June 20, 2011.

A&M completed loading dock bumper installation and installed temporary handrail system on May 27, 2011. Hillcrest facility began using loading dock May 30, 2011.

THP issued email on May 24, 2011 for A&M to install temporary handrail system once loading dock bumpers are installed. A&M to be completed with installation of loading dock bumper and temporary hand rail by May 27, 2011, so loading dock may be used by Hillcrest facility on May 30, 2011.

On May 23, 2011, THP approved and forwarded to Hamilton County, A&M request for payment, Pay Application No 1 in the amount of $38,068.24.

Hand Rail steel shop drawings approved (with exceptions noted) for fabrication on May 12, 2011. A&M expects for fabrication and installation of hand rail week of June 13, 2011.

Field Report No 002 issued on May 9, 2011.

Loading dock slab placement took place on May 6, 2011 with THP personnel on site to witness placement. Weather delayed the original April 28, 2011 placement.

New louver and new loading dock bumper approved to be purchased and installed out of project Cash Allowance via Change Order No 1, dated April 20, 2011, in the amount of $1,222.00. Total project budget remains unchanged.

Hamilton County requested price quote from A&M for a new louver and loading dock bumper, instead of reconditioning these existing items as part of the original scope. Pricing came in within the projects original Cash Allowance which Hamilton County chose to accept. THP issued Change Order No 01 on April 20, 2011 reflecting this modification.

The loading dock slab placement is estimated to take place on April 28, 2011. Weather pending.

April 13, 2011 A&M formed and placed concrete for the loading dock steps. THP reviewed the steel and formwork prior to concrete placement.

April 6, 2011 A&M formed and placed concrete for the loading dock walls. THP reviewed steel and formwork prior to concrete placement and issued Field Report No 01 for both concrete placements.

Submittals not complete as of April 25, 2011. Waiting on resubmittal of galvanized steel hand railing for loading dock.

April 1, 2011 A&M completed loading dock demo while maintaining emergency egress.

March 30, 2011 A&M mobilized on site and began project work.

March 23, 2011 THP held a pre-construction meeting on site with Hamilton County and A&M Contracting.

March 23, 2010 preconstruction meeting occurred. Demolition of loading dock to begin March 31 and closure of the dock will last 6 weeks.

February, 22 2011 A&M Contractors passed background checks and work can proceed. Bert has requested a schedule and submittals to be finalized.

January 27, 2011 A&M Contracting had a number of employees not pass the background check. Work cannot proceed until A&M Contracting project team has passed the background check.

Need to determine if project will occur in Spring or this Winter. If work occurs this winter a change order for heating will be required.

Submittals continue to arrive and be returned; last submittal return on December 7, 2010. THP has requested a number of additional items to be submitted.

October 26, 2010 THP returned initial submittals and requested preconstruction meeting.

October 20, 2010 Contract was approved and PO issued.

County reviewing contract for final approval and execution.

September 16, 2010 THP provided updated draft AIA contract per comments from County.

September 2, 2010 THP provided draft AIA contract.

County reviewed project bid.

August 10, 2010 THP recommended A&M Contracting Ink. based on the submitted bid. THP has no experience with A&M Contracting.

Bids received August 5, 2010. A&M Contracting Ink. submitted only bid. Bid was within 10% of published estimate.

Bids are due August 5, 2010.

THP plans to issue final addendum by July 29, 2010.

Pre-bid meeting held July 20, 2010.

THP updated documents and issued documents for bid on July 12, 2010.

County Responded with minor ITB comments THP to update documents.

THP Issued updated ITB documents on June 7, 2010 for review.

THP anticipates finalizing and sending Updated ITB documents to the County on June 1, 2010.

County finalized review of ITB documents and sent comments to THP on May 21, 2010.

THP finalizing bid set; anticipate release 4/5/2010.

THP complete design review meeting with County on 3/12/2010

Project anticipated for summer of 2010. (2/24/10)

Waiting on Owner comments to complete final documents. (2/24/10)

THP submitted owner review documents on 12/21/09

THP provided plans for an emergency deck for loading dock access on October 22, 2009.

THP submitted an SOSS for approval on October 23.

THP completed initial site visit with County and recommended closure of loading dock.


Hillcrest Training School Cottage Slab Review

Current Comments (5/26/2010)

Project complete and closed out.
Past Comments

L&T submitted closeout paper work. THP & L&T revising warranty language in submitted warranties; all other paper work acceptable.
L& T completing final project closeout paper work. (2/24/10)
L&T requested a project closeout, so additional grading work will not be accepted.(2/24/10)
All punch list work including additional interior repointing at interior/exterior wall joints reported to be complete. THP to review next week. (2/24/10)
All work complete except punch list. Punch list published and all work except grading except grading expected to be complete by February 19.
Contract was coordinated with the prosecutor’s office and final copies were sent to Hamilton County for review by the commissioners on September 14.
August 17 THP recommended Langenheim.
Bids received August 14 and Langenheim & Thomson was low bidder within the project budget.
Project in bid process with pre-bid meeting completed on July 28, 2009. Bid due August 11th.
Design development documents anticipated to County 6/26.
Partial front end specification sent to County for review on 6/22.
Kick off meeting completed and field survey’s complete.


Hillcrest – Security System Upgrade

Project Status

Project is Complete. All contractors have been paid.


11/11/04: Zeiser Pay Application #6 to Sharon via Courier.

11/04/04: Zeiser Change Order #4 for Punchlist items 1.19 and 1.20, new door hardware, in the amount of 690.00 typed and courier.

LaForce went to site today to complete punchlist items hardware repairs. However, Chris was unable due to audit, to show them location. Chris could you please coordinate with Zeiser to a date you would be available to show them locations.

Cottages grandmaster key – Contractors are working on resolution to be presented.

Post caps for doors stops are in and ready to be installed.


ICS to complete punchlist including rewiring system. Please update us to the progress.

12-07-04 – New master keys are on site for cottages. Zeiser to coordinate work with Hillcrest

01-04-05 – Zeiser and Laforce to complete cottage level master rekeying. Work to start on 01-05-04.

04-19-05 – Cottage level master rekeying complete. School has accepted suggested solution for fixing the bathroom keys. Laforce has install the proposed fix and is returning to make repairs.

04-2-05 – Bathroom rekey complete however one issue still exists. Dining hall locks have a total of four cylinders and only two where replaced. The second set can be defeted but according to Dave Jones with LaForce this is not possible because of the type of lock. They proposed replacing the sets and provided pricing accordingly. County review cost and felt that it was not their responsibility because this proposed fix was only a compromise that remedy the situation. Upon conference call with Diana, Mike, Jeff, Chris and Dave fix was agreed upon. Dave with LaForce to complete week by 05/06/05.


Hillcrest – Modular Building / Addition

08/08/05- C+R sent a preliminary project report to Hillcrest. TE to review and comment and provide updates to Spencer for inclusion.

06/16/05 – C+R, TE and HCDF had a follow up meeting at site to discuss layout issues. Spencer distributed possible floor plans for review. TE went over utilities that were underground and reported that electric, gas and water are all near the expansion area.

01/19/05 – Second Field trip to discuss cost and scope.


Hillcrest – Fire Alarm Replacement

12/26/06 – Project has been archived by TTC.

11/09/06 – Final permit was accepted

11/03/06 – TEI is working on closeout documentation. Final devices have been installed. Final approval has been obtained from Hamilton County Building Department.

10/27/06 – Final devices have been installed. Ginter to set up to get final approval

09/29/06 – Waiting on approval of revision from Hamilton County Building Department

09/22/06 – We have passed and approved (signed off and everything!) on 21 buildings out of the 22. Each building was permitted separately. The one building (grounds keeping) is on hold until the inspector sees the 501 model number heat detectors installed in the or we submit an engineering change that corrects the permit drawings with the 503 model number. They are in the process of getting the paper work together and will forward it to you. As far as the grounds building we are not being told that we can not use the building. Also we don’t believe they occupy it all day, mainly just storage and prep for the grounds care equipment.

9/12/06 – 21 buildings passed fire alarm inspection grounds keeping building approval on hold until heat detector issue is resolved.

9/6/06 – IPS adjust strobes in cottages that were coming back with a trouble condition. They were set to the wrong setting. Ginter adjusted devices.
8/30/06 – Hillcrest performed a drill test and a few troubles came back on the system. Chris and I contact IPS and Ginter to get trouble condition resolved.
8/25/06 – Hillcrest will be unable to have the cottage door delay egress maglocks repaired prior to the 8/29 test date. Test is rescheduled for 9/12
8/21/06 – Complete rewiring of audible devices in cottages so the do not sound when fire alarm system is in alarm.

04/06/06 – Mike Stiles has requested a change in the meeting schedule for 04/10 as he is out of the office on that day.

03/06/06 – Ginter email TEI that the county was requiring separate permits for each building. TEI call the Building department and clarified the situation. The county required 22 permit applications and 6 drawings. The minimum fee per permit is 257 dollars. Ginter indicated that they did not have an amount this large covered in there bid. TEI requested that Ginter submit a total for all cost so far in the project: Some of the additional fire alarm cable, additional permit fees and deletion of 10 in-ground boxes.

03/03/06 – Ginter Electric provided submittals, TEI approved submittals.

02/28/06 – Construction meeting #1, Please see attached.

02/03/06 – Ginter submitted pay application #1, TEI approved.

01/20/06 – Preconstruction at Hillcrest.

12/16/05 – County reported that this project has been approved. Wayne Trent will be project manager on this job and therefore the project is being moved to his task sheet.

02/01/06 – Ginter Electric started field investigation.

12/01/05 – Contract was approved.

11/01/05 – TEI prepared contract for Ginter Electric, Alternate #5 was accepted.

10/31/05 – Bid opening held, apparent low bidder was Ginter Electric.

10/24/05 – Issue addendum which delayed bid opening to October 31st. Addendum was the same form Re-bid.

10/14/05 – The rebid II prebid was held. No contractors showed up for the rebid. This might be expected because everyone has already been on site for 2 prebids already on this job. There can still be and will most likely be bidders on this project.

10/13/05 – The rebid has been rejected due to bond issues again and estimate tolerances. The second rebid is on the street with a revised bid opening of 10/25/05. MS has all the revised documents on the street. The prebid is 10/14/05 at 10a.m. MS will attend for TE.

10/11/05 – Project was issued for rebid.

10/05/05 – Bid was rejected by the prosecutors for basebid plus alternate because the awarding contractor complete the bid bond for only the base bid amount. TEI advise that alternate 7, which was cabling in the ground in conduit plus a spare, was very important. Since the contractor filled in the bid bond purchasing would only allow base bid. The next lowest contractor was reviewed and but there base bid plus the alternate was greater than 10% of the advertise amount. The county decided to rebid the project.

9/02/05 – Contract completed for AC Electric, base bid plus alternate 7. Submitted to Diana for approval.

08/25/05 – Bid opening held, apparent low bidder was AC Electric.

08/10/05 – Project prebid was conducted.

08/3/05 – Project was issued for Re-Bid.

07/01/05 – Diana, Gerald and TEI decided to hold off on rebiding project to gain more interest. MS contacted several contractors to find out marked conditions. The consensus was that the school at released a lot of work to be complete by the beginning of school and the contractors where very busy. It was also determined that breaking the project into several alternates would be required to make sure the project can be complete within budget.

06/28/05 – TTC and MS discussed the bids. A clarification on the addendum split up the labor versus the equipment. The bid is not over the original estimate of 298,000 however it is over the revised estimate. TE has to discuss with Diana when she returns our call. MS believes that this project should be rebid as the bids are simply too high for this type of project.

06/24/05 – Diana emailed TE today to say that the bids came in over 10% of the budget. TTC and MS to discuss with DES on Monday in a phone call.

05/19/05 – I spoke with Andy Boyd at IPS about the project. I discussed the three state term options with him which are: A) Purchasing material only and bidding out the installation to an electrical. IPS (or competitor) would be subcontractor.B) Purchasing material and IPS to set up the system and bid out electrical labor for the installation.C) Purchasing the entire project from IPS.Option C has changed because Purchasing indicated that they can only do the project if they do not use any subcontractors. I presented this to IPS and they gracefully turn this option down because they do not have the resources to complete a project of this size on there own. Option B was discussed. I ask the question, what advantage is it to the county to proceed. IPS indicated the main advantage was for them and that the cost under state term and bid day would be about the same. IPS and I agreed that this was not the best option for the county. Option A was discussed and believed to be the best solution. This makes the bidding process cleaner and only has one contractor to manage. The above is a recommendation but if you have any comments or suggestions please let me know. I would be happy to proceed the manner you prefer.

05/12/05 – TEI contact JoAnn with purchasing to discuss doing the project under state term. She indicated that IPS could not have an electrical subcontractor because the amount of the contract was to much and must be bid out publicly. Prepurchasing material was acceptable.

05/10/05 – TEI presented final drawings and Edwards material pricing. TEI indicated that the material price provided by Edwards was in budget with all the options requested but not with alternates. TEI indicated that Alternates should remain to insure the budget. TEI indicated that there where 3 way to bid the project: 1) Prepurchase material and bid out the installation. 2) Prepurchase material and ESI setup of system, bid out electrical labor to install system. 3) Purchase entire project under state term. Hillcrest indicated that they where interested in purchasing the entire project though state term because of the difficulty of trenching on the site and the responsibility to the electrical contractor.04/18/05 – TEI present near complete drawings written around a full system replacement of any manufacturer. Hillcrest review vendor options and indicated that the two favorites where Simplex and Edwards. Overall Edwards offered a security solution that could handle the need of the fire alarm system and replace the antiquated access control system. State term options was discussed and Hillcrest indicated the we receive compare the price for Edwards with almost all requested options and compare to original budget. Items to be included is: Voice System, Access control panel for each cottage, Access control software, Data base software, Computer based graphics package. Alternates for the project shall be: Voice system in lue of Tone system, Additional access control module for each cottage, Access control module for gates and upgrade spare conduit in ground with honeycomb type.3/30/05 – Vendor presentations.

03/29/05 – Vendor presentations.

03/09/05 – TEI complete field take off and composed 4 options with associated cost estimates to proceed with project. Cost estimate #1 was to completely remove existing system and replace with a new tone system of any manufacturer. Cost estimate #2 was to completely remove existing system and replace with new voice system of any manufacturer. Cost estimate #3 was to upgrade existing system Fire Alarm Control Panel with the same manufacturer, add additional devices as required per NFPA and replace all existing wiring. Cost estimate #4 was to upgrade the Fire Alarm Control Panel with the same manufacturer, add additional devices as required per NFPA and replace some of the existing wiring. Hillcrest to review options and requested vendor presentations be conducted to aid the school in there decision. TEI to continue with drawings as far as possible until decision is complete.

02/22/05 – Project Kick Off meeting. TEI to start field take off and review options. Original cost estimate presented was for a complete system replacement. TEI suggested reviewing the existing system and look at possible options.


HILLCREST – Electrical Upgrade Phase II

03/29/07 – TTE received final pay application and the ‘CONSENT OF SURETY COMPANY TO FINAL PAYMENT’ form from QC Electric. All of the final documentation will be forwarded to Hamilton County Facilities the second week in April of 2007. There are no open issues on this project.

03/23/07 – Closeout meeting was held on site today. There are no open issues for this project. Final documentation (O&M’s, As-Built, Final Pay App., Close Out Documentation) have been turned over to TTE. An Allowance Clarification letter needs to be written and signed before the final pay application can be processed. This letter will be sent to Queen City on Monday, March 26, 2007. When TE receives the signed letter back from QC all of

these materials will be turned over to Hamilton County Facilities. Dick Kuhn, Mark Donnelly, Bert Watts, Chris Hohmeister, Phil Lambing and Nick Richmond attended this meeting.

03/20/07 – Closeout meeting is scheduled for this Friday. Project is paid out except for retainage.

03/15/07 – I believe this project is ready for closeout. TTC to confer with PL and wrap this project up.

02/19/07 – Project is complete. TEI is starting closeout phase on this project. TEI will coordinate with Mike Jackson on closeout on this project.

11/10/06 – Concrete pads have been poured. ID studies and receptacle installations have been started in cottages. Project is still on schedule for December 1 shutdown.

10/27/06 – Contracts released, and preconstruction meeting was held. The schedule was approved. Coordination meeting with Duke Energy occurred 10/25/06.

09/22/06 – Still waiting for contracts.

09/15/06 – Bert called back to say contracts have not been received yet. Contracts are in review stage. QCE will not be on site next week but TE will coordinate a const mtg as soon as the paperwork is in place.

09/15/06 – Construction to start next week.

09/15/06 – Bert Watts said Contracts were released, ThermalTech has not yet received them.

09/08/06 – County reviewing contracts and insurance.



01/28/08 – This project was added late because it started before the website. After 4 years of futility, the County and ThermalTech finally agreed on a solution to the cooling and heating problems in the Security Building. Tony bought the new computer room ductless split system and County trades installed the unit. It is currently up and running and the space is doing great! EDI came to the site on Monday, Tuesday & Wednesday to install the AC unit. We had some unit problems as we received an AC unit and not a heat pump. Jeff did a great job returning the unit and getting the heat pump for us which amazingly was in stock at Habegger. We installed a 4-ton fan unit and 3.5 ton heat pump so that I could ensure full flow into the building since the ductwork sizing has always been a problem. The new unit and heat pack are up and running from the Carrier thermostat on the wall and the space has been 72, 74 and 77F when called to check on. Mike Jackson and I discussed the cold feet issues today and he is comfortable with the unit providing enough heat and said the space was 1000% better. I will recommend some radiant heating solutions to Hillcrest for their feet but I don’t believe that TE should have to pay for those. I believe this project could finally be complete.



04/23/08 – QCE had to fix the same AHU unit again when they cleaned up some of the school panel wiring. Project is nearing completion.

04/11/08 – MAH released the Const Mtg 05 Notes to the project partners.

03/13/08 – (3) full size sets of drawings and a letter stating that no loads were being added and that the feeders and panels are of sufficient size were delivered to IBI today. QCE should be able to obtain a permit with these actions.

03/06/08 – Today’s construction meeting has been rescheduled to 03/17 at 2pm. We also discussed the revised schedule for the project which I updated in a pseudo meeting minute memo in PDF format. TEI will process the IBI letter for no load increases during this job for QCE to give to IBI to get permit squared away. Bert also asked for a revised scheduled update which we have included in the memo.

02/26/08 – QCE called today to tell us that IBI denied our permits for this job. They now instead want TEI to provide a letter saying no loads have been changed or will change during this project and we are doing a direct panel for panel replacement. A shutdown is going to be scheduled in late March for school panelboard replacement(22 or 23).

01/09/08 – Today we decided that the pre const meeting notes were correct and the project completion date for this project is March 1, 2008. QCE will do and update project schedule with delivery details for TEI to approve and give to the owner very soon.

01/07/08 – MAH released the Const Mtg 02 Notes to the project partners.

11/28/07 – QCE provided an update to Mark Donnelly today with the following information: The ID portion of the project is going well. The (4) cottages are complete. Chapel, John Morrow building, Goundskeeping building, & Pole Barn are completed also with the ID. We still have to change locks and install the new locks on certain panels. Administration building is approximately 75% completed, the School is approximately 50% completed. Some work has been done in the Dining building, Rec building, & Gym. We have not been in Security yet. The transformer pad for the school transformer was poured 11/27/07. All Panels have been measured & verified. The order has been placed, but I don’t have delivery dates yet. As soon as I have the delivery dates I will put a schedule together and ask for a meeting to go thru what has to happen for the shutdowns & installation of the panels. TTC setup a const mtg for 12/10 at 2:30pm to follow up with panel order and schedule.

11/12/07 – TTC released the Pre Const Mtg Notes to the project partners.

08/31/07 – TEI still awaiting word from HCFD. This project remains in the Contract Phase.

08/17/07 – TEI awaiting word from HCFD in order to move forward with this project.

08/10/07 – TEI spoke with QCE about their bid. They were comfortable with the bid so TEI wrote the contract and forwarded to the Hamilton County Facilities Department. TEI awaiting word on signed contract so this project can move into the construction phase.

07/31/07 – TEI has obtained a copy of the bid package and is reviewing the information. Dick Kuhn of Queen City put together the bid; he will be back in his office on Wednesday, August 1. NR will contact him in the morning to discuss the bid.

07/27/07 – Today was the bid opening. NR attended for TEI. QCE and Ginter are the low bidders. QCE was at $138,000 for this project. TEI will check the bid to make sure everything is covered.

06/29/07 – No change.

06/22/07 – TEI obtained an ITB #099-07. This project is set to be issued on June 26, 2007. The pre-bid meeting is scheduled for July 3, 2007.

06/15/07 – TEI conducted a site visit on June 13, 2007. The final design strategies have been discussed and this project is moving into its final stages before being issued for bid. TE plans to issue this project for bid by next Friday, June 22 of 2007.

06/08/07 – This project was reviewed in a meeting between TEI and Mark Donnelly. There will be a set of drawings and a cost estimate delivered to Hamilton County Facilities Department this afternoon (June 8, 2007). A meeting will be held on site next week to review final conditions and then will go out for bid.

06/01/07 – NR will have a review set of drawings, specifications and a cost estimate ready for Mark Donnelly to review by the end of next week (June 8, 2007).

05/25/07 – Project is moving along as expected.

05/22/07 – Project review meeting was held today. Various points were discussed as described in the meeting notes. TEI is moving forward with this project while keeping the points discussed in today’s meeting in mind.

05/21/07 – TE has completed all field work. Drawings are 98% complete. A meeting to discuss the status of this project is scheduled for May 22, 2007 with Mark Donnelly.

04/07/07 – TEI has completed the device identification portion of this project and are currently inputting the data into CAD. The panel replacement portion of this project is moving along as expected.

04/20/07 – TEI has been on site working on the device identification portion of the project. The only areas left to complete are north and south wings of the School Building. TEI will be on site today, April 20 2007, to work on the panel replacement portion of this project.

04/13/07 – TEI will be on site beginning next week to work on the electrical device identification portion of this project. Design for panel replacement is moving along as expected.

04/06/07 – Project moving along as expected.

03/28/07 – Kick-off meeting held on site. Buildings to be included in the Device Identification portion of this phase are Administration, School, Chapel, Security, Maintenance, Storage, Dining & Recreation and Gym & Pool. Efforts are aiming for August of 2007 for the shut-down and transformer pad replacement to coincide with the week that school is out of session. Attendees of this meeting include: Mark Donnelly, Bert Watts, Chris Hohmeister, Phil Lambing, Nick Richmond and Ron from Facilities. The next meeting is scheduled for May 2nd at 10:00 on location.

03/23/07 – Design meeting to be held the last week of March 07.

03/20/07 – TE to assist in deciding scope on this project and settle on which buildings are left to do on this site.

03/15/07 – SOSS received and posted to the website. A kickoff meeting will soon be scheduled.

03/14-07 – NR created project journal and filled in back data from previous meetings. Project awaiting signed SOSS to be returned to TE.



03/13/08 – TEI awaiting scheduling of suppression system test. Will witness on request.

03/09/08 – TTC sent closeout paperwork to DK and request for Credit on project.

03/05/08 – DK tested the electrical portion of the system successfully today. All relays operated as required. A test of the fire suppression system will now be scheduled by FES. This will need to be done when no cooking is occurring under the hood. A tank of ordinary air will be connected to simulate the suppressant.

02/26/08 – DK reported that this project is complete and ready for final fire alarm test. This is being coordinated through Hillcrest and MJ. TE may attend this final fire alarm test.

01/28/08 – We have received Payapp 2, O&M’s and an approved permit for this job from DK. MJ requested and TTC asked DK to setup a FA test showing the working components of the system. MH has the O&M’s to approve for this job.

01/08/07 – Earl Moon approved the project yesterday and TTC spoke to Kevin Broxterman today about the allowance usage and remaining allowance monies on this job. TEI will schedule a punchlist visit to the site and close out this project. TTC tried to contact Huntwork to do the punchlist today while he is on site with another job. TTC instructed DK to do as-builts and give them back to TEI for inclusion into the CAFM system. TEI will send an email outlining all the closeout procedure documents for this job. MD has requested closeout before next Monday. TEI finished the punchlist today and distributed this to everyone.

12/12/07 – MAH talked to Kevin Broxterman and the disconnects are scheduled to be installed on Friday. He asked if there had been any discussion about his using some of the allowance for this work. I directed him to Bert Watts while Mark D is out.

11/30/07 – The electrical inspector has an issue with the disconnect location because of access to the them over the equipment. MH went to the site today to review and find a new location based on our discussions with Kevin B at DK and Mark D at the County. We want to find a new location on the first floor for the disconnects. MH located a space behind the wall of the existing equipment. We will have to relocated the existing fire extinguishers, sign and fire blanket so everything can share the space but it appears very do-able. TEI is sketching up the photo right now for changes. The inspectors wanted the disconnects in the basement but the owner does not. DK submitted their first pay app on the job and TTC approved it for $8941.60.

11/13/07 – Debra-Kuempel has reported updates as well as Mike Jackson from Facilities. Debra says they should finish tie-ins this Thursday. Greg will schedule with IBI and IPS for inspections and fire alarm tie-ins for the end of this week or early next week. Greg of DK reported that they had no problems or questions at this time. MJ reported that the tie-in in the basement Rec Ares is complete. Also that the new disconnects for the steamer and the oven are complete. DK scheduled a shutdown last Friday on 11/09/07 and all went well. DK has not been above the ceiling in the crawl space yet nor are they on site on today. To TEI, this project seems to be going well. TTC will check progress with MH at TEI to make sure we are on target for this job and possibly schedule a field trip this week to check on progress. This field trip may be combined with meeting EDI at the site to discuss the AC installation and also for TEI to review the computer room installation that Facilities is currently working on.

11/12/07 – TTC checked in with Kevin Broxterman. Equipment is ordered and DK will be on site this week not last week to do the work per the schedule agreed to with County Facilities.

10/12/07 – Pre Const Mtg held and meeting notes distributed to all attendees. Super timeline crunch has been averted by Facilities passing the accreditation testing with the understanding that this hood system is being upgraded right now. DK will order parts and scheduled work very soon.

08/13/07 – TTC interviewed Debra Kuempel and confirmed their bid price was good. TEI wrote the contact and TTC marked it up. It will be in County hands on Tuesday for approval. The PO can be acquired at the same time. TEI will apply for the IBI permit on Tuesday as well.

08/10/07 – This project was sent out to bid on August 2, 2007. A mandatory pre-bid meeting was held on site on August 3, 2007. This project opens for bid on September 3, 2007 and is to be completed by September 15, 2007.

07/27/07 – This week TTC & MH began working on the final design this week. MH visited the site for fieldwork and created the construction drawings. TTC is hoping that

05/14/07 – SOSS has been received and logged. TEI asked for a kickoff meeting around 05/28-05/31. We will await MD direction on this.

04/27/07 – TEI awaiting signed SOSS.

04/17/07 – MJ called to say that fan belts were replaced and airflow in the hoods is up significantly. One motor has been replaced previously. TTC is writing the SOSS for the report.

04/13/07 – TTC & TA met MD & BW at the site for a walk through and looked at existing conditions of the kitchen hood and electrical requirements. TTC found that the EF did not have any nameplate date on them. Airflow to hoods seemed low. MD requested traversing the hoods to confirm airflow.


HILLCREST Electric Panelboard Replacement

10/16/08 – Final payapps & O&M’s delivered and sent to Bert Watts. Consent received and mailed to Bert.

08/26/08 – Closeout meeting finished. QCE submitted two change orders for this project. TTC sent change order request to Bert for this job. Closeout documents turned over to TEI for approval.

07/30/08 – TEI has finished punchlist and distributed to contractor. Contractor to finish punchlist, submit all final paperwork and prepare final O&M’s for the project.

Allowance: $15,000

Approved change orders 1 and 2. Should be sent today. Then final pay app can be submitted.


HILLCREST – Oak Cottage Wall Investigation and Hillcrest Maple Cottage Wall Investigation

3/29/12 – Langenheim & Thomson install block walls per THP details.

County reviewed THP proposed test of replacing block wall with reinforced stud wall system. Due to security concern, wall is to remain a block wall system. THP to complete test installation of block wall system with control joints.

3/29/12 – THP issued preliminary summary letter and will follow-up with detailed letter report.

3/30/12 – THP issued final report with all test repairs complete.


HILLCREST – CT Cabinets Upgrade

09/28/10 – SOSS received. Design documents are complete and ready to be sent to County for review. TEI QC complete. Should be bid next week.

10/28/10 – TTC has to release for bidding. The job is ready to bid.

01/24/11 – DeBra Kuempel won the bid and we turned in a contract for approval. I am not sure of the current status. Confirm with Bert when the pre const mtg should be scheduled. DK should have everything turned in for approval.

03/30/11 – The contract is either approved or close to approval so we can start the work very soon.

04/27/11 – Bert sent the contract to Debra Kuempel and let them know that background checks should be gotten and given to Bert to start the project.

05/06/11 – Bert asked about background checks for DK.

06/29/11 – Kickoff meeting was held Monday. TEI is authoring the meeting notes. Work scheduled to be done late July or early August.

07/28/11 – ThermalTech returned all DK submittals on the project. No exceptions.

08/22/11 – Work scheduled for early September. DK sent a preliminary timeline for Bert to share with contractor. Power permit is being coordinated with Duke.

02/28/12 – ThermalTech is in receipt of the closeout documents that you have submitted for this project, along with a request for final payment. Per the requirements of the project manual, there are items that must be submitted prior to processing the final payments. Upon receipt of the following items, closeout documents can be processed: Notarized statements of warrantee, AIA Document G706 contactors affidavit of payment of debts and claims, AIA Document G706 A Contractor’s affidavit of release of leans, and AIA Document G707 consent of surety to final payment.

6/26/12 – Mike Barth spoke with Tom Helmes of DK, indicating that the Notorized warranty is the only item that is holding up the closeout package. Tom indicated that he would drop off the warranty on 6/26 for final processing. Upon receipt, ThermalTech will process the closeout documentation.


HILLCREST – Underground Fuel Tank Replacement

09/28/10 – SOSS received and project out for bidding. Pre bid is Oct 4.

01/24/11 – Pre Const Mtg schedule for next week with Valor Construction and Indrolect electrical company. The County asked us to finalize a temporary design for power to the maintenance building that broke during the winter.10/28/10 – Bids are opened. Valor has passed their pre award meeting questionnaire. Contract has been sent to Tony for approval.

02/22/11 – Submittals are approved for the tank and sent back to Valor for construction and ordering. Indrolect submittals are approved and sent back. Indrolect has additional background checks because they are going to bore the new conduit under the driveway on this job. Valor has to come to HamCo to do their backgrounds checks. We notified Valor by email of this requirement.

03/30/11 – Contractors are on site and beginning work the first week of April.

04/27/11 – The old tank is out. The new tank is in. Valor ordered the wrong sensors so they have been reordered. IBI rejected the disconnect install so that has been reworked and TEI will pay for that and submit a change order for the all the engineering changes on this project under the new contract for Tony to approve. If everything goes well Indrolect and Veeder-root rep should have the Veeder-root and TopKat working on Thursday. Valor will arrive on Friday to put the interstitial sensor in. On Monday, Indrolect will arrange the final inspection and by next Tuesday everything should be complete.

05/02/11 – Valor has completed the tank install. All sensors are installed and wired. Indrolect is complete with all wiring and hookups but will be back to backfill the grounds this week. We have IBI permit tag in place. Reliable is on site programming the TopKat with Denton. Theoretically, the system is ready to use if needed. I believe all permits and inspections are complete. Now it is down to getting Veeder-root programming complete and making the TopKat communicate with the pump properly and that has to be coordinated with Denton and Reliable I believe.

05/09/11 – The new software is on it’s way and Gene from Reliable is on site working on comm issues. TTC spoke with Denton this morning.

05/13/11 – Michael Denton asked if any word on the software for the ball vent yet. TTC replied that Valor had been asked how to remove the ball valve with no response, but the software was on it’s way and should be on his computer before the 12 day test runs out IMO.

05/25/11 – Valor must complete punch list items and train on software.

05/31/11 – Valor will finish the punch list on the tank next week and get factory trained representation training for Denton. He promises to get this work done ASAP. Closeout documents are in process and may be at ThermalTech this week.

06/14/11 – Indrolect instructed to use May 17, 2011 for date on final paperwork. Substantial completion will be typed up by LKV.

06/16/11 – TTC asked Liam for an update on the following punch list items: Has the lip at tank fill been fixed? TEI will apply laminated tags for Valor if they can get the rest of the project completed. The ball valve has been removed. Factory training on the TopKat system needs to be done, asked if Gasboy will send someone. Local rep says he will get everything up and running.

06/21/11 – TTC and Liam spoke and the closeout packet is coming to TEI for final closeout. Liam believes the punch list was left with Denton and is complete except for factory training which Liam is currently working on. Denton had said that the only thing they thought Valor did last week was remove the ball valve, but Liam believes that Loren completed everything but training. Some confusion on who has the punch list and how much of it has been checked off completed by vendor and witnessed by Owner. If everything is complete, TTC will make a punch list visit and sign off to completion and send Tony all that documents that Valor is sending to TTC for approval.

06/29/11 – We contacted Valor again this week to try to get them to rectify the punch list issues.

07/05/11 – Valor called back today and said they would be down this week for fix the water leaking into the fill sump issue.

07/11/11 – TTC contacted Valor and Indrolect about closeout documentation.

07/16/11 – TTC Received a new payapp from Indrolect. Signed and sent to Bert Watts with closeout documents.

07/29/11 – Valor has completed punchlist work on the tank. We still need that extra training from the factory. Some water got into a fill sump again Denton noted. ThermalTech has to investigated.

08/08/11 – Valor payapp #2 for 15690 arrived but TEI believes the county will want to put this on hold for now, until water issues and training are complete. Tony Matre stated to hold the invoice until items are resolved.

08/22/11 – Uploaded Closeout Documents for Tony. PDF

08/30/11 – Still having some issues at Hillcrest with water in the fill and the training portion of the TopKat System. TTC would like Valor to check with Mike Denton to discuss these issues and a timetable to rectify. TTC should be copied on all correspondences.

08/31/11 – TTC has sent photos of the water issue to Christ Batch at Valor. Chris will also check into the training that was scheduled for 8/18 and they thought was completed. Chris will get back with TTC soon to discuss options.

09/08/11 – New OPW spill bucket has been ordered by Valor, should solve the water issue.

09/20/11 – Valor has replaced Emco bucket with O.P.W. The O.P.W. bucket lid is much heavier and does have a gasket. It is a nice install. The plastic was removed from the new concrete and the lid opened. The inside of the bucket was damp, but this could be because the work was done in the rain.

09/21/11 – Mike Denton says that the factory authorized training has not happened yet. It was scheduled but was never followed through with. This training is on how to use the TopKat software and reporting. The initial installer had never worked on the system before. TTC has asked Valor to please check on this since it is the only open item remaining. The sump pump was dry as of this morning.

09/27/11 – The supplier e-mailed Valor Contracting and stated that the training issue would be resolved within 24 hours.

10/05/11 – TopKat has set training up for October 6th between 9:00am-10:00am through Chuck Storey.

10/06/11 – Training by TopKat was completed today.


Hamilton County EMA

THP work complete. Refer to TTE for project status.

December 3, 2010 THP provided a summary review letter of siren pole wind loading review.

November 18, 2010 Thermal Tech requested siren pole wind load review. Work to proceed hourly.


Courthouse 2010 Lighting Repairs

Refer to TTE for final closeout status.

March 7, 2011 THP published review of completed stone work.

Work complete in late February.

Once approved, THP will close out the project.

March 3, 2011 THP completed review of work and sealant work is acceptable. THP did not see the installation, but received written confirmation that backer rod and shims are installed in the joint.

January 27, 2011 County approved contract. Ram reviewing contract with corporate office. Ram insurance paperwork submitted and approved.

December 28, 2010 Preconstruction meeting occurred.

November 4, 2010 bid received. Thermal Tech and County processing bid.

Prebid meeting completed October 28, 2010.

Prebid meeting scheduled for October 28, 2010.

October 7, 2010 THP sent draft documents to TTE for review.

THP anticipates returning the documents October 7, 2010.

September 28, 2010 THP received draft documents for THP review and design of sealant repairs.

September 2, 2010 provide an email proposal to Thermal Tech to provide consulting services.

August 26, 2010 County, Thermal Tech and THP had kick off meeting to review lighting repair project. THP requested to provide assistance with sealant joint repairs. THP to be a consultant to Thermal Tech.


Courthouse 2010 Service Drive Water Infiltration Investigation

Project updates included with F12.

August 26, 2010 Tony Matre requested option #2 repair (full repair) be completed. To expedite 2010 repair schedule work to be included with F12 as an add alternate. THP to bill design services to remaining hours on F13; then review additional service proposal.

THP delivered report August 25, 2010.

Past Comments

THP delivered report August 26, 2010

THP to complete report.

SOSS approved June 8, 2010.

THP submitted SOSS for review on May 11, 2010.

THP and County completed kick-off meeting.


Courthouse – Exterior Lighting Project

04/11/05 – TE still trying to get final paperwork from Mike Cavanaugh. Once we receive this final paperwork, we will send to Mark Hardin.

04/11/05 – ICS completed As-builts and sent to ThermalTech and TE sent to MWH in middle march.

04/15/05 – Received final pay application and closeout documents. Signed and sent to Hamco.


Courhouse – Video Surveillance Project

07/12/05 – Stapleton turned in final pay application, change order and final paperwork. TTC will review this week and get down to the County.

12-20-04 – Construction meeting #2 see meeting minutes.

12-02-04 – Construction meeting #1 see meeting minutes.

11-0-04 – Pre construction meeting took place, reviewed standard county items.

10-05-04 – AIA Contract for Stapleton typed and mailed to Sharon Chamberlin.

Project bid open had taken place. Mike Stiles has contacted Stapleton Rep. and confirmed the package and Sharon has authorized the contract to be prepared. Mike Stiles requested a copy of Stapleton’s bid from JoAnn in purchasing. Waiting for her to let us know when its available for pick up.

01-20-05 – Construction Meeting – see meeting notes below.

03-07-05 – Construction Meeting – see meeting notes below.

04-08-05 – Completed punch list and sent to contractors to complete – see list below.

04-29-05 – Received call from Paul Wilson with CSI. He indicated that the items on the punch list are complete. He also indicated that their time exceeded the bid amount for setting up the NAS device and switch. He request a change order for the additional time. TEI indicated to submit the additional hours and money requested for review by HC and TEI.


Courthouse Archive Room Renovation

08/08/05 – TTC received and approved the final paperwork for CH Archive from Ginter. Warranty and other paperwork received. TTC will hand deliver on 085/09.

08/01/05 – Diana called today because only 2 of the closeout documents have been received. TTC is checking with Ginter.

07/11/05 – TE received Ginter’s final paperwork for this job. TTC will review and sent to County.

07/07/05 – TE and PHB worked out the payapp details today by phone. TE had payapp #4 in house but wasn’t logged into the website. Issues are now resolved. Final paperwork will be sent to the county on Friday.

07/05/05 – TE received final pay application for project. There are some details that TE has to discuss with PHB before TTC will submit to County. Owner’s address is incorrect and has been on previous apps – TTC missed that. The previous payments don’t add up. Line 7 (previous payapps) doesn’t match Line 6 on previous payapp. TTC will stop in at PHB in morning and go over with Lee Anne.

06/21/05 – PHB (LeAnn) called to say they are sending a revised final payapp to closeout the project. TTC coordinate final numbers and instructed the to apply for final.

06/15/05 – Dorn Sprinkler Invoice arrived for $220.00 to clean the sprinkler head covered with insulation. TE will mail to Diana today.

06/10/05 – Peck sent a Word document confirming final allowance usages. TE instructed them to sign the change order sent and send back all paperwork. Bernie said he is on it. TTC emailed Dorn Sprinkler about invoice for cleaning the sprinkler head.

05/24/05 – TTC and PHB discussed final paperwork and O&M’s. PHB is finalizing project for closeout. Paperwork should be received in a week or so.

05/11/05 – DES advised to closeout the project and have PHB return the balance of the allowance. TE will write the credit change order for everyone to sign of $1358 and send for everyone to sign. TE also updated the allowance usages for this project below.

04/26/05 – Ginter and Honeywell ran the FA test this morning. Everything passed with flying colors. Larry Keifer asked us to replace one sprinkler head that had some spray foam insulation on it from the ceiling work by JB Schmitt. TTC called Dorn and instructed them to get with Mike Bartlett and take care of the issue ASAP. TTC told Ginter to send in all their final paperwork for the job for approval.

04/22/05 – TTC contacted Bernie about final paperwork and O&M’s for this job and finalizing allowance work too. Bernie said that the would work on it right away.

04/18/05 – Ginter called today to inform us that the city has finalized our final inspection date of April 26 at 6:00 a.m. at the site. JC Schmitt does not believe they need to be there so TE, Ginter, Honeywell and Mike Bartlett will be required to run this test for approval. All the work is already complete and the pretest already passed. Please make arrangements for this test and I believe Jim Ginter already called to let you know about this test.

04/11/05 – TTC called city directly today to inquire about status of permit. It has been two months since we submitted engineering change to this project and TE pointed out that we need get this project finalized. TTC is waiting to hear back from city. City mailed to Jim Ginter on 04/08/05. We need this permit to call for final inspection according to Amit. Jim did you receive the permit in the mail yet?

04/05/05 – Permit Number for Engineering Change to Fire Protection: 2004P07997.

04/0405 – Ginter called and got approval on the permit from Amit Gosh. Once more approval or actually waiting for the city to call us to tell us to pick up the permit – we will be ready to go.

03/24/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/16/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/03/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

02/22/05 – Ginter reported that all new fire alarm devices are installed in the Archive Room. Joe Dale of Honeywell is working on the new Fire Alarm permit drawings which have to be submitted to the city as an engineering change to the existing permit drawings. Mike Stiles will assist Ginter if needed to submit these plans. Honeywell was supposed to turn in these drawings on Mon or Tues of this week to the city. They are not turned in as of this email and Jim Ginter is calling Honeywell with word that we want them turned in ASAP so that we can pay out the GC in two weeks and have final inspections complete. TE received signed Change Order #002 from Ginter today.

02/16/05 – Ginter reported that new fire alarm devices are being installed today. They will coordinate with building officials on final inspection. MS wrote a change order for Ginter to have Copp provide the power supply, ~$700, that was left off of the door hardware schedule.

02/10/05 – TE has coordinated with the City and Inspector. Ginter has submitted the new fire alarm permit drawings to the city. TE will attempt to keep everyone apprised on the new status.

02/09/05 – GBBN issued punchlist to JB Schmidt for this project.

02/01/05 – Change order #01 for Ginter was issued by ThermalTech to Contractor and County.

01/20/05 – Final Inspection scheduled for February 3, 2005. Building Inspector and Fire Marshal should be present. 6:00 a.m. Larry Keifer requested additional strobes. Te to coordinate with Honeywell to get this work done.

01/18/05 – The Archive room passed its pretest this morning no problem according to Mr. Bob Ginter. Honeywell has to hook up the outdoor air actuator still and that won’t happen until next week according to joe dale (via b ginter). then we can set final inspection. Peck has to modify the outdoor air duct so that it is not connected directly to the unit and blank off penetration at unit as well.

01/11/04 – TE checked on job progress today. Peck reported that unit could not be completely started up because two receiver tanks were missing from the condenser on the roof. Peck found them in their shop and brought them onsite. They will try to have their crew at the building on Wednesday to install the tanks and the associated piping. Peck also noted that the OA duct is connected to the return plenum at an access door that will make unit maintenance difficult from that side of the unit. TTC discussed with Mike and conferred with other TE engineers and there is no reason to bring in 200 cfm of outdoor air for this space. TE will direct Peck to remove the OA connection at the unit and just terminate the outside air duct at the wall with a grille. This will allow a small amount of OA to enter the space to meet the one person 20 cfm requirement by code. If in the future, we need too, we can add an additional exhaust fan to the OA to bring in small quantities directly into the room. This is a better design and use of the equipment in my opinion and the factory startup technician agrees. Please advise TE that this is acceptable to the County.

01/13/05 – Ginter faxed pencil copy of revised Application #1 which has been approved.

01/12/05 – Called Ginter and Spoke with Michael. Michael remembers the problem with App. #1 and didn’t know that the Application was hold and should be resubmitted. Michael thought that it had been sent to Hamilton County and was upset that he would have to resubmit and wait 30-60 days for payment. Michael asked to revise Application #2 combining the Payment Due amounts of App 1 & 2 and resubmit the Application as #1. We’ll skip submission of Application #2 and use the already submitted application #3 which will be on hold for the final documents.

01/11/05 – Ginter Applications #2 & 3 reflects previous payment for Application #1. Application #1 has been on hold since November due to the incorrect change order amount. Contacted Ginter, spoke to Jim Ginter and explained why I could not submit Application #2, he said that I should talk to Michael who is not in the office and that he would have Michael call me.

01/10/05 – Ginter Pay App. #’s 2 & 3 received.

01/03/04 – JB Schmidt is installing the doors and security hardware on Monday and Tuesday and Ginter will hook up the security right afterwards.

12/30/04 – TTC & MB walked the site and MB pointed out that the Danfoss valves being installed may cause more trouble than they are worth as the piping is sold old and the connection to it could result in steam leaks in the piping below the concrete slab. TE will direct PHB on how to proceed

12/24/04 – TTC talked to Peck for and update. The piping and sheetmetal guys should be back on the job on Monday, 12/27. They are currently under the gun at the Quest building downtown and this is pushing our schedule back. TE emphasized that this project is scheduled to be completed at the end of the year and Doug Turner said a significant portion of the remaining work could be completed this week. Hopefully they do make it on site and a lot of the ductwork and piping gets completed before our next construction meeting. TE has to contact Ginter because communications questioned the width of the cable tray Ginter installed. I don’t see making it bigger now, but let me see what they say. All of the remaining HVAC equipment including controls and steam “donfoss” type valves are on site or Peck has them for installation. Mike Bartlett questioned the rubber pads under the feet of the Liebert units. TTC will check into this item with Peck.

11/29/04 – Liebert unit and crane delivery scheduled for this Saturday. TE forwarded request for CH personnel for this weekend work onto the County. TE requested that TE, PHB and Mike Bartlett get together later this week

11/10/04 – TTC coordinated with Peck on the feet of head available from the condensate pump on the Liebert unit and what is called for on the drawings. Pump has 20 ft of head so overhead piping can be ran about 15ft above the floor and along the water line route to the shaft and into the drain per the drawings. AC unit is still schedule to be delivered the 3rd week of November. Peck will still be coordinated the crane and equipment lifts for that weekend. If anything changes TE will let everyone know.

11/09/04 – Ginter Pay Application #1 received with incorrect Chg. Order Amount Jim of Ginter notified and will resubmit.

11/09/04 – Ginter Change Order #1 sent via courier to Ginter for signature and then to Hamilton County.

10/21/04 – Mitch from HCCD at last meeting requested an EMT conduit or cable tray from existing telephone data closet to archive computer room. TEI to get revised sketch to Ginter Electric for pricing.

10/08/04 – TE visited the site. We noted the walls are up to the deck, the caps are on the ceiling cavities per our last meeting. The Liebert HVAC unit is not expected to ship until late in November. Pet.

09/30/04 – Rough in inspection complete for Electrical, Some supports installed, waiting for GC to install doors for door contact work.

09/15/04 – Construction Meeting complete. Project on schedule. See meeting notes below.

09/13/04 – Submittals received from PHB on Mechanical Equipment.

09/13/04 – Permit ready. JC Schmidt to pick up and pay for with reimbursement from PHB

09/03/04 – JB Schmidt on site. Moving studs up into space.

10/08/04 – Sheetmetal caps were installed on existing openings in ceiling.


Courthouse – Mediation Department

07/14/05 – Apparently a change order for the VFD drive was never encumbered and this has prevented Reupert from getting paid. Diana said that Mediation was going to work on getting it taken care of.

03/25/05 – Reupert submitted closeout documents and paperwork for final pay application for project. TTC approved and sent to Ms. Supe.


CH – Pumped Condensate Project

04/20/05 – All final paperwork arrived for TE approval and submission on this project.

03/16/05 – Process will have final paperwork and payapps in this week.

03/14/05 – TTC contacted Process about final paperwork and payapps.

02/01/05 – Process contacted County to state that project is complete.

01/25/05 – Project punchlist sent out by TE.

01/18/05 – Process has completed pumped condensate work including the summer boiler relief vent. TE to stop by and punchlist on Thursday.

01/14/05 – TE approved 1-1/2″ insulation on all condensate for this project. AKM approved a $1600 invoice for Process to tie an existing summer relief vent into and existing 1-1/2″ pipe with a ball valve. Process is to submit an invoice for this work to Mike Bartlett today for approval so they will not have to pull of the job to do the work.

01/10/05 – Process came on site and began work on new piping for pumped condensate.


Courthouse – VFD Upgrade on CT-3 and CT-4

Courthouse – VFD Upgrade on CT-3 and CT-4 – SOSS TE-101 | 10700.283 | PM: TTC ME: TTC EE: RJ

08/08/05 – Final Pay Application was received. Allowance was straightened out. Ginter will keep $1000 allowance and pay DK $680 for additional programming work at the site to set the condenser water temperature to a constant number.

07/18/05 – O&M’s arrived to from Ginter Electric. TE will deliver to County offices on Tuesday by hand delivery.

07/06/05 – TTC delivered final paperwork and O&M’s to the County today. They were hand delivered to Martin.

06/29/05 – Ginter delivered the O&M’s to ThermalTech today. TE will give to the County.

06/16/05 – MB reported that the new cooling tower sequence is still maintaining a very stable condenser water setpoint. All parties are very happy at this time. TE emailed Ginter about final paperwork and punchlist work.

06/06/05 – TTC emailed punchlist to Ginter for final work and sent closeout documents with it for final pay application.

05/20/05 – TTC completed punchlist.

05/10/05 – MB reported today that VFD startup was completed by DK and Stoermer’s Gary John. The County reset the vibration safety switch and cleaned on of the low level sump alarms and CT-3 and CT-4 respectably. Ginter said they would check the wiring if that didn’t fix the alarm issue at the panel. The new program to control CWS/R temperature is in place and working from all first test. As the County brings on all 4 towers

04/22/05 – TTC and Dennis Faul worked on finalizing the cooling tower programming. MB and Mike Denton approached TTC about concerns about cooling tower control algorithms especially where concerned about the screw chiller. To that end, TTC has revised the cooling tower sequences such that a steady state condenser water will be more achievable. It is my hope that this alleviates many of the concerns expressed about this cooling plant. Here is the revised control sequence for approval and comments PDF. This should make commissioning much easier as this control strategy is much easier than the original.

03/15/05 – Programming is partially complete. Testing of program must wait till warmer weather. Project is in a holding pattern until that time. Ginter confirmed today that this was the


02/01/05 – TE ready to punchlist these projects within the week.

01/25/04 – Dennis Faul returned TE’s call yesterday. The DK technician is onboard running the electric for new controls. We cannot test the system until we reach summer conditions but DK is still going to put program in soon.

01/21/05 – TTC contacted Jim Ginter of Ginter Electric in regards to 11,640.00 controls work which was not completed and should not be included in the payment due amount of the pay application. Michael Ginter delivered a revised Application #1.

01/13/05 – Ginter faxed Apps 1 & 2 to TEI for review prior to mailing. Changes required to both. Marked-up and faxed back to Ginter.

01/12/05 – Called Ginter and spoke to Michael. Asked him to resubmit both Apps 1 & 2 with allowance included. Michael to redo & fax copy for approval.

01/11/05 – Called Ginter, Jim Ginter said I would need to talk to Michael who is not in the office and he would have Michael call TEI.

01/10/05 – App 1 @ 22,024.80 and App 2-final @ 1,915.20 rec’d. w/o allowance included in contract price or on Sched of Values. 01/03/05 – Ginter provided a Change /Order deduct of $458 if we allow them to leave existing disconnects in place. They are worried that new dimensions will force changes to existing PVC conduit and this could result in problems with breakage of brittle material. County to advise TE on how we want to proceed.

11/18/04 – TE approved the design of all 4 cooling towers being routed from the main MCC in the penthouse. TE will provide new stamped design drawings for Ginter to submit to IBI showing the new work.

11/05/04 – Pre Construction meeting.

10/27/04 – Contract approved, TE to setup kickoff meeting.

09/24/04 – Mailed Contract to Tony for approval.

Prebid complete. Addendum 1 distributed.


CH – Relocate Basement FCU

03/30/07 – DK called today to inquire why they weren’t paid their final monies on the bsmt FCU project. TTC resent the 07/14 data to MWH.

07/14/05 – Received final invoice from Debra Kuempel today. It will be voided because of billing errors.

04/11/05 – TE talked to Michelle Combs today. DK sent an invoice in December for this project. TE does not have a record of this receipt. Michelle will resend a full invoice today for TE approval. TE will send to MWH immediately afterwards.

11/18/04 – Project complete. No punchlist items to complete. All units are running and trades finished up wall and electric work.

11/10/04 – TTC talked to Ed Enderle today. DK is hoping to be 95% complete by the end of the day on Thursday. TTC will visit the site and check on progress. DK is also going to seal up the walls during this project. HamCo will have to reconnect the electric if they haven’t already done so before units are operations

11/08/04 – TE visited the site on Friday afternoon. Fan coils are relocated and piping is in the midst of being changed over. Ductwork has been removed from through wall condition. Wall penetrations are open. Ceilings are partially disassembled for FCU relocation, ceiling tiles are stored inside the rooms.

10/21/04 – Pre Construction meeting complete. Debra-Kuempel scheduled to begin the first week of November. See notes below.

10/14/04 – TTC messed up the pre construction and it has to be rescheduled. Debra-Kuempel and TE will be onsite on October 21 at 8:30 am for the next meeting.

Bidding is complete.

County: ThermalTech is awaiting Notice to Proceed


Courthouse/Justice Center – 2003 Boiler Plant Upgrades

11/02/05 – Triton called to say that the certified payroll will be delivered tomorrow. This should wrap up the project completely.

10/18/05 – Triton’s final payapp came in. TTC approved and will give to Tony Matre on Wed via hand delivery.

10/03/05 – The valves and spool piece have been replaced. Triton is getting an insulator down to the site for final work. The final pay application is en route to ThermalTech for final approval.

09/22/05 – Triton is scheduled to replace the defective valves in the Courthouse this week on Saturday. They will being working on Saturday morning. The CH will turn off the steam on Friday to allow the plant to cool down.

09/13/05 – Triton came down and measured for the spool piece today. Mike really wants it installed by 10/01 and Ed Makin said they would try for two weeks. Mike asked about putting back an isolation valve in the spool. TTC said that TE prefers to keep the main wide open with no restrictions plus the pneumatic valve provides isolation.

09/12/05 – Today we kicked of the low pressure steam flow meter project. Process will do all the work during the first week of October. Triton is coming down tomorrow to measure the spool piece for the Courthouse valve replacement.

07/18/05 – TTC contacted Brian Pastura by both email and voicemail today to discuss resolution on the boiler plant valves and back pressure regulator. No return contact as of this email. TTC will continue to follow up with Triton.

07/06/05 – TTC found a quote from JCI to control the bypass valve on the supply to the JC. Based on the Spirax/Triton quote this may be the best solution to the steam supply issue in the

wintertime because the full flow works as proven last year through the bypass. By automating this valve to try to match what the JC needs, we can control total flow and still have the shutdown of the spirax valve to prevent carryover. PDF. TTC sent out Add01 for the Steam Flow meter project. The new bid opening date is July 12 at NOON. PDF

07/05/05 – TTC emailed Triton and Spirax trying to get the repair price out of them for the steam valve and to find out how to independently control the valve without DDC as TTC requested and Mike Bartlett questions (TTC did not have an answer for).

06/15/05 – TTC emailed Alpha Controls about new bid project: Mike, We spoke about 3 or 4 months ago about replacing the existing flowmeter at the Hamilton County Courthouse. I now have to bid out the replacement of the flow meter at the Courthouse. I will need the specs on the valve that you believe we should install. If you remember our final decision was to abandon the valve in the vertical and go with a steam flow valve downstream of the PRV in the low pressure supply to the building. This will allow us access to a large header from the floor with significant straight length. We will provide all the bid documents for three to five mechanical contractors to give us pricing on the job. I will ask them to hire you to provide the meter and do the control hook up. I will need a budget price from you now for your portion and I will need your help on spec’ing the right model number and getting the right flanges and connections in to accomplish what we are trying to do. You will have take the existing control leads and reconnect them to the new flow meter we are providing. As you already know – we don’t need exact lb/hr reading for some sort of process control but rather a flow meter that can give reasonable accurate readings from 0 – 10,000 lbs/hr or so.

06/15/05 – TTC and Pete Rorison talked by phone today. The existing valve will not and can not work under the existing conditions. Spirax can provide a Globe Control valve that can work in this situation. They cannot however, provide this valve at no cost to the County. Pete said the original valve was designed at what they believed was the correct design conditions at the time. Spirax will provide a new control valve, actuator and positioner and re-use the existing pilot positioner tap for control. Either the County or Triton will need to run a new pneumatic line over to the controller/actuator for the motive force to operate the new control valve. TTC instructed Pete to coordinate with Triton and get a best and lowest price together to do this work. There will be a credit for returning the old valve in lieu of the one we may install. When TE receives this new price, we will pass it along with a recommnedation.

06/14/05 – Pete Rorison emailed TE back today. Many of the issues in his memo were not correct. TTC sent a response email to rebut what was said and to try to clear up exactly what Spirax/Sarco is ready to do to make things right. TTC will follow up with Triton tomorrow for final resolution.

06/10/05 – TTC emailed Brian and Pete with a final direction on the Steam Control Valve. TE told them to have a solution in place by June 30th. PDF

06/09/05 – Triton called today to inquire about status of project. Per TTC’s conversation with Tony Matre on 6-7-05 there needs to be some resolution on the Spirax/Sarco valve issue before final payment is received. TTC will call Pete Rorison to see if another valve will work. TE sent out an email to discuss with County PDF.

05/20/05 – TTC and MB discussed the findings of Randy from PHB. The final word is that the steam valve is working correctly. TE next will make a recommendation about increasing the steam pressure in the boiler plant up to compensate for the low flow through the Spirax Valve. Another option is to set the max flow with the bypass open a certain % (field verified) to meet the max demand of the JC in the winter time. PHB stated that the 20:1 spring is the correct spring for the pilot and that the pressures seen on the inlet and outlet of the valve are correct for this valve.

05/10/05 – Triton’s revised invoice arrived. TTC discussed with Jim Drye the steam valve and decided to talk to Rich DeBat, a former Armstrong rep, about this control valve. TTC also contacted PHB about troubleshooting the existing valve. TE had Bernie provide a name of the their steam control valve tech (Randy) who is scheduled to meet Mike Bartlett on Friday at 6:30 a.m. to look at the valve and maybe do some troubleshooting for us. I told him to bill ThermalTech for now (as a reimbursable expense for now) We can judge how much of this time and work we need to help us out. After he gives us a synopsis, TE will provide a permanent fix for the boiler plant and we understand this has to happen this spring/summer.

04/20/05 – Triton and TE discussed in great deal when Triton is going to get paid for their remaining outstanding balance on this project. TE sent an email to Tony about what Triton wants to do – bill the whole contract now at this point. We are waiting on the County to decide about this.

04/19/05 – AKM & TTC talked about how to resolve PRV issue and Tony said that he has $10,000 to get these final two issues resolved through Triton Services. They will charge to the existing PO’s that Tony has available for this job. TE promised to have these issues taken care of this summer before winter ever gets here.

03/25/05 – TTC and County discussed solutions for the PRV valve since Spirax cannot provide a valve big enough to handle our load. TTC wants to bring in another engineer to discuss placing a second control valve in line to meet the maximum demand and still protect the Courthouse from steam spikes of the Justice Center.

03/14/05 – Triton delivered all final paperwork to ThermalTech for this project. TE to review and deliver to County.

03/11/05 – ThermalTech, County, Triton & Spirax/Sarco met to discuss the regulator valve issues at the CH. Several problems were discussed and solutions were also discussed. The bottom line is that the valve is not sized large enough to handle the full load required at the JC and still maintain the steam pressure required at the JC. Per Mike Bartlett’s excellent suggestions, TTC is pursuing JCI about automating the bypass valve to make up for this discrepancy. More info is coming in the near future to further discuss this issue.

03/08/05 – TTC has spoken to Brian Pastura about getting chuck on site to make these changes and they will do what they can but would like to have a meeting between TEI, Mike, Tony and Triton to decide on how to handle the workings of this valve for now and in the future. i agreed with brian that this would be a good idea and asked him to set up this meeting.

03/07/05 – Mike reported that Alpha’s service called yielded some design and equipment problems with the existing flow meter: Mike Maloney from Alpha Controls came in Thursday to check the Courthouse Yokogawa Steam Flow Meter. He found the sensor not working and recommended replacement. He estimated a replacement cost of just the sensor at $1500.00. After looking at the meters specifications, I found that this meter does not read steam flow under 3000 lbs/hr. An 8 inch valve is set up for reading between 3000 lbs/hr to 40000 lbs/hr. We believe that the Courthouse steam flow would be no more than 10000 lbs/hr. He recommended downsizing the valve to 4 inches and redoing the piping to allow for more straight piping that is recommended. He says that right now the 8 inch meter does not have enough straight piping for correct flow readings. He estimated a complete 4 inch meter at $2000.00. TE recommends doing this work to get this additional flow added to the Master flow readings on the front end for use and troubleshooting reasons.

02/24/05 – Pete Rorison called today to basically say that ThermalTech gave him the wrong design conditions for the valve and that it will meet the design conditions of 65 psig inlet and 50 psig outlet at 16,000 pph. The problem is that the JC suffers and 8 lb pressure drop at full flow so the valve should be sized at 65 psig inlet and 58 psig outlet. ThermalTech sent their rebuttal and the Submittals data stating the the valve was to operate between 50 – 80 psig.

02/15/05 – TTC contacted Triton and Spirax about pressure regulating valve and its proper workings. We will advise when we have more information.

02/01/05 – TE contacting the manufacturer about valve failures. Plants scheduled to be shutdown and worked on when OAT is above 50 degree F.

01/25/05 – Triton tried to fix the pilot positioners but the Powell Steam Valves did not hold. TE requested that Triton contact the mfgr and get new valves to replace the defective valves. The spec and the contractor wanted positive shutoff valves in this system. MB may talk to JC about a steam plant shutdown to replace the valves and do the pilot control work on the steam pressure regulator/backpressure control valve.

01/14/05 – Tim Cancilla is supposed to be onsite today to readjust the controller valve for correct steam flow and to rectify pilot tubing installations.

01/12/05 – Called Tim @ Triton to verify final doc list rec’d. Tim said finals will take awhile to get together and that he needs a Affidavit of Contractor Prevailing Wage form (2 Affidavits mailed), App 5 sent via courier to AKM.

01/11/05 – Revised Triton Apps 5 & 6 rec’d, 6 on hold for final docs, 5 signed.

01/10/05 – TTC request I notify Triton of incorrect contract amount on Apps. Tim w/Triton notified that Apps 5 & 6 should be resubmitted w/correct contract amount & that 6 will be held for final docs, refaxed final doc list.

01/04/05 – Tim with Triton, faxed a pencil copy (Application No. 6) requesting $27,206.76 with zero balance. Pencil copy approved.

01/04/05 – Email AKM regarding status of Triton App 5 pymt.

01/04/05 – 010405 App 7 received w/zero balance ( Requested 9,571.00 Change Order amount). App 7 voided due to incorrect $’s on G702 and G703, Called & faxed mark-up of 7, requested a revised App be sent as App 6.

12/10/04 – Follow-up with Triton Services, Timothy Herzog regarding final documents. Mr. Herzog requested that ThermalTech re-fax the list of required documents (which was done in the am of this date).

120605. T. Herzog contacted & fax sent informing that App 6 on hold for final docs.

11/22/04 – Triton Services, Timothy Herzog contacted regarding final documents and a list of documents required faxed to Triton.

11/16/04 – Process has replaced the traps and they are insulating them. Kerry approved contract changes and TE will deliver the documents to Tony for final work, Triton will bring diamond plate down for installation and have their insulator finish up the piping on the boiler room. Mark Neary returned my phone call and I have to follow up with him on flow meter recalibration.

11/04/04 – TE contacted Triton about finalizing the punchlist on this project including the diamond grate, insulation, flow meter, summer boiler vent line. Process has been told to replace the steam traps asap even pending the outcome of the purchase order submitted. TE instructed Doug Doxsey that in addition to the traps and the pumped condensate lines that a steam trap to drain off the summer boiler may also be a change order the County is looking to get prices on.

10/27/04 – TE sent change order request to Process for steam trap replacement. PDF

10/22/04 – TE scheduled to perform punchlist on project. Insulator has wrapped up majority of work. Steel grating not installed yet.

10/21/04 – TE has to get a steam trap decision to Mike Bartlett this week. TTC has to send data to Michigan to get answer from our special Armstrong Engineering.

10/21/04 – TE and MB discussed having process add the summer boiler vent to the new venting scheme instead of floor connected.

10/20/04 – Three Original AIA 101 Process Construction Contract document and Original plus two copies of submitted contract documents, from Process, sent via UPS to Hamilton County Dept. of Facilities.

10/19/04 – Original Contract documents received from Process Construction.

10/14/04 – TE approved Change Order #2 and sent 4 copies to Brian Pastura for Triton signing and then onto Tony Matre. Also the insulator is making steady progress on finishing the rest of the boiler plant insulation work. Once complete TE will do final punchlist.

09/29/04 – Triton is starting on concrete work, Neary provided controls programming earlier this week and appear to have done a nice job, Mike Bartlett reported that a recent Justice Center steam spike of 7000# reduced JC pressure from 55 to 14psig but the CH plant went from 55 to 51psig thus the new regulating valve appears to have worked as the plant maintained nominal working conditions. TE to schedule their punchlist as soon as insulators, concrete and controls are complete by Triton.

09/24/04 – Tony approved bid for Process at $9680 for pumped condensate work. TE will write contract.

09/24/04 – Water hammer is occurring at the CR-2. Stoermer-Anderson and TE will have to study this scenario and advise the County as to what might be occurring and how to eliminate this effect. Theoretically with no “live steam” going to the receiver, we should not have steam hammer at this piece of equipment.

09/24/04 – Triton submitted the NO BID paperwork so the bid tab is correct. How should we proceed. Please advise.

09/23/04 – Triton updated TE on project status today. Concrete contractor should be done in the next couple of days. PCI should be onsite today starting insulation. Ed Makin will be back down to put the bypass around the city water to the CR-2 on Friday maybe.

09/23/04 – Bids are in for pumped condensate work. Process is low with $9680 bid. Bid Tab – PDF

09/13/04 – Triton has completed work inside the boiler room and removed all tools. Room has been cleaned in general. Outstanding items include: Controls work, final inspection, concrete work, lead/lag switch on CR-2, plumbers must fix softeners so that we can start boiler plant, air control regulator must be replaced, insulators must do their work, steel grating must be installed at entrance to CR-2 room.

09/08/04 – Inspectors questions flash tank drain going directly to floor drain. TE to coordinate in morning on 10th. Code states there must be an air gap of 1″. Existing condition did not have it installed that way. TE to resolve.

09/03/04 – SA instructed us to take the new 3″ pumped condensate into all 3 of the water inlets in the top of the tank. This is a changes from the design documents. Triton will have to submit a change order for this work.

09/02/04 – MB reported today that Triton installed the 3″ make-up from CR-2 to DA into a 3-inch opening without a spray head. After much discussion, it was decided that Triton has to install the piping such that the 3″ CR-2 feed must tie into (2) existing city water connections both using spray heads. The remaining opening can stay in place for an emergency city water feed control to the DA.

09/02/04 – TTC and MB decided that after the plant is complete, we will begin to run the plant at 65 psig just like it has always been designed to operator and therefore we won’t need the PRV’s installed at this time.

09/02/04 – Dorn contacted TE today. TE gave them a PO that will be reimbursed during subsequent invoice from TE to do work on moving sprinkler line on T&M basis in CH walk-in cooler.

09/02/04 – Mike Bartlett report that Spirax stated that strainers must go in front of the PRV valves. I checked with Todd and Gary and the HTX are rated for 125 lb service and they should be able to take the higher pressure in the boiler plant. I am recommending that instead of incurring more cost, we just return the valves and take the credit to the job to offset some other change order cost on blowdown lines.

09/01/04 – TE talked to Spirax engineers. Both PRV’s must be installed for HTX’s, flow meter can go in vertical, they will provide flow differentials based on varying steam pressure.

09/01/04 – Stoermer Anderson confirmed that summer boiler relief valve is already set at either 80 psig or 100 psig, so no additional work is required.

Triton Services will be back on site August 16 to begin finalizing the project work

CR-2 and DA piping will require a FULL plant shutdown scheduled for August 20th.

Summer boiler must be online independent of the main plant by August 20.

TE will coordinate with Stoermer Anderson for final Summer boiler startup details

Summer boiler will use treated city water during 3 week shutdown as work is completed.

TE and Triton to coordinate schedule with Mark Neary for new control sequences and SA for boiler work and equipment changeout and use.

County will provide an airline close to top of CR-2 so that project doesn’t incur additional change orders

Boiler plant work scheduled to be finished and back online by September 10, 2004.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: There are (2) 3″ inlet to the DA tank, can we remove the 2″x3″ bushings and run the pumped condensate directly into the DA tank. Response: The bushing and utilize one 3″ inlet for the 3″ pumped condensate from CR-2. Since all material are now steel, the existing dielectric flange for the brass fitting will not be required. Keep one of the existing connections in place and utilize the existing city water connection and controller as backup or emergency fill if ever needed.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Triton asked TE to locate the circuit setters shown on the drawings. Response: TE advised Triton to NOT install the circuit setters shown on the drawings and save the money to use on the blowdown piping.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Can 6-inch piping to steam regulator on catwalk be maintained in size. Response: TE and Triton met onsite to discuss field issues and many small items were resolved including the piping layout for the new back pressure regulator valve on the catwalk service the JC. 6 inch piping can feed the JC and increase to 8-inch immediately after the final isolation valve just like existing piping is installed. Existing Justice Center main will now be bypass. Per the contract documents Triton will replace regulator on wall with new equipment to operate bypass valve.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: How do the new control valves get place and installed on the project? Response: The new control valves for the CR-2 and DA-1 shall be installed per the direction of Stoermer Anderson. The existing city water feed to the DA can remain in place as on 3″ outlet can be used for the feed from the CR-2 unit and the other 3″ (bushing down to 2″) can remain as emergency feed to DA.

Vendor: Triton Services | Date Issued: 08/24/04 | Date Answered: 08/24/04 Question: The original separator location was in the Justice Center. Now the separator is being installed in the Courthouse. This is a change in the contracts. Please advise. Response: Brain, you are correct the original location was in the Justice Center but it was to be installed over 15ft high in the air requiring lifts and extra manpower to reach it and install it at the location shown on the drawings. All of this work is included in your base bid. It is my belief that the new location provides a significant reduction in manpower and installation time. Therefore I see no reason to incur additional cost to the project for this project change as we can just use that savings to apply it to the additional work to reroute the condensate line which should only include 2 90 degree elbows and their associated welds. The rest of the scope remains the same.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: There is one 3″ flange on top the DA that has significant pitting. Should this be replaced? Response: Please replace the flange at the weld neck. Make it suitable for the blind flange required in the construction documents.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: The separator has a 1-1/2″ outlet on it. The TE drawings show a 1″ outlet with a 3/4″ condensate drain. How should this be remedied? Response: Please route the condensate drain as 1″ to the condensate system.

Vendor: Triton Services | Date Issued: 08/26/04 | Date Answered: 08/26/04 Question: What pressures should the new flow meters be calibrated too? Response: Set the Courthouse flow meters to 65 psig and the Justice Center flow meters to 55 psig. Name them CH-FM3 and CH-FM4 and JC-FM1 and JC-FM2


Courthouse/Justice Center Separator Installations

05/04/06 – Tony called today because Geiler’s final payapps arrived unsigned by ThermalTech. TTC drove down picked them up, signed them, made copies and sent them back down to the County for payment. Geiler did not sign the copies so TTC just copied the original after he signed it. Once paid this project should be complete.

04/10/05 – Indrolect inspection at the Justice Center is finally complete and approved. We are now ready to finalize and archive this project.

02/17/06 – I don’t think the inspection will be next tuesday because indrolect just got to the bottom of Geiler’s permit applications. apparently geiler took out two permits even though they didn’t know it. the city doesn’t even have a copy of the jc one in their file but on the computer it says it was approved. ibi will accept the previous plans exam as valid for the work but now has to schedule another field trip to sign off on the jc wiring install. when ibi decides they can come out, indrolect will call us and let us know and I will call you and let you know.

02/114/06 – Dave Memory of IBI approved the CH install with TTC, MB and Don Ritter in attendance.

01/11/06 – Geiler and Indrolect have submitted for all permits. As soon as plans exam is finished, Indrolect will call for final inspection. Should be within a week or so.

01/01/06 – The job cannot be closed out because of permitting issues from the contractor. Indrolect and Geiler have been given the preconstruction meeting minutes where we state that permits are required on this project. TTC sent Geiler an email today saying that the additional $1024 in allowance money that is being paid to Geiler should be used to acquire all the permits on this job. TE will provide additional stamped sets of drawings as necessary for permits.

11/28/05 – TTC received all final paperwork and will deliver to Tony Matre on 11/30/05. This job is complete.

11/03/05 – Lisa Huber said that all paperwork has been sent to ThermalTech for this project. Certified Payroll will go directly to Tony after they get it on Friday.

11/01/05 – Finalized final billing strategy for final bill. Tony said they don’t have any certified payroll and they need final pay applications. TTC emailed everyone to let them know.

10/20/05 – Lisa from Geiler returned TTC’s phone calls about project closeout. TTC returned on 10/24 at 5:30 p.m.

10/18/05 – TTC emailed and faxed the change to Michelle Wells again in the hopes the final payapp will soon be processed.

10/14/05 – TTC contacted Geiler again for the correct final pay app. No work yet.

10/03/05 – Still no word back from Geiler on corrected pay application.

09/21/05 – We received the final pay application for this project as application #2 this time. This payapp is not right because of the change order amount is wrong versus the allowance. TTC contacted Geiler to advise them on how to fix it.

08/08/05 – TTC reviewed and will deliver the final paperwork to Tony on 08/09. We have not received a final pay request nor the signed punchlist. TTC emailed Barry Bowman to inquire about final pay application.

07/25/05 – TTC received in his mailbox the final paperwork.

07/18/05 – Contacted Barry Bowman about finalizing this project and sending all paperwork in.

06/10/05 – Relisted Punchlist PDF.

05/20/05 – After cleaning the system several times and removing large amounts of rust from the piping inlet to the separator and disassembling and reassembling the 4″ supply line to the separator, the team was finally able to start the separator and see significant flow at the cooling towers. The system when new provide 40 psig of gauge pressure at the outlet. TTC stated that with this much pressure and flow out of the nozzles that 2 towers could be scrubbed at once and the normal operation would be to run the separator with CT-1, 2 for a week then switch it to CT-3,4 for the next week. The punchlist will be sent but Geiler took care of issues at the site during startup. The strainer’s and filters should be changed a couple of times a day for the week or so. Monitor how much cleaning the SEP-1 unit provides and modify the cleaning schedule as needed.

05/17/05 – Startup for CH separator is scheduled for 05/20/05 at 8:30 a.m. with TE, Geiler and Trane.

05/16/05 – MB emailed today to say that all four CH cooling towers are online and ready for separator startup. TTC emailed the contractors and requested the dates. TTC approved application #01 for payment and us mailed to Tony Matre.

05/02/05 – Geiler promised that Bob or his associate would be at the JC today working on the punchlist items for the JC separator.

04/22/05 – TE contacted Geiler for a third time about completing punchlist work at JC. HS confirmed again that no punchlist work has been done. Geiler wants paid badly and keeps inquiring about payment and TE keeps pointing out deficiencies in payapp procedure. TTC believes that we have it resolved now. CH staff believes that their towers will be online next week for startup and testing.

02/11/05 – Punchlist distributed to Barry, Tony, Howard and Mike.

02/10/05 – Geiler asked to see the original meeting minutes where we told them to put a drain and valve into the Courthouse Separator system so it could be used in wintertime operation. TE to send out from original pre construction meeting.

02/01/05 – TE ready to punchlist these projects within one week. Punchlist is complete. TTC to distribute.

01/25/05 – Indrolect made the tie-in into the bus bar on the elevator equipment room yesterday without incident. We powered up the disconnect and then removed the fuses awaiting startup by Geiler. TE contacted Barry about status of project and if it was ready for punchlisting.

0/1/21/05 – No response to the previous day conversation. Followed-up with a fax to Geiler about Application #1.

01/20/05 – Pay Application #1 placed on hold. The Application was not signed or notarized and also did not include a completed, notarized, Affidavit of Original or Subcontractor. Michelle Wells contacted and told of the reason the pay application would be put on hold. She asked for a voice message to be left for Lisa explaining why the application is being put on hold (which was done).

12/03/04 – County approved Geiler change order allowance usage for redesign of PVC system for separator. $3976 will come out of the allowance on this job.

11/22/04 – TTC stopped at the Justice Center this morning and reviewed the existing piping layout with Joe Phillips, Howard Seal, Barry Bowman. What we are proposing to modify the design so that each tower can be isolated from the system and cleaned to prevent any of the “dirt” from the system from getting entrained into the condenser water supply to the chillers. Each separator could be piped separate with PVC and ran an determinant amount of time to clean the tower daily and then return to normal operation. TTC will provide a revised drawing to the Geiler company for this work today.

11/12/04 – TE contacted Geiler about completing the Courthouse and redesigning the Justice Center to meet actual design request. TE will follow up with Geiler.

11/04/04 – TE contacted Barry about field progress on this job.

10/20/04 – Geiler needs to install a drain and valve on CT-2,3,4 mains for winter draining. Install a drain on the main unit. Possibly relocate the existing control panel 90 degrees rotated to meet NEC requirements.

09/29/04 – CH CT 1 & 2 piping is complete. CT -3 & 4 and JC work to be scheduled in the upcoming weeks.

09/24/04 – Geiler estimates that the courthouse should be done is 1 – 1.1/2 weeks.

09/23/04 – TE emailed Geiler to inquire about project schedule update since progress is slow.

09/13/04 – Piping is being installed in Courthouse cooling tower #1. Looks good.

08/01/04 – Geiler asked how do they fit piping under fill? on 08/10/05, TE advised to decrease manifold size to 2″ and slide piping under existing fill. This will maintain integrity of cooling tower. TE provided a new highly detailed dwg of the tower piping for Geiler to use on all towers for this project.

Geiler has been onsite doing the work.

ThermalTech received written approval for sweeper design.

2″ pipes will be slid under existing fill.

Contract should be back on site soon.

Cliff Wingerter is on medical leave so all questions should be routed to Barry Bowmen.


Courthouse – Steam Flow Meter Project

01/03/06 – TTC received final paperwork on this project. TE will process it and turn it into Tony for closeout on this project.

12/02/05 – TTC contacted Process again for final paperwork.

11/28/05 – TTC was at the CH today and discussed with Mike Bartlett final job status and any punchlist items. MB did not report any further punchlist items and TTC did not record any incomplete from original walk through that asked for the insulation to be complete and the programming and scaling of the steam flow from the new meter to be re-adjusted. These issues are complete and therefore no further punchlist items are required at this time. TTC approved allowance usages to repair a defective gasket on an isolation valve during the project for $1000 which included some miscellaneous other travel charges and such to get parts for this replacement and a start/stop scenario while ThermalTech gathered information to allow replacement to take place. TTC will contact process for final payment.

11/08/05 – Mark Neary reprogrammed the Teletrol system, Mike Bartlett confirmed the flow readings were correct. Insulation on the old steam flow meter spool piece is not complete yet but everything seems to be working fine.

11/07/05 – Process sent in pencil copy payapp. TTC contacted Mark Neary for online programming of the Teletrol trying to save paying a trip from Columbus for work.

11/01/05 – Project is complete aside from reprogramming the Teletrol system to the new range of inputs from the low flow meter. TTC is working with Mark Neary to get him to dial in and change the expected input to match the actual flow. Mike Bartlett reported today that the partial flow was around 1900 lbs/hr.

10/24/05 – TTC emailed Mike & Clem to get project update.

10/21/05 – Mike reported that the main steam flow was back up and running for the JC. The flow meter eccentric reducers were being welded in. One was already installed.

10/18/05 – Process is all set to begin and finish work this week on the project. The steam plant is down and cooling off. TTC will stop at the site on Wed morning to coordinate and then go to a prebid on a different project. Process will attempt to do the work on the existing steam meter first and get the plant back up and running on Wednesday.

10/14/05 – Personnel issues are preventing Process from starting but Clem reported that they can start on Wed 10/19 at the CH.

10/04/05 – We were trying to be done with the Steam Flow meter the first week of October. TTC has contacted Process about availability of parts and getting this work completed.


Courthouse – LIP Schematic

10/14/05 – TTC has inquired about status of project. If the new door is being handled within, then this project is complete. Diana and Mark please advise as to the status of this project.

05/12/05 – We all discussed alternate schemes for a new door to meet state and local requirements and TTC sent this color scheme with 2 options to GBBN for review and for GBBN to forward to the County for approval. The County shall make the final decision as to which way they want to proceed. PDF

04/26/05 – TE recommended provide a new door in the space to meet code requirements from state and city per this PDF. TE must wait for County to approve a way to go on this issue before project is complete.

02/2/05 – TTC talked to Mike Bartlett and found that complaints have subsided since Greg greatly reduced the total amount of air being fed into the space. This makes sense since the space was being overcooled because the thermostat was on the second floor in the courtroom. Mike and I agree that this should work until warmer weather where the airflow into the space will have to be increased and further substantiates the need for a revised VAV system in this space. As it right now, Mike and Greg are acting as a manual VAV system by increasing and decreasing air into the space as needed. This also shows us that if you are really crunched on money, you may be able to only put a couple of VAV boxes into the first floor area and begin to achieve some control with the hopes that a full redesign can be done later.

02/25/05 – Client reported that space is still not conditioned during the evening time properly. The daytime loads can still never be fixed without a new VAV system and HW reheat (see SOSS TE-115)

02/10/05 – TE has not seen a balance report or payapp for the additional work.

02/01/05 – MB reported today that Room 169 reported no comfort problems with the space except when the main ahu turned on. PHB still has to balance the new air device. TE to punchlist this job shortly.

01/31/05 – MB reported that PHB finished the job last week. TE has to confirm. We will not be able to assess whether this project was a success as of yet because the units are being operated 24/7 in that area due to other heating and cooling considerations.

01/24/05 – Peck started the job. They need a different transition and will be fabricating that in their shop and finishing the job in one or two days.

01/18/05 – Peck said they should start on Wed or Thur, Jan 19, 20 per Ken Shad.

01/03/05 – Peck should be onsite at the end of the week or beginning of next week to start after hours work.

12/24/04 – TE has given Peck a scheme to price up. Assuming they have the best price, they are the odds on favorite to do the work since they have an unused $5000 allowance at the CH in the Archive Room.

12/10/04 – Price came back from B&J at $6000, TE is now inviting Peck to give us a price since they have allowance money left on 2 County jobs. TTC will bring down Peck for a price quote next week. We will coordinate the visit with Mike Bartlett.

11/05/04 – TE met with Mike Bartlett, Bill Scholl & Jim (Bailiff’s) and reviewed the general layout of the 5-ton split system into the Bailiff’s area but the actual unit would be in the Room #167 space (this would have to be approved by LIP), ductwork would have to routed above the 167 space and over to the 169 area, 6 diffusers at 250 cfm would have to be installed to meet the 2000 cfm requirement of the 5-ton unit with a return ductwork. the final design is still under review. TE has asked B&J Jacobs to take a look at the design to provide their opinion about what is easiest and most cost effective. TTC will provide prices associated with each of these schemes so that the County can make an informed decision.

11/04/04 – TE discussed a ductless split system installation with Mike Bartlett, Bill Scholl, Diana Supe and left a message for Tony Matre in the 24 hour bailiff’s area today. TE will follow up with Tony about funding this project ASAP.

10/25/04 – Bill Scholl called today. A full size duct won’t fit through above the ceiling into the bailiff’s area due to a water pipe on the other side of the wall but a duct sized at 12″ x something will. I told them to find out and maximize the size of this duct and proceed with the design. We are trying to get as much air into the space as possible without spending much money. This is the first stage of the “fix”. Other things may have to be done later to enhance the space control.

10/14/04 – Mike Bartlett showed TTC B-12 which is actually the unit that serves the Concessions area. This unit is served by both heating hot water and chilled water. There is a temperature sensor in the supply duct outlet that monitors the leaving supply air temperature. TE believes that this is the sensor that Bill Scholl was talking about in our 10-12-04 meeting. When there is no building chilled water, the DX unit provides cooling for B-12. Our drawing and scheme will still work as the timeclock will divert air to the room #169 space starting at 7:00 p.m. – 6:00 a.m. and regardless of how B-12 is providing cool air it will get into 169 to assist in cooling the space. The system should still be modified as design on the ThermalTech packet.

10/12/04 – TTC met with DES, MB and Bill Scholl to discuss the reworking of the ductwork on unit B-9 to included Room #169 during the evening time. Trades and HVAC will do the work in the next couple of weeks. The plan will have B-9 provide cooling to the space after hours between 7pm and 6am. TE also stated that the ceiling work required should be done after hours as this will be a messy job especially in the kitchen area above the sink. We will monitor the temperature during the first few months and see if the County needs to add additional controls to the new air damper later. Bill reported that a Paragon timeclock will be used as they have one of those instock and the damper and actuator have been ordered. TE is fine with this information.

10/08/04 – TEI received a letter asking a question from the plan examiner. MS needs to talk with him because a letter will answer his question. However, he is out of the country till the 18th. They directed me to his supervisor Bill Spurling. TE have been playing phone tag with him for the past couple of days.

10/07/04 – TE visited the site and TTC has sketched up a possible solution to the Room #169 heating issue. TTC will deliver to the County and look at making the changes utilizing existing equipment for the work on 10-08-04.

09/13/04 – Fire Test at 6:00 a.m. Space did not pass. Egress into LIP has been questions by Larry Keifer. Security drawing was not included in permit set to City. Mike Stiles to rectify ASAP.

09/03/04 – TE delivered a set of permit prints to Brian Meyer for use on this project.

Security issues are ongoing between LIP and Recorders.

Trades doing all work on this projects.

10/18/04 – TEI received question from plan examiner and TEI responded. Plan examiner was out of the country until Monday October 18th. TE waiting for response.

4/19/05 – TEI recieved call from Mark Harden indicating that the State Agency that governs the computer system information will not accept the security level provided by the rear door in the computer room. However, the motion sensor that was added was per the the state building code. Mark and TEI to discuss options. TEI contacted Ben with GBBN to discuss the option of another 2nd path of egress. Ben to respond.